### v2

# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

Abstract
This work proposes a policy learning algorithm for seeking generalised feedback Nash equilibria in $N_{P}$-players non-cooperative dynamic games. We consider linear-quadratic games with stochastic dynamics and design a best-response dynamics in which players update and communicate a parametrisation of their state-feedback policies. Our approach leverages the System Level Synthesis framework to formulate each player’s update rule as the solution to a tractable robust optimisation problem. Under certain conditions, rates of convergence to a feedback Nash equilibrium can be established. The algorithm is showcased in exemplary problems ranging from the decentralised control of unstable systems to competition in oligopolistic markets.
Index Terms:
Non-cooperative games, Feedback Nash equilibrium, Best-response dynamics, System level synthesis
I Introduction
Modern cyber-physical systems are often comprised of interacting subsystems operated locally by selfish decision-making agents. Ideally, agents operate these systems according to feedback policies that optimise local objectives, while satisfying global requirements and being robust to their rivals’ interference. However, the large-scale and decentralised nature of most applications, alongside the usual lack of coordination between agents, hinders most traditional approaches to policy design. Dynamic game theory provides an alternative framework based on the concept of competitive equilibria (e.g., the Nash equilibrium [1, 2]), which describe locally optimal, yet strategically stable, operating conditions for each non-cooperative agent. A methodology for policy design based on Nash equilibrium seeking thus presents a promising venue for enabling a decentralised control of cyber-physical systems.
In general, solving a non-cooperative game can concern distinct goals: i) For the unbiased observer, to examine the behaviour of rational agents from a local to a global scale; ii) for the game designer, to actuate the self-interested players towards system-wide objectives; and iii) for the player, to determine locally optimal and robust policies in competitive environments. Regardless, computing a Nash equilibrium (NE) solution is a notoriously difficult task [3]. Algorithmic game theory thus emerges as the field concerned with designing procedures to bridge this computational gap [4]. An important class of algorithms are those which place to the players the task of converging towards a non-cooperative equilibrium: These include best-response dynamics (BRD, [5, 6, 7, 8, 9, 10]) and no-regret learning (NRL, [11, 12, 13, 14, 15]). Recently, online feedback optimisation has also been employed for equilibrium seeking problems [16]. In general, these are fixed-point iteration methods centred on players updating their strategies, simultaneously and independently, using only the information available to them. In particular, best-response dynamics stands out as a simple, yet fundamental, model of policy learning for uncoordinated but communicating players. These algorithms have become important tools in economics and engineering, with applications including the design and control of networked systems [17, 18, 19], robotics [20, 21], and resource management [22, 23, 24].
For dynamic games, any solution concept depends qualitatively on a description of the data available to each player; a notion termed information pattern [1]. Under open-loop information patterns, when players only access the initial state of the game, the analysis and solution of NE problems are well-understood. Conversely, games under closed-loop (or feedback) information patterns, when players monitor the state of the game, are still under active research. A feedback Nash equilibrium (FNE) solution is often more attractive than its open-loop counterpart: It renders the players robust to disturbances and decision errors. However, designing routines for computing FNE is demanding for all but simple cases. A notably challenging class of games consists of generalised feedback Nash equilibrium (GFNE, [2]) problems: When players’ policies are required to satisfy coupled constraints (e.g., restrictions on the state of the game). To the best of our knowledge, methods for FNE seeking exist only for linear-quadratic games under restrictive assumptions [25, 26, 27, 28], unconstrained nonlinear and control-affine games [29, 30, 31, 32], and application-specific problems [33, 34]. Recently, [35] expanded these results and presented a systematic (approximate) solution for GFNE problems. Despite remarkable, these efforts are restricted to finite-duration games and thus exclude decision-making processes in which agents are continuously operating the underlying system. Finally, a systematic solution is also lacking for GFNE games described by stochastic dynamics.
In this work, we investigate algorithms for GFNE seeking in (stationary) difference games with linear-stochastic dynamics. Leveraging the System Level Synthesis (SLS, [36]) framework, we propose a best-response approach based on a parametrisation of all stabilising policies for the game. The algorithm consists of players iteratively, and simultaneously, updating the parameters of their individual policies, then announcing these changes through some communication network (Figure 1). Under our approach, each player’s update rule is formulated as the solution to a tractable robust optimisation problem. Moreover, these problems allow for constraints to be imposed directly on the policies’ structure, thus having the ability to encode information patterns at the synthesis level. Finally, this proposed learning dynamics do not depend on the state and actions applied to the system, and thus can be performed simultaneously with the game’s execution. In summary, our contributions are:
  1. (i)
    A system-level best-response dynamics (SLS-BRD) algorithm for GFNE seeking in linear-stochastic dynamic games. Specifically, we design a class of best-response maps which compute an optimal system level parametrisation of a player’s policy in response to its rivals’ choices.
  2. (ii)
    A realisation of the best-response mappings as robust finite-dimensional programs amenable to numerical solutions. Considering these best-responses, the SLS-BRD algorithm is formally an $\varepsilon$-GFNE seeking algorithm for an equilibrium gap $\varepsilon>0$ which can be explicitly established.
  3. (iii)
    For a specific but important class of dynamic games, an analysis of the conditions and rates of convergence of the SLS-BRD to feedback Nash equilibria.
This policy learning algorithm is showcased in simulated experiments on the decentralised control of an unstable network and price management in a competitive oligopolistic market.
Figure 1: SLS-BRD: Control architecture and the learning dynamics.
The paper is organised as follows: Section II overviews the classes of (generalised) static and dynamic games, and the best-response dynamics algorithm. In Section III, we provide a system level parametrisation for linear-stochastic games, then design a best-response dynamics for GFNE seeking. Finally, Section IV illustrates this learning dynamics in simulated examples, and Section V provides some concluding remarks. Towards a concise presentation, only some essential theorems are proved in the main text: The remaining are in the Appendix.
I-A Notation
We use Latin letters to denote vectors (lowercase) and mappings (uppercase), and use boldface to distinguish signals, operators, and their respective spaces. Sets are in calligraphic font; exceptions are the usual $\mathbb{R}$ and $\mathbb{N}$, and the sets of $N\times N$ symmetric ($\mathbb{S}^{N}$), positive semidefinite ($\mathbb{S}^{N}_{+}$), and positive definite matrices ($\mathbb{S}^{N}_{++}$). In particular, sequences are denoted $\bm{x}=(x_{t})_{t\in\mathcal{I}}$ for a countable set $\mathcal{I}\subseteq\mathbb{N}$, or $\bm{x}=(x_{t})_{t=0}^{T}$ if =$\mathcal{I}=\{0,\ldots,T\}$. For $p\in(0,\infty)$, we use the space of $N_{x}$-dimensional vector-sequences $\ell^{N_{x}}_{p}(\mathcal{I})=\{\bm{x}:\|\bm{x}\|_{\ell_{p}}=(\sum_{t\in\mathcal{I}}\|x_{t}\|^{p})^{1/p}<\infty\}$, with $\ell^{N_{x}}_{\infty}$ and ,$\ell^{N_{x}}_{\infty,e}$ the space of all bounded sequences and all sequences, respectively. $\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{Y}})$ is the set of all bounded linear operators $\bm{A}:\bm{\mathcal{X}}\to\bm{\mathcal{Y}}$ and we sometimes denote transformed signals by $\bm{A}\bm{x}=(Ax_{t})_{t\in\mathcal{I}}$. Whenever $(\bm{\mathcal{X}},\bm{\mathcal{Y}})$ are finite-dimensional, we let $M_{A}$ refer to a matrix representation of $\bm{A}$. We use the standard definition of Hardy spaces $\mathcal{H}_{\infty}$ and $\mathcal{RH}_{\infty}$, and write $\frac{1}{z}\mathcal{RH}_{\infty}$ to refer to the set of real-rational strictly proper transfer functions. Finally, some standard signals and operators used in this paper are: The impulse signal $\bm{\delta}=(\delta_{t})_{t\in\mathcal{I}}$, the identity operator $I$ and the identity matrix $I_{N_{x}}$, and the shift operator $\bm{S_{+}}:(x_{0},x_{1},\ldots)\mapsto(0,x_{0},\ldots)$.
We distinguish set-valued mappings from ordinary functions using the notation $F:\mathcal{X}\rightrightarrows\mathcal{Y}$. For any tuple $s=(s^{p})_{p\in\mathcal{P}}\in\mathcal{S}$ we frequently write $s=(s^{p},s^{-p})$ to highlight the element $s^{p}$; this should not be interpreted as a re-ordering. Similarly, if $\mathcal{S}=\prod_{p\in\mathcal{P}}\mathcal{S}^{p}$, we define the product $\mathcal{S}^{-p}=\prod_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\mathcal{S}^{\tilde{p}}$.
II Non-cooperative games and best-response dynamics
A (static) $N_{P}$-player game, denoted by a tuple
$\mathcal{G}\coloneqq(\mathcal{P},\{\mathcal{S}^{p}\}_{p\in\mathcal{P}},\{L^{p}\}_{p\in\mathcal{P}}),$ (1)
defines the problem in which players $p\in\mathcal{P}=\{1,\ldots,N_{P}\}$ each decides on a strategy $s^{p}\in S^{p}(s^{-p})\subseteq\mathcal{S}^{p}$ to minimise an objective function $L^{p}:\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}\to\mathbb{R}$. The strategy spaces $S^{p}$ ($\forall p\in\mathcal{P}$) determine the actions available to the players, with the mappings $S^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ restricting this choice based on the actions from their opponents. As such, both the players’ objectives and feasible strategies depend explicitly on their rivals’ strategies. Finally, the players are assumed to be rational, non-cooperative, and acting simultaneously.
A solution to the game $\mathcal{G}$ is understood as a strategy profile $s=(s^{1},\ldots,s^{N_{P}})\in\mathcal{S}$, $\mathcal{S}=\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}$, having some specified property that makes it agreeable to all players if they act rationally. In non-cooperative settings, a widely accepted solution concept is that of a generalized Nash equilibrium: The game is solved when no player can improve its objective by unilaterally deviating from the agreed strategy profile. Formally,
Definition 1.
A strategy profile $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})\in\mathcal{S}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}$ if
$L^{p}(s^{p^{*}},s^{-p^{*}})\leq\min_{s^{p}\in S^{p}(s^{-p^{*}})}L^{p}(s^{p},s^{-p^{*}}).$ (2)
holds for every player $p\in\mathcal{P}$.
In general, the set of GNEs that solve a game $\mathcal{G}$,
[…]
is not a singleton. As such, players might favour different solutions based on their individual objectives. We thus seek to characterise which solutions are “acceptable”, in the sense that no other equilibrium can improve the objective of all players simultaneously. Consider the partial ordering of $\Omega_{\mathcal{G}}$ defined by
$s^{*}\prec\tilde{s}^{*}\leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ L^{p}(s^{p^{*}})\leq L^{p}(\tilde{s}^{p^{*}}),\quad\forall p\in\mathcal{P},$ (3)
with at least one inequality being strict, for any pair $s^{*},\tilde{s}^{*}\in\Omega_{\mathcal{G}}$. A solution $s^{*}\in\Omega_{\mathcal{G}}$ is then characterised as an admissible GNE if there are no other equilibrium $\tilde{s}^{*}\in\Omega_{\mathcal{G}}$ such that $\tilde{s}^{*}\prec s^{*}$. Note that an admissible GNE can still be favourable to only a subset of players (that is, $\tilde{s}^{*}\nprec s^{*}\nprec\tilde{s}^{*}$), and thus it is possible that $\mathcal{G}$ has no solution that is “fair” to all players. The game might also not admit any GNE (i.e., $\Omega_{\mathcal{G}}=\emptyset$), then being characterised as unsolvable. Hereafter, we ensure that the problems being discussed are well-posed by considering the following conditions on their primitives:
Assumption 1.
For each player $p\in\mathcal{P}$,
  1. a)
    the objective $L^{p}:\mathcal{S}^{1}\times\cdots\mathcal{S}^{N_{P}}\to\mathbb{R}$ is jointly continuous in all of its arguments and convex in the $p$-th argument, $s^{p}\in S^{p}(s^{-p})$, for every $s^{-p}\in\mathcal{S}^{-p}$.
  2. b)
    the mapping $S^{p}:\mathcal{S}^{p}\rightrightarrows\mathcal{S}^{-p}$ takes the form
    $S^{p}(s^{-p})\coloneqq\{s^{p}\in\mathcal{S}^{p}:(s^{p},s^{-p})\in\mathcal{S}_{\mathcal{G}}\},$
    where $\mathcal{S}_{\mathcal{G}}$ is some global constraint set shared by all players. Moreover, $\mathcal{S}^{p}$ and $\mathcal{S}_{\mathcal{G}}$ are both compact convex sets and they satisfy $\mathcal{S}_{\mathcal{G}}\cap(\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}})\neq\emptyset$.
Under Assumption 1, a generalisation of the Kakutani fixed-point theorem ensures that $\mathcal{G}$ has a GNE, that is, $\Omega_{\mathcal{G}}\neq\emptyset$ [37]. In practice, these conditions consider each objective to have a unique optimal value, while imposing the feasible set of strategies to be nonempty and coupled only through a common constraint. Although restrictive, these assumptions still cover a broad class of problems of practical relevance.
We investigate algorithms for solving $\mathcal{G}$. A direct computation of a GNE is equivalent to solving $N_{P}$ optimisation problems simultaneously, as implied by Definition 1. Such an approach would require players to be coordinated and their objectives to be public. Conversely, we consider adaptive procedures in which the players learn their GNE strategies independently. In this direction, consider that $\mathcal{G}$ admits episodic repetitions, and let $s_{k}\coloneqq(s_{k}^{1},\ldots,s_{k}^{N_{P}})$ be the strategy profile taken by players $\mathcal{P}$ at the $k$-th episode. A prototypical learning procedure for equilibrium seeking is outlined in Algorithm 1, where
  • $T^{p}:\mathcal{S}^{p}\times\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ describes how the $p$-th player updates its strategy, based on its opponents’ predicted next actions and given its individual objective;
  • $R^{p}:\mathcal{S}\rightrightarrows\mathcal{S}^{-p}$ describes how the $p$-th player’s predicts its opponents’ strategies for the next episode, based on the strategy profile currently being played.
Input: Game $\mathcal{G}\coloneqq(\mathcal{P},\{\mathcal{S}^{p}\}_{p\in\mathcal{P}},\{L^{p}\}_{p\in\mathcal{P}})$
Output: GNE $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})$
1
2 Initialize $s_{0}=(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k=0$;
3 for $k=1,2,\ldots$ do
4       if $s_{k}\in T(s_{k})$ then return $s_{k}$;
5       for $p\in\mathcal{P}$ do
6             Update $s^{p}_{k+1}\in T^{p}(s^{p}_{k},R^{p}(s_{k})\mid L^{p})$;
7
8
Algorithm 1 Prototypical learning dynamics
The procedure in Algorithm 1 belongs to the class of fixed-point methods: Its termination implies that $s^{*}$ is a fixed-point of both $T^{p}$ and $R^{p}$, that is, $s^{*}\in T^{p}(s^{p^{*}},R^{p}(s^{*}))\subseteq T^{p}(s^{p^{*}},s^{-p^{*}})$. In its general form, the conditions (and convergence rates) for these learning dynamics to approach an equilibrium are difficult to establish. In this work, we build upon a fundamental instance from this class of algorithms: The best-response dynamics (BRD). This routine is overviewed in the following.
Best-response dynamics
Let the mapping $BR^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$,
$BR^{p}(s^{-p})\coloneqq\operatorname*{arg\,min}_{s^{p}\in S^{p}(s^{-p})}L^{p}(s^{p},s^{-p})$ (4)
denote the best-response of $p\in\mathcal{P}$ to other players’ strategies. Collectively, $BR(s)\coloneqq BR^{1}(s^{-1})\times\cdots\times BR^{N_{P}}(s^{-N_{P}})\subseteq\mathcal{S}$ is the joint best-response to any given profile $s\in\mathcal{S}$. Then,
Theorem 1.
A strategy profile $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})\in\mathcal{S}$ is a GNE for $\mathcal{G}$ if and only if $s^{*}\in BR(s^{*})$ or, equivalently,
$s^{p^{*}}\in BR^{p}(s^{-p^{*}}),\quad\forall p\in\mathcal{P}.$ (5)
The task of computing a Nash equilibrium can thus be translated into searching for a fixed-point of the set-valued mapping $BR:\mathcal{S}\rightrightarrows\mathcal{S}$. The set of GNE solutions for $\mathcal{G}$ is the set of all such fixed-points, $\Omega_{\mathcal{G}}\coloneqq\{s^{*}\in\mathcal{S}:s^{*}\in BR(s^{*})\}$. A natural procedure for GNE seeking thus consists of players adapting their strategies towards best-responses to their rivals’ strategies, which they assume will remain constant. Formally,
$T^{p}(s^{p}_{k},R^{p}(s_{k}))\coloneqq(1{-}\eta)s^{p}_{k}+\eta BR^{p}(R^{p}(s_{k})),$ (6)
given $R^{p}(s_{k})=s_{k}^{-p}$ and a learning rate factor of $\eta\in(0,1)$. This learning dynamics, summarised in Algorithm 2, is known as best-response dynamics or fictitious play.
Input: Game $\mathcal{G}\coloneqq(\mathcal{P},\{\mathcal{S}^{p}\}_{p\in\mathcal{P}},\{L^{p}\}_{p\in\mathcal{P}})$
Output: GNE $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})$
1
2 Initialize $s_{0}=(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k=0$;
3 for $k=0,1,2,\ldots$ do
4       if $s_{k}\in BR(s_{k})$ then return $s_{k}$;
5       for $p\in\mathcal{P}$ do
6             Update $s^{p}_{k+1}\in(1{-}\eta)s^{p}_{k}+\eta BR^{p}(s^{-p}_{k})$ ;
7
8
Algorithm 2 Best-Response Dynamics (BRD)
After each episode, the strategy profile gets updated as
$s_{k+1}=T(s_{k})=(1-\eta)s_{k}+\eta BR(s_{k}),$ (7)
given the global update rule $T=(1-\eta)I+\eta BR$. Notably, the mappings $T$ and $BR$ share the same set of fixed-points: The GNEs $\Omega_{\mathcal{G}}$. We can then establish the following result.
Theorem 2.
Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be a non-expansive mapping. Then, the best-response dynamics $s_{k+1}=T(s_{k})$ converge monotonically to a GNE solution $s^{*}\in\Omega_{\mathcal{G}}$, that is,
$\lim_{k\to\infty}\inf_{s^{*}\in\Omega_{\mathcal{G}}}\|T(s_{k})-s^{*}\|=0$ (8)
[…]
This convergence result stems from fixed-point theory, where the BRD algorithm is interpreted as belonging to the class of averaged (or Mann-Krasnosel‘skii) iteration methods [38, 39]. For non-generalised quadratic problems, we show in Section III-B that the best-response mapping $BR$ is non-expansive under specific conditions on $L^{1},\ldots,L^{N_{P}}$ Importantly, if the best-response mapping is a contraction, then the BRD converge geometrically to a GNE solution $s^{*}\in\Omega_{\mathcal{G}}$, which is unique.
Theorem 3.
Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be $L_{BR}$-Lipschitz, $L_{BR}<1$. Then, from any feasible $s_{0}\in\mathcal{S}$, the best-response dynamics $s_{k+1}=T(s_{k})$ converge to the unique GNE $s^{*}\in\Omega_{\mathcal{G}}$ with rate
$\cfrac{\|s_{k}-s^{*}\|}{\|s_{0}-s^{*}\|}\leq\big{(}(1{-}\eta)+\eta L_{BR}\big{)}^{k}$ (9)
[…]
Importantly, these results might not hold in practice whenever the best-response maps, $\{BR^{p}\}_{p\in\mathcal{P}}$, are only approximated (e.g., by solving Eq. (4) numerically). However, such inexact averaged operators are still known to converge under specific assumptions on the accuracy of this approximation [40]. The learning rate $\eta$ also plays a central role in the numerical stability of the BRD algorithm: A careful choice is required to ensure that strategy updates do not escape the feasible set, that is, to ensure that $T(s_{k})\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$ for all $s_{k}\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$. Otherwise, the learning dynamics become undefined and the aforementioned convergence results do not apply. For non-generalised games, $T$ trivially satisfy the constraints for any $\eta\in(0,1)$ and thus a careful design of the learning rate might not be necessary. In such cases, $\eta\to 1$ is the optimal choice if $BR$ is a contraction and the convergence rate in Eq. (9) simplifies to $L_{BR}^{k}$.
Finally, we note that the stopping criteria in Algorithm 2 can be modified to allow for earlier termination. In this case, interrupting the best-response dynamics at some episode $k_{f}>0$ will produce a strategy profile $s_{k_{f}}\in\mathcal{S}$ for which
$L^{p}(s_{k_{f}}^{p},s_{k_{f}}^{-p})\leq\min_{s^{p}\in S^{p}(s_{k_{f}}^{-p})}L^{p}(s^{p},s_{k_{f}}^{-p})+\varepsilon$ (10)
[…]
II-A Infinite-horizon dynamic games
An infinite-horizon dynamic $N_{P}$-player game, denoted by tuple
$\mathcal{G}_{\infty}\coloneqq(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}),$ (11)
is defined by the stochastic linear dynamics
$x_{t+1}=Ax_{t}+\sum_{p\in\mathcal{P}}B^{p}u^{p}_{t}+w_{t},\quad x_{0}\text{ given},$ (12)
describing how the state of the game, $\bm{x}=(x_{t})_{t\in\mathbb{N}}\in\bm{\mathcal{X}}$, evolves in response to the players’ actions $\bm{u}^{p}=(u_{t}^{p})_{t\in\mathbb{N}}\in\bm{\mathcal{U}}^{p}$ ($\forall p\in\mathcal{P}$) and the additive random noise $\bm{w}=(w_{t})_{t\in\mathbb{N}}\in\bm{\mathcal{W}}$. For each realisation $\bm{w}\in\bm{\mathcal{W}}$ and initial $x_{0}$, the state is explicitly expressed as $\bm{x}=\bm{F_{w}}\bm{u}$ via the causal affine operator
$\bm{F_{w}}:\bm{u}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\displaystyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{u}^{p}+\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\Big{)},$ (13)
with $\bm{A}:\bm{x}\mapsto(Ax_{t})_{t\in\mathbb{N}}$ and $\bm{B}^{p}:\bm{u}^{p}\mapsto(B^{p}u^{p}_{t})_{t\in\mathbb{N}}$. Because known, the dependency on $x_{0}$ is omitted to simplify notation. Moreover, we assume $\mathrm{E}w_{t}=0$ and $\mathrm{E}(w_{t}w_{t^{\prime}}^{{\mkern-1.5mu\mathsf{T}}})=\delta_{t-t^{\prime}}\Sigma_{w}$, given a covariance matrix $\Sigma_{w}\in\mathbb{S}_{++}^{N_{x}}$, for every $t,t^{\prime}\in\mathbb{N}$. Finally, the sets $\bm{\mathcal{X}}$, $\bm{\mathcal{U}}^{p}$ ($\forall p\in\mathcal{P}$), and $\bm{\mathcal{W}}$ define all permissible state, action, and noise sequences; they take the form
$\bm{\mathcal{X}}$$\coloneqq\{\bm{x}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):x_{t}\in\mathcal{X},\leavevmode\nobreak\ t\in\mathbb{N}\};$
$\bm{\mathcal{U}}^{p}$$\coloneqq\{\bm{u}^{p}\in\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):u^{p}_{t}\in\mathcal{U}^{p},\leavevmode\nobreak\ t\in\mathbb{N}\};$
$\bm{\mathcal{W}}$$\coloneqq\{\bm{w}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):w_{t}\in\mathcal{W},\leavevmode\nobreak\ t\in\mathbb{N}\},$
given sets $\mathcal{X}\subseteq\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}\subseteq\mathbb{R}^{N_{u}^{p}}$ ($\forall p\in\mathcal{P}$), and $\mathcal{W}\subseteq\mathbb{R}^{N_{x}}$.
In dynamic stochastic games, each player decides on a plan of action $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$ to minimize its objective functional
$J^{p}(\bm{u}^{p},\bm{u}^{-p})\coloneqq\mathrm{E}\left[\sum_{t=0}^{\infty}L^{p}(x_{t},u^{p}_{t},u^{-p}_{t})\right],$ (14)
defined by cost function $L^{p}:\mathcal{X}\times\mathcal{U}^{1}\times\cdots\times\mathcal{U}^{N_{P}}\to\mathbb{R}$. The mappings $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ restrict the permissible strategies for each player based on their rivals’ strategies. Under this formulation, the dynamic game $\mathcal{G}_{\infty}$ is stationary and can be interpreted as a static game defined on the appropriate functional spaces. A plan of action $\bm{u}=(\bm{u}^{1},\ldots,\bm{u}^{N_{P}})\in\bm{\mathcal{U}}$ can then be characterised as a generalised Nash Equilibrium solution to $\mathcal{G}_{\infty}$ when no player can improve its objective by unilaterally deviating from this agreed profile. Formally,
Definition 2.
A strategy profile $\bm{u}^{*}=(\bm{u}^{1^{*}},\ldots,\bm{u}^{N_{P}^{*}})\in\bm{\mathcal{U}}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}_{\infty}$ if
$J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{*}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{*}})$ (15)
holds for every player $p\in\mathcal{P}$.
As before, the set of GNEs that solve $\mathcal{G}_{\infty}$,
[…]
is not necessarily a singleton and can include non-admissible equilibria. The game is considered unsolvable if $\Omega_{\mathcal{G}_{\infty}}=\emptyset$. The following assumptions are taken for this class of games:
Assumption 2.
For each player $p\in\mathcal{P}$ and noise $\bm{w}\in\bm{\mathcal{W}}$,
  1. a)
    the cost functional $J^{p}:\bm{\mathcal{U}}^{1}\times\cdots\bm{\mathcal{U}}^{N_{P}}\to\mathbb{R}$ is jointly continuous in all of its arguments and convex in the $p$-th argument, $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$, for every sequence $\bm{u}^{-p}\in\bm{\mathcal{U}}^{-p}$.
  2. b)
    the mapping $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ takes the form
    $U^{p}(\bm{u}^{-p})\coloneqq\{\bm{u}^{p}\in\bm{\mathcal{U}}^{p}:(\bm{u}^{p},\bm{u}^{-p})\in\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\},$
    where $\bm{F_{w}}^{\dagger}$ is a left-inverse of the operator in Eq. (13) and
    $\bm{\mathcal{U}}_{\mathcal{G}}=\{\bm{u}\in\textstyle\prod_{p\in\mathcal{P}}\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):(u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},\leavevmode\nobreak\ t\in\mathbb{N}\}$
    is a global constraint set shared by all players. The sets $\mathcal{U}^{p}$, $\mathcal{U}_{\mathcal{G}}$, and $\mathcal{X}$ are all non-empty, compact, and convex. Finally, we have that $\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})\neq\emptyset$.
These conditions are analogous to those of Assumption 1: They ensure the existence of GNE solutions to $\mathcal{G}_{\infty}$, that is, $\Omega_{\mathcal{G}_{\infty}}\neq\emptyset$. In this case, the players not only share a global constraint $\bm{\mathcal{U}}_{\mathcal{G}}$, but are also required to ensure that state trajectories lie in a feasible set, $x\in\bm{\mathcal{X}}$, against all possible noise realisations. These constraints often describe operational desiderata and/or limitations concerning the game $\mathcal{G}_{\infty}$. Moreover, note that Assumption 2 implies that the operator $\bm{F_{w}}$ is injective, that is, that distinct control signals cannot produce the same state trajectory. A sufficient condition for this property consists on the matrix $B=[B^{1}\leavevmode\nobreak\ B^{2}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ B^{N_{P}}]$ being full-rank.
Any equilibrium $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ incorporates an open-loop information pattern: Actions $u_{t}^{*}=(u_{t}^{1^{*}},\ldots,u_{t}^{N_{P}^{*}})$ depend explicitly only on initial $x_{0}\in\mathcal{X}$ and stage $t\in\mathbb{N}$. A plan of action with such representation is undesirable, as players become sensible to noise disturbances and decision errors. Conversely, a state-feedback decision policy, $\bm{u}^{*}=K(\bm{x})$ for some $K:\bm{\mathcal{X}}\to\bm{\mathcal{U}}$, often carry desirable robustness properties. In this sense, a feedback (respectively, open-loop) representation of $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ is said to be strongly (weakly) time consistent [1]. In this work, we investigate state-feedback solutions to the game $\mathcal{G}_{\infty}$.
We consider a closed-loop information pattern and assume that each $p$-th player’s actions are represented as
$\bm{u}^{p}\coloneqq\bm{K}^{p}\bm{x},\quad\bm{K}^{p}:\bm{x}\mapsto\bm{\Phi^{p}}*\bm{x},$ (16)
given a causal operator $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\subseteq\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}}^{p})$ defined by its convolution kernel $\bm{\Phi}^{p}=(\Phi^{p}_{n})_{n\in\mathbb{N}}\in\ell_{1}(\mathbb{N})$. The sets $\{\bm{\mathcal{C}}^{p}\}_{p\in\mathcal{P}}$ describe the operators that satisfy some $\mathcal{G}_{\infty}$-related restrictions (e.g., information patterns incurred by communication, actuation, and sensing delays). In this setting, players do not devise their actions explicitly but rather by designing a state-feedback policy profile $\bm{K}\coloneqq(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})\in\bm{\mathcal{C}}$, with $\bm{\mathcal{C}}=\bm{\mathcal{C}}^{1}\times\cdots\times\bm{\mathcal{C}}^{N_{P}}$. The solution concept that naturally arises is that of a generalized feedback Nash equilibrium.
Definition 3.
A policy profile $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$ is a generalized feedback Nash equilibrium (GFNE) for $\mathcal{G}_{\infty}$ if
$J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{*}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{*}}),$ (17)
where $\bm{u}^{*}\in\mathrm{Ker}(I-\bm{K}^{*}\bm{F_{w}})$, holds for every $p\in\mathcal{P}$.
The set of GFNE that solve $\mathcal{G}_{\infty}$ is defined as
[…]
We consider $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ to be admissible only if it renders the game stable, that is, if the closed-loop evolution
$\bm{x}^{*}=\big{(}I-\bm{S_{+}}(\bm{A}-\textstyle\sum_{p\in\mathcal{P}}\bm{B}^{p}\bm{K}^{p^{*}})\big{)}^{-1}\big{(}\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\big{)}$
is bounded ($\bm{x}^{*}\in\ell_{\infty}^{N_{x}}$) for all bounded noise ($\bm{w}\in\ell_{\infty}^{N_{x}}$). A policy satisfying this requirement is said to be stabilising. From Assumption 2, we have $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}\neq\emptyset$ when $\bm{\mathcal{C}}=\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}})$. For a more specific $\bm{\mathcal{C}}\subset\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}})$, establishing the existence (and, especially, uniqueness) of a solution is demanding [41]. In practice, the set $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ can be constructed from open-loop equilibria $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ by i) parametrising the set of all possible trajectories $\{\bm{x}_{\bm{w}}^{*}=\bm{F_{w}}\bm{u}^{*}\}_{\bm{w}\in\bm{\mathcal{W}}}$, then ii) identifying policies $(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$ that satisfy $\{\bm{u}^{p^{*}}=\bm{K}^{p^{*}}\bm{x}_{\bm{w}}^{*}\}_{\bm{w}\in\bm{\mathcal{W}}}$, $p\in\mathcal{P}$. Highlighting this equivalence, we refer to such $\bm{u}^{*}$ as the open-loop realisation of the closed-loop policy $\bm{K}^{*}$, and vice-versa.
Best-response dynamics for GFNE seeking
The mapping
$BR^{p}(\bm{u}^{-p})\coloneqq\operatorname*{arg\,min}_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p})}J^{p}(\bm{u}^{p},\bm{u}^{-p})$ (18)
is the best-response of $p\in\mathcal{P}$ to other players’ plan of action. Under its feedback representation, $\bm{u}^{p}=\bm{K}^{p}\bm{x}\in BR^{p}(\bm{u}^{-p})$ is a solution to the infinite-horizon control problem
$\underset{\bm{u}^{p}\coloneqq\bm{K}^{p}\bm{x}}{\text{minimize}}\quad$ $\mathrm{E}\left[\sum_{t=0}^{\infty}L^{p}(x_{t},u^{p}_{t},u^{-p}_{t})\right]$ (19a)
$\underset{\forall t\in\mathbb{N}}{\text{subject to}}\quad$ $x_{t{+}1}=Ax_{t}+\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}u^{\tilde{p}}_{t}+w_{t},$ (19b)
$x_{t}\in\mathcal{X},\leavevmode\nobreak\ \leavevmode\nobreak\ u^{p}_{t}\in\mathcal{U}^{p},\leavevmode\nobreak\ \leavevmode\nobreak\ (u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},$ (19c)
$\bm{K}^{p}\in\bm{\mathcal{C}}^{p},$ (19d)
$(x_{0}\text{ given}).$ (19e)
While posed in terms of action signals ($\bm{u}^{p}$, $p\in\mathcal{P}$), Problem (19) should be interpreted as the direct search for a best-response policy $\bm{K}^{p}$ against the (fixed) plan of action from other players, $\bm{u}^{-p}\coloneqq(\bm{u}^{\tilde{p}})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. We slightly abuse notation and let $BR^{p}(\bm{K}^{-p})$ be its solutions, when the problem is parametrised by $\bm{u}^{-p}\coloneqq\bm{K}^{-p}\bm{x}=(\bm{K}^{\tilde{p}}\bm{x})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. The mapping $BR:\bm{\mathcal{C}}\rightrightarrows\bm{\mathcal{C}}$, defined by $BR(\bm{K})=BR^{1}(\bm{K}^{-1})\times\cdots\times BR^{N_{P}}(\bm{K}^{-N_{P}})$, is the joint best-response to a strategy profile $\bm{K}$. The GFNEs of $\mathcal{G}_{\infty}$ are thus the fixed-points of this mapping: That is, $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}=\{\bm{K}^{*}\in\bm{\mathcal{C}}:\bm{K}^{*}\in BR(\bm{K}^{*})\}$. Due to constraints ($\mathcal{X},\mathcal{U}^{p},\mathcal{U}_{\mathcal{G}}$) and $\bm{\mathcal{C}}^{p}$, an analytical solution to Problem (19) does not exist. Moreover, because infinite-dimensional, its numerical approximation cannot be obtained.
The BRD method for GFNE seeking is given in Algorithm 3. As $\mathcal{G}_{\infty}$ is dynamic and stationary, the procedure needs not episodic repetitions of the game. Instead, the learning dynamics occurs simultaneously with the game’s execution: Players learn and announce their new policies at stages $t\in\{(k{+}1)\Delta T\}_{k\in\mathbb{N}}$. $\bm{K}_{k}\coloneqq(\bm{K}_{k}^{1},\ldots,\bm{K}_{k}^{N_{P}})$ denotes the strategy profile after $k\in\mathbb{N}$ updates. The period $\Delta T\geq 1$ defines the rate at which policies are updated, reflecting some communication structure (e.g., the time needed for each $p\in\mathcal{P}$ to collect $\{\bm{K}_{k}^{\tilde{p}}\}_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$).
Input: Game $\mathcal{G}_{\infty}\coloneqq(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$
Output: GFNE $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$
1
2 Initialize $\bm{K}_{0}=(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k=0$;
3 for $t=0,1,2,\ldots$ do
       /* Players apply actions $\{u^{p}_{k,t}=K^{p}_{k}x_{k,t}\}_{p\in\mathcal{P}}$/
4       if $\bm{K}_{k}\in BR(\bm{K}_{k})$ then return $\bm{K}_{k}$;
5       if $t=(k{+}1)\Delta T$ then
6             for $p\in\mathcal{P}$ do
7                   Update $\bm{K}^{p}_{k+1}\in(1{-}\eta)\bm{K}^{p}_{k}+\eta BR^{p}(\bm{K}^{-p}_{k})$ ;
8
9            $k\coloneqq k+1$;
10
11
Algorithm 3 BRD for GFNE seeking (BRD-GFNE)
If we verbally execute Algorithm 3, we have the following:
  • The players $p\in\mathcal{P}$ act on $\mathcal{G}_{\infty}$ according to the policies
    $\bm{u}^{p}_{k}=\bm{K}_{k}^{p}\bm{x}_{k},\quad k\in\mathbb{N},$
    where $\bm{u}^{p}_{k}=(u^{p}_{t})_{t\in\mathcal{T}_{k}}$ and $\bm{x}_{k}=(x_{t})_{t\in\mathcal{T}_{k}}$ are the signals restricted to the interval $\mathcal{T}_{k}=[k\Delta T,(k{+}1)\Delta T)$.
  • At $t=(k{+}1)\Delta T$, every $p$-th player updates its policy,
    $\bm{K}_{k+1}^{p}\in(1{-}\eta)\bm{K}^{p}_{k}+\eta BR^{p}(\bm{K}^{-p}_{k}),$
    which is then announced to the other players.
The BRD-GFNE induces an operator $T=(1-\eta)I+\eta BR$ which is equivalent to the update rule of its static counterpart. Thus, it possesses the same properties: The learning dynamics converge if $BR$ is non-expansive and the convergence rate is geometric if $BR$ is also a contraction (Theorems 23). These properties can also be stated in terms of stage indices $t\in\mathbb{N}$ by substituting $k=\lfloor t/\Delta T\rfloor$. As in the static case, a careful choice of the learning rate $\eta\in(0,1)$ is required to ensure that this fixed-point iteration is well-defined. Moreover, the BRD-GFNE can be interrupted at any episode $k_{f}>1$ (producing an $\epsilon$-GFNE, $\epsilon>0$) and, if necessary, a central coordinator assigned to project the resulting strategy profile $\bm{K}_{k_{f}}$ onto the feasible set $\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})$. Finally, we note that such considerations can be disregarded when the game is not generalised, that is, when $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$.
III Best-response dynamics via system level synthesis
In this section, we present an approach for GFNE seeking in (stationary) stochastic dynamic games. Firstly, we introduce the system level parametrisation of the players’ feedback policies ($\bm{K}^{p}$, $p\in\mathcal{P}$) and reformulate their best-response mappings ($BR^{p}$, $p\in\mathcal{P}$) through finite-dimensional robust optimisation problems. Then, a modified BRD-GFNE procedure is proposed and its convergence properties are investigated.
We focus on $N_{P}$-players linear-quadratic stochastic games $\mathcal{G}_{\infty}^{\text{LQ}}=(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$ with dynamics
$x_{t{+}1}=Ax_{t}+\sum_{p\in\mathcal{P}}B^{p}u_{t}^{p}+w_{t},\quad x_{0}\text{ given},$ (20)
and objective functionals
$J^{p}(\bm{u}^{p},\bm{u}^{-p})=\mathrm{E}\Bigg{[}\sum_{t=0}^{\infty}\Big{(}\|C^{p}x_{t}\|_{2}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}u^{\tilde{p}}_{t}\|_{2}^{2}\Big{)}\Bigg{]},$ (21)
defined by matrices $C^{p}\in\mathbb{R}^{N_{z}\times N_{x}}$ and $D^{p\tilde{p}}\in\mathbb{R}^{N_{z}\times N_{u}^{\tilde{p}}}$ with dimension $N_{z}\geq N_{x}+N_{u}$. The following assumptions ensure that GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ are admissible (that is, stabilising):
Assumption 3.
For each player $p\in\mathcal{P}$,
  1. a)
    The pair $(A,B^{p})$ is stabilisable;
  2. b)
    The pair $(C^{p},A)$ is detectable;
  3. c)
    The matrix $D^{pp}$ is full column rank, i.e., $D^{pp^{{\mkern-1.5mu\mathsf{T}}}}D^{pp}\in\mathbb{S}^{N_{u}^{p}}_{++}$. Moreover, $D^{p\tilde{p}^{{\mkern-1.5mu\mathsf{T}}}}C^{p}=0=C^{p^{{\mkern-1.5mu\mathsf{T}}}}D^{p\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$.
Finally, the constraint sets $\mathcal{X}$ and $\mathcal{U}^{p}$ ($\forall p$) are convex polyhedra and contain their respective origins: $0\in\mathcal{X}$ and $0\in\mathcal{U}^{p}$.
The class $\mathcal{G}_{\infty}^{\text{LQ}}$ describe problems in which $N_{P}$ non-cooperative agents have to agree on stationary policies that jointly stabilise a global system, robustly to the noise process, while penalising state- and input-deviations differently. While representative of many practically relevant problems, this choice is not restrictive. Our derivations should follow similarly for any collection of cost functions $\{L^{p}\}_{p\in\mathcal{P}}$ satisfying Assumption 2.
III-A System-level best-response mappings
System level synthesis (SLS, [36]) is a novel methodology for controller design which lifts the search for optimal controllers into the search of optimal closed-loop responses to disturbances. Under this approach, the synthesis of stationary control policies can be equivalently posed as the solution to scalable numerical problems, even when operational constraints on the state and control signals need to be enforced. Additionally, the method allows for constraints to be enforced directly on the structure of the control policies. In this section, we present a system-level parametrisation for the best-response mappings in $\mathcal{G}_{\infty}^{\text{LQ}}$.
We start by assuming a stabilising profile $(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})$, guaranteed by Assumption 3. Each policy is associated with a transfer matrix $\bm{\hat{K}}^{p}\in\mathcal{RH}_{\infty}$, $\bm{\hat{K}}^{p}=\sum_{n=0}^{\infty}\frac{1}{z^{n}}\Phi_{n}^{p}$, which defines the state-feedback $\bm{\hat{u}}^{p}=\bm{\hat{K}}^{p}\bm{\hat{x}}$ in the frequency domain. Considering the linear dynamics Eq. (20),
$z\bm{\hat{x}}$ $=A\bm{\hat{x}}+\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{u}}^{p}+\bm{\hat{w}};$ (22a)
$\bm{\hat{u}}^{p}$ $=\bm{\hat{K}}^{p}\bm{\hat{x}},\quad(\forall p\in\mathcal{P}),$ (22b)
the signals $(\bm{\hat{x}},\bm{\hat{u}}^{1},\ldots,\bm{\hat{u}}^{N_{P}})$ can be expressed in terms of $\bm{\hat{w}}$,
$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}^{1}\\ \vdots\\ \bm{\hat{u}}^{N_{P}}\end{bmatrix}$ $=\begin{bmatrix}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\\ \bm{\hat{K}}^{1}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\\ \vdots\\ \bm{\hat{K}}^{N_{P}}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\end{bmatrix}\bm{\hat{w}};$
$=\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}\bm{\hat{w}},$ (23)
where the introduced transfer matrices ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}$) are denoted as system level responses or closed-loop maps. Under this representation, the following result holds.
Theorem 4 (System level parametrisation, [36]).
Consider the dynamics Eq. (22) under state-feedback $\bm{\hat{u}}^{p}=\bm{\hat{K}}^{p}\bm{\hat{x}}$ ($\forall p\in\mathcal{P}$). The following statements are true:
  1. a)
    The affine space
    $\begin{bmatrix}zI-A&{-}B^{1}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ {-}B^{N_{P}}\end{bmatrix}\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}=I,$ (24)
    with $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}\in\frac{1}{z}\mathcal{RH}_{\infty}$, parametrizes all system responses from $\bm{\hat{w}}$ to $(\bm{\hat{x}},\bm{\hat{u}}^{1},\ldots,\bm{\hat{u}}^{N_{P}})$ achievable by internally stabilising policies $(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$.
  2. b)
    Any response $(\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}})$ satisfying Eq. (24) is achieved by the policies $\bm{\hat{K}}^{p}=\bm{\hat{\Phi}_{u}}^{p}\bm{\hat{\Phi}_{x}}^{-1}$ ($\forall p\in\mathcal{P}$), which are internally stabilising and can be implemented as
    $z\bm{\hat{\xi}}$ $=\bm{\tilde{\Phi}_{x}}\bm{\hat{\xi}}+\bm{\hat{x}};$ (25a)
    $\bm{\hat{u}}^{p}$ $=\bm{\tilde{\Phi}_{u}}^{p}\bm{\hat{\xi}},$ (25b)
    with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$ and $\bm{\tilde{\Phi}_{u}}^{p}=z\bm{\hat{\Phi}_{u}}^{p}$.
Proof.
To simplify the notation, we momentarily define the matrices $B\coloneqq[B^{1}\ B^{2}\ \cdots\ B^{N_{P}}]$, $\bm{\hat{\Phi}_{u}}\coloneqq\big{[}\bm{\hat{\Phi}_{u}}^{1^{{\mkern-1.5mu\mathsf{T}}}}\leavevmode\nobreak\ \bm{\hat{\Phi}_{u}}^{2^{{\mkern-1.5mu\mathsf{T}}}}\cdots\bm{\hat{\Phi}_{u}}^{N_{P}^{{\mkern-1.5mu\mathsf{T}}}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}$, and $\bm{\hat{K}}\coloneqq\big{[}\bm{\hat{K}}^{1^{\mkern-1.5mu\mathsf{T}}}\leavevmode\nobreak\ \bm{\hat{K}}^{2^{\mkern-1.5mu\mathsf{T}}}\cdots\bm{\hat{K}}^{N_{P}^{\mkern-1.5mu\mathsf{T}}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}$. For the first statement, consider internally stabilising policies $(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$ such that $\bm{\hat{u}}^{p}=\bm{\hat{K}}^{p}\bm{\hat{x}}$ ($\forall p\in\mathcal{P}$). Then, from Eq. (23), we have that
$\begin{bmatrix}zI-A&-B\end{bmatrix}\begin{bmatrix}(zI-A-B\bm{\hat{K}})^{-1}\\ \bm{\hat{K}}(zI-A-B\bm{\hat{K}})^{-1}\end{bmatrix}$
$\hskip 15.00002pt=(zI-A)(zI-A-B\bm{\hat{K}})^{-1}{-}B\bm{\hat{K}}(zI-A-B\bm{\hat{K}})^{-1}$
$\hskip 15.00002pt=(zI-A-B\bm{\hat{K}})(zI-A-B\bm{\hat{K}})^{-1}$
$\hskip 15.00002pt=I.$
For the second statement, we first show that $\bm{\hat{K}}$ achieves the desired response, and then that these policies are internally stabilising. Note that Eq. (24) implies that $\bm{\hat{\Phi}_{x}}$ has the leading spectral component $\Phi_{x,1}=I_{N_{x}}$, which is invertible, and thus $\bm{\hat{\Phi}_{x}}^{-1}$ exists. Then, $\bm{\hat{K}}=\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1}$ is well-defined, and we have
$\bm{\hat{x}}$ $=(zI-A-B\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1})^{-1}\bm{\hat{w}}$
$=\bm{\hat{\Phi}_{x}}\left((zI-A)\bm{\hat{\Phi}_{x}}-B\bm{\hat{\Phi}_{u}}\right)^{-1}\bm{\hat{w}}$
$=\bm{\hat{\Phi}_{x}}\bm{\hat{w}},$
due to Eq. (24), and $\bm{\hat{u}}=\bm{\hat{K}}\bm{\hat{x}}=\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1}\bm{\hat{\Phi}_{x}}\bm{\hat{w}}=\bm{\hat{\Phi}_{u}}\bm{\hat{w}}$. Thus, $\bm{\hat{K}}$ achieves the response ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}$), or, equivalently, $(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$ achieve $(\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}})$. To show that this policy is internally stabilising, consider its equivalent representation $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}$ with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$, $\bm{\tilde{\Phi}_{u}}=z\bm{\hat{\Phi}_{u}}$, and $\bm{\tilde{\Phi}_{y}}=I$, as depicted in Figure 2.
Figure 2: Feedback structure for the policy $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}=\bm{\hat{\Phi}_{u}^{p}}\bm{\hat{\Phi}_{x}^{-1}}$, equivalent to the internal representation in Eq. (25).
Introducing external perturbations $\{\delta_{x},\delta_{u},\delta_{\xi}\}\subseteq\ell_{\infty}$ into the game, it suffices to verify that the transfer matrices from $(\bm{\hat{\delta}_{x}},\bm{\hat{\delta}_{u}},\bm{\hat{\delta}_{\xi}})$ to $(\bm{\hat{x}},\bm{\hat{u}},\bm{\hat{\xi}})$ are stable. In this case, we have
$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}\\ \bm{\hat{\xi}}\end{bmatrix}=\begin{bmatrix}\bm{\hat{\Phi}_{x}}&\bm{\hat{\Phi}_{x}}B&\bm{\hat{\Phi}_{x}}(zI-A)\\ \bm{\hat{\Phi}_{u}}&I+\bm{\hat{\Phi}_{u}}B&\bm{\hat{\Phi}_{u}}(zI-A)\\ \frac{1}{z}I&\frac{1}{z}B&\frac{1}{z}(zI-A)\\ \end{bmatrix}\begin{bmatrix}\bm{\hat{\delta}_{x}}\\ \bm{\hat{\delta}_{u}}\\ \bm{\hat{\delta}_{\xi}}\end{bmatrix}.$ (26)
Since $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}\in\frac{1}{z}\mathcal{RH}_{\infty}$, the transfer matrices in Eq. (26) are all stable. We thus conclude that the policy $\bm{\hat{K}}$ is internally stabilising, and so are the policies $(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$. ∎
We refer to the system level responses through their kernels, $\bm{\Phi_{x}}=(\Phi_{x,n})_{n\in\mathbb{N}}\in\ell_{2}(\mathbb{N})$ and $\bm{\Phi_{u}}^{p}=(\Phi^{p}_{u,n})_{n\in\mathbb{N}}\in\ell_{2}(\mathbb{N})$, for all $p\in\mathcal{P}$. Due to strict causality, $\Phi_{x,0}=0$ and $\Phi_{u,0}^{p}=0$. From Theorem 4, the operators $\{\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\}_{p\in\mathcal{P}}$ and the transfer matrices $\{\bm{\hat{K}}^{p}\in\mathcal{RH}_{\infty}\}_{p\in\mathcal{P}}$ are equivalent representations of the feedback policies. Hence, provided there is no confusion, we use exclusively the first notation. In particular, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($p\in\mathcal{P}$) denotes the policy parametrised by $(\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{p})$ and $\bm{K}=(\bm{\Phi_{u}}^{1},\cdots,\bm{\Phi_{u}}^{N_{P}})\bm{\Phi_{x}}^{-1}$ denotes the corresponding profile. A time-domain characterisation of $\bm{K}$ is given in the following.
Corollary 4.1.
A policy $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($\forall p\in\mathcal{P}$) is defined by the kernel $\bm{\Phi}^{p}=\bm{\Phi_{u}}^{p}*\bm{\Phi_{x}}^{-1}$, and implemented as
$\xi_{t}$ $=-\textstyle\sum_{\tau=1}^{t}\Phi_{x,\tau{+}1}\xi_{t-\tau}+x_{t};$ (27a)
$u^{p}_{t}$ $=\phantom{-}\textstyle\sum_{\tau=0}^{t}\Phi^{p}_{u,\tau{+}1}\xi_{t-\tau},$ (27b)
using an auxiliary internal state $\xi=(\xi_{n})_{n\in\mathbb{N}}$ with $\xi_{0}=x_{0}$.
The system level parametrisation enables a methodology for policy synthesis consisting of searching the space of stabilising policies (in)directly through $(\bm{\Phi_{x}},\bm{\Phi_{u}}^{p})$, $p\in\mathcal{P}$. In particular, this parametrisation can be leveraged to reformulate the best-response dynamics maps in $\mathcal{G}_{\infty}^{\text{LQ}}$ as tractable numerical programs. In this direction, consider that players design stabilising policies $\bm{K}=(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})$ by choosing their desired system level responses $\bm{\Phi_{u}}=(\bm{\Phi_{u}}^{1},\ldots,\bm{\Phi_{u}}^{N_{P}})$ simultaneously. From the affine space Eq. (24), the signal $\bm{\Phi_{x}}$, common to all players, satisfies the (deterministic) linear dynamics
$\Phi_{x,n+1}=A\Phi_{x,n}+\sum_{p\in\mathcal{P}}B^{p}\Phi_{u,n}^{p},\quad\Phi_{x,1}=I_{N_{x}},$ (28)
or $\bm{\Phi_{x}}=\bm{F_{\Phi}}\bm{\Phi_{u}}$ given the causal affine operator
$\bm{F_{\Phi}}:\bm{\Phi_{u}}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\displaystyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{\Phi_{u}}^{p}+\bm{\delta}I_{N_{x}}\Big{)},$ (29)
Using Parseval’s Theorem [42], we substitute Eq. (23) into Eq. (21) to redefine the objective functionals of $\mathcal{G}_{\infty}^{\text{LQ}}$ as
$J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})$
$\quad=\mathrm{E}\Big{[}\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}w_{n}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}w_{n}\|_{F}^{2}\Big{)}\Big{]}$
$\quad=\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}.$
The game $\mathcal{G}_{\infty}^{\text{LQ}}$ thus induces a system-level dynamic game,
$\mathcal{G}_{\infty}^{\Phi}\coloneqq(\mathcal{P},\bm{\mathcal{C}}_{x},\{\bm{\mathcal{C}}_{u}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}),$ (30)
defining the problem in which players $p\in\mathcal{P}$ each plans a closed-loop response $\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\subseteq\bm{\mathcal{C}}_{u}^{p}$ to minimise its individual cost functional $J^{p}:\bm{\mathcal{C}}_{u}^{1}\times\cdots\times\bm{\mathcal{C}}_{u}^{N_{P}}\to\mathbb{R}$. Here, the set-valued mappings $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ are defined as
$U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\coloneqq\{\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}:(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})\in\bm{F_{\Phi}}^{\dagger}(\bm{\mathcal{C}}_{x}),\\ \bm{\Phi_{u}}^{p}*\bm{w}\in U^{p}(\bm{\Phi_{u}}^{-p}*\bm{w})\},$
which incorporate the constraints $U^{p}$ from the original $\mathcal{G}_{\infty}^{\text{LQ}}$. The sets $(\bm{\mathcal{C}}_{x},\bm{\mathcal{C}}_{u}^{p})$ are designed to enforce the policy constraints $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}$ directly through the kernels $(\bm{\Phi_{x}},\bm{\Phi_{u}}^{p})$: They are related as $\bm{\mathcal{C}}_{u}^{p}=\{\bm{K}^{p}\bm{\mathcal{C}}_{x}:\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\}$. We refer to a joint response $\bm{\Phi}_{u}=(\bm{\Phi_{u}}^{1},\ldots,\bm{\Phi_{u}}^{N_{P}})\in\bm{\mathcal{C}}_{u}$, $\bm{\mathcal{C}}_{u}=\bm{\mathcal{C}}_{u}^{1}\times\cdots\times\bm{\mathcal{C}}_{u}^{N_{P}}$, as a system-level strategy profile. Finally, the set of (open-loop) system-level GNEs for this game is denoted as $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$.
The best-response mappings for $\mathcal{G}_{\infty}^{\Phi}$ take the form
$BR_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\coloneqq\operatorname*{arg\,min}_{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})}J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p}),$
consisting of the set of closed-loop maps $\bm{\Phi_{u}}^{p}$ which are best-responses to the maps of other players, $\bm{\Phi_{u}}^{-p}=(\bm{\Phi_{u}}^{\tilde{p}})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$. They are solutions to the system level synthesis problem
$\underset{\bm{\Phi_{u}}^{p}}{\text{minimize}}\quad$ $\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}+\Big{\|}\displaystyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\Big{\|}_{F}^{2}\Big{)}$ (31a)
$\underset{\forall n\in\mathbb{N}^{+}}{\text{subject to}}\quad$ $\Phi_{x,{n+1}}=A\Phi_{x,n}+\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (31b)
[…]
$\bm{\Phi_{x}}\in\bm{\mathcal{C}}_{x},\quad\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p},$ (31f)
$\Phi_{x,1}=I_{N_{x}}.$ (31g)
Collectively, the mapping $BR_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$, defined by $BR_{\Phi}(\bm{\Phi_{u}})=BR^{1}_{\Phi}(\bm{\Phi_{u}}^{-1})\times\cdots\times BR^{N_{P}}_{\Phi}(\bm{\Phi_{u}}^{-N_{P}})$, is the joint best-response to a system-level strategy profile $\bm{\Phi_{u}}$. As before, the GNEs of $\mathcal{G}_{\infty}^{\Phi}$ are equivalent to the fixed-points of this map, $\Omega_{\mathcal{G}_{\infty}^{\Phi}}=\{\bm{\Phi_{u}}^{*}\in\bm{\mathcal{C}}_{u}:\bm{\Phi_{u}}^{*}\in BR_{\Phi}(\bm{\Phi_{u}}^{*})\}$. Considering how $\mathcal{G}_{\infty}^{\text{LQ}}$ induces $\mathcal{G}_{\infty}^{\Phi}$, a relationship can be established between the best-responses $BR$ and $BR_{\Phi}$ and, consequently, between their fixed-points, $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$ and $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$. In Section III-B, we formalise this relationship and propose a learning dynamics for GFNE seeking based on the system-level best-response mappings.
Remark 1.
Problem (31) is independent of the matrix $\Sigma_{w}^{1\over 2}$ due to its linear-quadratic structure and the fact that $\Sigma_{w}\in\mathbb{S}^{N_{x}}$. Hence, we remove it from $J^{p}$ ($\forall p$) to simplify notation.
The best-responses $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ are still intractable: i) They are defined by infinite-dimensional problems with no general solution and ii) that require full knowledge of the noise process $(w_{n})_{n\in\mathbb{N}}$ to formulate the constraints $U_{\Phi}^{p}$. In the following, we tackle both issues and provide a class of finite-dimensional robust optimisation programs that approximate Problem (31). We conclude the section by presenting a class of system level constraints which enforce a richer feedback information pattern.
Finite-horizon approximation
The programs in $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ can be made finite-dimensional by restricting the closed-loop maps to the set of finite-impulse responses (FIR),
$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}\in\ell_{2}[0,N]:\Phi_{x,n}\in\mathcal{C}_{x,n},\leavevmode\nobreak\ n{\in}[0,N),\leavevmode\nobreak\ \Phi_{x,N}=0\};$
$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N):\Phi_{u,n}^{p}\in\mathcal{C}^{p}_{u,n},\leavevmode\nobreak\ n{\in}[0,N)\},$
given horizon $N\in[2,\infty)$. We enforce $\bm{\Phi_{x}}\in\bm{\mathcal{C}}_{x}$ and $\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}$ in Problem (31) by adding the terminal constraint $\Phi_{x,N}=0$, and restricting the spectral factors to satisfy $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N-1}$. The constraint sets $\mathcal{C}_{x,n}\subseteq\mathbb{R}^{N_{x}\times N_{x}}$ and $\mathcal{C}_{u,n}^{p}\subseteq\mathbb{R}^{N_{u}\times N_{x}}$ ($\forall p$) have no restriction other than being compact convex sets. The resulting problems are of the form
$\underset{\bm{\Phi_{u}}^{p}=(\Phi_{u,n}^{p})_{n=1}^{N-1}}{\text{minimize}}\quad$ $\sum_{n=1}^{N}\Big{(}\|C^{p}\Phi_{x,n}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\|_{F}^{2}\Big{)}$ (32a)
$\underset{\forall n\in[1,N)}{\text{subject to}}\quad\leavevmode\nobreak\ \leavevmode\nobreak$ $\Phi_{x,{n+1}}=A\Phi_{x,n}+\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (32b)
[…]
$\Phi_{x,n}\in\mathcal{C}_{x,n},\quad\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (32f)
$\Phi_{x,1}=I_{N_{x}},\quad\Phi_{x,N}=0,$ (32g)
which define finite-dimensional convex programs that can be solved numerically. Since $(\bm{\mathcal{C}}_{x},\bm{\mathcal{C}}_{u}^{p})$ are finite-dimensional, the operators $(\bm{F_{\Phi}},\bm{K})$ and the convolutions in Eq. (32e) can be represented by matrix multiplications: Consider any kernel $\bm{\Phi}\in\{\bm{\Phi_{x}}\}\cup\{\bm{\Phi}^{p},\bm{\Phi_{u}}^{p}\}_{p\in\mathcal{P}}$ and some signal $\bm{z}\in\ell_{\infty}$. Then,
$\bm{z}^{\text{out}}=\bm{\Phi}*\bm{z}\leavevmode\nobreak\ \leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ z^{\text{out}}=M_{\Phi}z,$
given block-Toeplitz matrix $M_{\Phi}=[\Phi_{i-j}]_{i\in[1,2N+1],j\in[1,N+1]}$ and $(z,z^{\text{out}})$ being a vector representation of these signals, that is, $z=\texttt{col}(z_{0},\ldots,z_{N})$ and $z^{\text{out}}=\texttt{col}(z^{\text{out}}_{0},\ldots,z^{\text{out}}_{2N})$. Moreover, $z^{\text{out}}_{n}=(\Phi*z)_{n}=[M_{\Phi}]_{n}z$, where $[M_{\Phi}]_{n}$ is the $n$-th block-row of $M_{\Phi}$. Thus, Problem (32) can be reformulated into a program requiring only matrix algebra. Finally, we remark that policies ($\bm{K}^{p}$, $p\in\mathcal{P}$) are implemented either by using Corollary 4.1 or by directly computing kernels $\bm{\Phi}^{p}=([M_{\Phi_{u}^{p}}]_{n}\Phi_{x}^{-1})_{n\in\mathbb{N}}$.
Although realising Problem (31) into a tractable program, the Problem (32) is only feasible when the pair $(A,B^{p})$ is full-state controllable. This is a difficult requirement in multi-agent settings, as often $N_{u}^{p}\ll N_{x}$ for all $p\in\mathcal{P}$, leading to Problem (32) becoming overdetermined. Furthermore, enforcing FIR constraints is known to result in deadbeat policies: Control actions are excessively large in magnitude for small $N<\infty$. Alternatively, we restrict the system level responses to the sets
$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}{\in}\ell_{2}[0,N]:\Phi_{x,n}{\in}\mathcal{C}_{x,n},\leavevmode\nobreak\ n{\in}[0,N),\leavevmode\nobreak\ \|\Phi_{x,N}\|_{F}^{2}\leq\gamma\};$
$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}{\in}\ell_{2}[0,N):\Phi_{u,n}^{p}{\in}\mathcal{C}^{p}_{u,n},\leavevmode\nobreak\ n{\in}[0,N)\},$
with $\|\Phi_{x,N}\|_{F}^{2}=\sum_{i}\sigma_{i}(\Phi_{x,N})^{2}\leq\gamma$ for some factor $\gamma\in(0,1)$ and $\sigma_{i}(\cdot)$ denoting the $i$-th largest singular value of a matrix. These are reffered to as the set of approximately FIR for $N>0$. The $p$-th player best-response map thus takes the form
$\underset{\bm{\Phi_{u}}^{p}=(\Phi_{u,n}^{p})_{n=1}^{N-1}}{\text{minimize}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{n=1}^{N-1}\Big{(}\|C^{p}\Phi_{x,n}\|_{F}^{2}{+}\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\|_{F}^{2}\Big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \hskip 121.00024pt+\|C^{p}\Phi_{x,N}\|_{F}^{2}$ (33a)
$\underset{\forall n\in[1,N)}{\text{subject to}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (33b)
[…]
$\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,n}\in\mathcal{C}_{x,n},\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (33f)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,1}=I_{N_{x}},\leavevmode\nobreak\ \leavevmode\nobreak\ \|\Phi_{x,N}\|_{F}^{2}\leq\gamma.$ (33g)
The solutions to Problem (33) approximate those of the infinite-horizon Problem (31): With respect to $N$, the performance of the former converges to that achieved by the latter [36]. In this case, feasibility only requires $(A,B^{p})$ stabilisable and a sufficiently large horizon $N$ to ensure that $\|\Phi_{x,N}\|_{F}^{2}\leq\gamma$ is achievable for some $\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$. Computationally, this is still a finite-dimensional convex problem which can be solved numerically. Here, we let $\widehat{BR}_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ be the solutions of Problem (33) parametrised by $\bm{\Phi_{u}}^{-p}$. The map $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$, $\widehat{BR}_{\Phi}(\bm{\Phi_{u}})=\widehat{BR}_{\Phi}^{1}(\bm{\Phi_{u}}^{-1})\times\cdots\times\widehat{BR}_{\Phi}^{N_{P}}(\bm{\Phi_{u}}^{-N_{P}}),$ is the joint (approximately)best-response to the system-level profile $\bm{\Phi_{u}}$.
Conversely to $BR_{\Phi}$, the fixed-points of $\widehat{BR}_{\Phi}$ do not coincide with the set of GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, but are rather contained in the set of $\varepsilon$-GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ for some equilibrium gap $\epsilon>0$. This is clear from the fact that the original Problem (31) and the approximation Problem (33) have different optimal values. Under certain conditions, this fact can be shown explicitly.
Theorem 5.
Consider a fixed-point $\bm{\Phi_{u}}^{\varepsilon}\in\widehat{BR}_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$ and assume that $\|\Phi_{x,N}^{\star}\|_{F}^{2}\leq\gamma$ for $\bm{\Phi_{x}}^{\star}=\bm{F_{\Phi}}\bm{\Phi_{u}}^{\star}$ obtained from the original best-response $\bm{\Phi_{u}}^{\star}\in BR_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$. Then, the profile $\bm{\Phi_{u}}^{\varepsilon}=(\bm{\Phi_{u}}^{1^{\varepsilon}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\varepsilon}})$ is an $\varepsilon$-GNE of $\mathcal{G}_{\infty}^{\Phi}$ satisfying
$J^{p}(\bm{\Phi_{u}}^{p^{\varepsilon}},\bm{\Phi_{u}}^{-p^{\varepsilon}})\leq\min_{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{\epsilon}})}J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p^{\varepsilon}})+\varepsilon$ (34)
with $\varepsilon=\max_{p\in\mathcal{P}}\gamma J^{p}(\bm{\Phi_{u}}^{p^{\varepsilon}},\bm{\Phi_{u}}^{-p^{\varepsilon}})$ for every player $p\in\mathcal{P}$.
The equilibrium gap associated with $\bm{\Phi_{u}}^{\varepsilon}\in\widehat{BR}_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$ is thus proportional to the parameter $\gamma$, assuming that the terminal constraint Eq. (33g) also holds for solutions to the original Problem (31). Hereafter, $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ ($\forall p$) are assumed to incorporate the pair ($\bm{\mathcal{C}}_{x}$, $\bm{\mathcal{C}}_{u}^{p}$) defining the set of (soft-constrained) FIR approximations defined above for a $N>0$. We write $U_{\Phi,N}^{p}$ whenever this needs to be made explicit (e.g., to distinguish this choice from a more general $\bm{\mathcal{C}}_{x}$ and $\bm{\mathcal{C}}_{u}^{p}$).
Robust operational constraints
From Assumption 3, the sets $\mathcal{X}$, $\mathcal{U}^{p}$, and $\mathcal{U}_{\mathcal{G}}$, can be expressed by linear inequalities,
$\mathcal{X}$ =$=\{x_{t}\in\mathbb{R}^{N_{x}}:G_{x}x_{t}\leq g_{x}\};$
$\mathcal{U}^{p}$ =$=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\leq g_{u}^{p}\};$
$\mathcal{U}_{\mathcal{G}}$ =$=\{u_{t}\in\textstyle\prod_{\tilde{p}\in\mathcal{P}}\mathbb{R}^{N_{u}^{\tilde{p}}}:\textstyle\sum_{\tilde{p}\in\mathcal{P}}G_{\mathcal{G}}^{\tilde{p}}u^{\tilde{p}}_{t}\leq g_{\mathcal{G}}\},$
given some matrices $G_{x}\in\mathbb{R}^{N_{\mathcal{X}}\times N_{x}}$, $G_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}^{p}}\times N_{u}^{p}}$ and $G_{\mathcal{G}}^{\tilde{p}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}\times N_{u}^{\tilde{p}}}$ ($\forall\tilde{p}\in\mathcal{P}$), and vectors $g_{x}\in\mathbb{R}^{N_{\mathcal{X}}}_{\geq 0}$, $g_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}}^{p}}_{\geq 0}$, and $g_{\mathcal{G}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}}_{\geq 0}$. The map $U^{p}(u^{-p})$ is then equivalent to the actions $\bm{u}^{p}=\bm{\Phi_{u}}^{p}*\bm{w}$ whose associated response $\bm{\Phi_{u}}^{p}$ satisfy
$[G_{x}]_{i}(\Phi_{x}*w)_{n}$ $\leq[g_{x}]_{i},\quad i=1,\ldots,N_{\mathcal{X}};$ (35)
$[G_{u}^{p}]_{j}(\Phi_{u}^{p}*w)_{n}$ $\leq[g_{u}^{p}]_{j},\quad j=1,\ldots,N_{\mathcal{U}^{p}};$ (36)
$\textstyle\sum_{\tilde{p}\in\mathcal{P}}[G_{\mathcal{G}}^{\tilde{p}}]_{l}(\Phi_{u}^{\tilde{p}}*w)_{n}$ $\leq[g_{\mathcal{G}}]_{l},\quad l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}},$ (37)
with $([G_{x}]_{i},[g_{x}]_{i})$, $([G_{u}^{p}]_{j},[g_{u}^{p}]_{j})$, and $([G_{\mathcal{G}}^{\tilde{p}}]_{l},[g_{\mathcal{G}}]_{l})$ denoting, respectively, the $i$-th, $j$-th, and $l$-th rows of the corresponding matrices and vectors. In this work, players are assumed to synthesise policies that satisfy these constraints for any $\bm{w}\in\bm{\mathcal{W}}$. At the expense of conservativeness, we cast Problem (33) as a robust optimization problem by considering the worst-case realisation of the noise. Specifically, we reformulate Eq. (35),
$\sum_{n^{\prime}=0}^{N}\sup_{w\in\mathcal{W}}[G_{x}]_{i}\Phi_{x,n^{\prime}}w\leq[g_{x}]_{i},\quad i=1,\ldots,N_{\mathcal{X}}$ (38)
and similarly for Eqs. (36)–(37), then exploit our knowledge of $\bm{\mathcal{W}}$ to obtain an analytical solution for the supremum.
A common instance of $\mathcal{G}_{\infty}^{\text{LQ}}$ considers the problem in which $(w_{n})_{n\in\mathbb{N}}$ is uniformly bounded. In these problems, the inequality Eq. (38) can be expressed in terms of a dual norm on the appropriate vector space. We highlight two important cases.
  • $\mathcal{W}$ is an ellipsoid centred at zero. Formally, the set is
    $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|Pw_{t}\|_{2}\leq 1\},$
    given a matrix $P\in\mathbb{S}_{++}^{N_{x}}$. This corresponds to noise processes with uniformly bounded energy, as encoded by $\|\cdot\|_{2}$. For such cases, Eqs. (35)–(37) can be enforced by the second-order conic (SOC) constraints
    $\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{x}]_{i},\quad(\forall i);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{u}^{p}]_{j},\quad(\forall j);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{\mathcal{G}}]_{l},\quad(\forall l).$
  • $\mathcal{W}$ is a polyhedron, symmetric around zero. Formally,
    $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|Pw_{t}\|_{\infty}\leq 1\},$
    given a full-rank matrix $P\in\mathbb{R}^{N_{\mathcal{W}}\times N_{x}}$. This corresponds to noise processes with uniformly bounded intensity, as encoded by $\|\cdot\|_{\infty}$. For such cases, Eqs. (35)–(37) can be enforced by the first-order conic constraints
    $\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{x}]_{i},\quad(\forall i);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{u}^{p}]_{j},\quad(\forall j);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{\mathcal{G}}]_{l},\quad(\forall l).$
Remark 2.
If $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}=\mathbb{R}^{N_{u}^{p}}$, and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, these constraints are trivially satisfied for any $w\in\bm{\mathcal{W}}$, and thus can be removed from Problem (31). Conversely, if $\mathcal{W}=\mathbb{R}^{N_{x}}$ when either $\mathcal{X}$, $\mathcal{U}^{p}$, or $\mathcal{U}_{\mathcal{G}}$ is bounded, then no stabilising policy can enforce those constraints for all possible $\bm{w}$.
Structural constraints
The sets $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ ($\forall p\in\mathcal{P}$) are designed to impose some structure directly on the policy $\bm{K}^{p}$ (Corollary 4.1), often in the form of sparsity constraints. An important class of such structural constraints encodes information patterns incurred by the presence of actuation and sensing delays: Let $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) satisfy the restrictions
$\bm{\mathcal{C}}^{p}=\{\bm{K}^{p}\in\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}}^{p}):$ Actions [$\text{Actions }[B^{p}(K^{p}x)_{t}]_{i}\text{ respectively}$ (39)
affect and feedback the component
$[x_{t}]_{i}\text{ with the delays }d_{a},d_{s}>0\},$
and consider the operators $\bm{S}_{x}:\bm{\Phi_{x}}\mapsto(S_{x,n}\odot\Phi_{x,n})_{n\in\mathbb{N}_{+}}$ and $\bm{S}_{u}^{p}:\bm{\Phi_{u}}^{p}\mapsto(S_{u,n}^{p}\odot\Phi_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ ($\forall p$), given the matrices
$(S_{x,n})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}(A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{s}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
$(S_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{s}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
with $\texttt{Sp}(\cdot)$ denoting the sparsity pattern of a matrix [^1] The operator $\texttt{Sp}(A)$ produces a binary matrix such that $[\texttt{Sp}(A)]_{i,j}=1$ if $[A]_{i,j}\neq 0$ and $[\texttt{Sp}(A)]_{i,j}=0$ otherwise.. It can be shown that $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) if its parametrisation satisfy $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n\in\mathbb{N}^{+}}$ and $(\Phi^{p}_{u,n}\in\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ with
$\mathcal{C}_{x,n}$ =$=\{\Phi_{x,n}\in\mathbb{R}^{N_{x}\times N_{x}}:\Phi_{x,n}=S_{x,n}\odot\Phi_{x,n}\};$ (40a)
$\mathcal{C}_{u,n}^{p}$ =$=\{\Phi_{u,n}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}:\Phi_{u,n}^{p}=S_{u,n}^{p}\odot\Phi_{u,n}^{p}\}.$ (40b)
The constraints Eq. (40) enforce that the closed-loop response to the noise obeys an information pattern induced by the dynamics of the game $\mathcal{G}_{\infty}^{\text{LQ}}$. Specifically, $[S_{x,n}]_{i,j}=0$ (respectively, $[S_{u,n}^{p}]_{i,j}=0$) indicates that any disturbance to the $j$-th component of the state, $[x_{t}]_{j}$, does not affect $[x_{t+n}]_{i}$ (does not feedback into action $[u_{t+n}^{p}]_{i}$). Finally, note that $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ are compact convex sets, as they correspond to the kernels of the linear operators ($I-\bm{S}_{x})$ and $(I-\bm{S}_{u}^{p}$).
The ability to impose a desired policy structure using convex constraints is a central feature of the SLS framework. In the context of dynamic games, it allows for describing and, most importantly, solving problems where players have asymmetric information patterns; a major challenge for feedback Nash equilibrium problems [43, 41]. For $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ from Eq. (40), conditions for the existence of GFNE (i.e., $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$), can be stated in terms of the delay parameters $d_{a}$ and $d_{s}$. Unless stated otherwise, the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ (and $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$) are assumed to include these constraints with $d_{a}=d_{s}=1$
III-B System-level best-response dynamics
A learning dynamics based on the system-level best-responses $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ relies on the following central result.
Theorem 6.
A policy profile $\bm{K}^{*}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{*^{-1}}\in\bm{\mathcal{C}}$, is a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ if $\bm{\Phi_{u}}^{*}\in BR_{\Phi}(\bm{\Phi_{u}}^{*})$, or, equivalently,
$\bm{\Phi_{u}}^{p^{*}}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{*}}),\quad\forall p\in\mathcal{P}.$ (41)
Proof.
Consider an arbitrary fixed-point $\bm{\Phi_{u}}^{*}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{*})$. From Theorem 4, we have $\bm{\Phi_{x}}^{*}=\bm{F_{\Phi}}\bm{\Phi_{u}}^{*}$. Now, consider policies $\bm{K}^{p^{*}}=\bm{\Phi_{u}}^{p^{*}}(\bm{\Phi_{x}}^{*})^{-1}$, $p\in\mathcal{P}$. Clearly, $\bm{\Phi_{u}}^{p^{*}}=\bm{K}^{p^{*}}\bm{\Phi_{x}}^{*}$. As a consequence, for any $\bm{w}\in\bm{\mathcal{W}}$,
$\bm{\Phi_{u}}^{p^{*}}\bm{w}=\bm{K}^{p^{*}}\bm{\Phi_{x}}^{*}\bm{w}\leavevmode\nobreak\ \leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{u}^{p^{*}}=\bm{K}^{p^{*}}\bm{x}^{*}.$
and, by definition, $\bm{\Phi_{u}}^{p^{*}}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{*}})$ imply $\bm{u}^{p^{*}}\in U^{p}(\bm{u}^{-p^{*}})$. Thus, $\bm{u}^{*}$ is the open-loop realisation of the policy $\bm{K}^{*}$. Finally, since $J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\cong J^{p}(\bm{\Phi_{u}}^{p^{*}},\bm{\Phi_{u}}^{-p^{*}})$ and $\bm{\Phi_{u}}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, we conclude that no player can obtain an admissible policy that unilaterally improves its cost, that is, $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$. ∎
The relationship between $BR$ and $BR_{\Phi}$ implies that a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be obtained analytically from a GNE of $\mathcal{G}_{\infty}^{\Phi}$. This fact allows us to adapt the BRD-GFNE procedure (Algorithm 3) and propose a procedure for GFNE seeking in constrained infinite-horizon dynamic games based on the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$. This system-level best-response dynamics (SLS-BRD) approach is given in Algorithm 4. We remark on some technical aspects:
  • The pair $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{u,k}}^{p})$ defines the $p$-th player’s parametrisation after $k\in\mathbb{N}$ updates, that is, $\bm{\Phi_{x,k}}^{p}=(\Phi_{x,k,t}^{p})_{t\in\mathbb{N}}$ and $\bm{\Phi_{u,k}}^{p}=(\Phi_{u,k,t}^{p})_{t\in\mathbb{N}}$. The update index $k\in\mathbb{N}$ should not be mistaken for the stage index $t\in\mathbb{N}$.
  • Updating the policy $\bm{K}_{k+1}^{p}$ consists of either employing Corollary 4.1 with maps $(\bm{\Phi_{x,k+1}}^{p},\bm{\Phi_{u,k+1}}^{p})$, or directly computing its kernel $\bm{\Phi_{k+1}}^{p}=([M_{\Phi_{u,k+1}^{p}}]_{n}\Phi_{x,k+1}^{-1})_{n\in\mathbb{N}}$.
  • Responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ are most likely distinct at $k<\infty$, that is, $\bm{\Phi_{x,k}}^{p}\neq\bm{\Phi_{x,k}}^{\tilde{p}}$ for $p\neq\tilde{p}$. Consequently, the system level parametrisation (Eq. 24) might not hold for the profile $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ and this could lead to stability issues. However, this policy still satisfies a robust variant of Theorem 4 when the distances $\|\bm{\Phi_{x,k}}^{p}-\bm{F_{\Phi}}\bm{\Phi_{u,k}}\|$ ($\forall p\in\mathcal{P}$) are sufficiently small [36].
Input: Game $\mathcal{G}_{\infty}^{\text{LQ}}\coloneqq(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}^{p}}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$
Output: GFNE $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$
1
2 Initialize $\bm{K}_{0}=(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k=0$;
3 for $t=0,1,2,\ldots$ do
       /* Players apply actions $\{u^{p}_{k,t}=K^{p}_{k}x_{k,t}\}_{p\in\mathcal{P}}$/
4       if $\bm{\Phi_{u,k}}\in BR_{\Phi}(\bm{\Phi_{u,k}})$ then return $\bm{K}_{k}$;
5       if $t=(k{+}1)\Delta T$ then
6             for $p\in\mathcal{P}$ do
[…]
8            $k\coloneqq k+1$;
9
10
Algorithm 4 System-level BRD (SLS-BRD)
The SLS-BRD induces an operator $T_{\Phi}=(1-\eta)I+\eta BR_{\Phi}$, which defines the global update $\bm{\Phi_{u,k+1}}$ from $\bm{\Phi_{u,k}}$. As before, $T_{\Phi}$ and $BR_{\Phi}$ share the same fixed-points: The GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$. From Theorem 6, convergence to a response $\bm{\Phi_{u}}^{*}\in T_{\Phi}(\bm{\Phi_{u}}^{*})$ then implies convergence to a policy $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$. Hence, this learning dynamics is a formal procedure for GFNE seeking.
In this general form, Algorithm 4 is still unpractical due to $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ being intractable. The SLS-BRD can be adapted to consider instead $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$: The players’ updates become
$\bm{\Phi_{u,k+1}}^{p}\coloneqq(1{-}\eta)\bm{\Phi_{u,k}}^{p}+\eta\widehat{BR}^{p}_{\Phi}(\bm{\Phi_{u,k}}^{-p}).$
The global update rule induced by this modified algorithm is $\widehat{T}_{\Phi}=(1-\eta)I+\eta\widehat{BR}_{\Phi}$. In this case, the fixed-points of $\widehat{T}_{\Phi}$ coincide with those of $\widehat{BR}_{\Phi}$. As such, this (approximately)best-response dynamics is a procedure for $\epsilon$-GFNE seeking, $\varepsilon>0$.
Convergence of the SLS-BRD
The convergence of Algorithm 4 depends on $BR_{\Phi}$ (or $\widehat{BR}_{\Phi}$, for its tractable version) being at least non-expansive (Theorems 23). Formally,
Theorem 7.
Let the map $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$ be $L_{\widehat{BR}_{\Phi}}$-Lipschitz, with $L_{\widehat{BR}_{\Phi}}<1$. Then, the SLS-BRD $\bm{\Phi_{u,k+1}}=\widehat{T}_{\Phi}(\bm{\Phi_{u,k}})$ converge to the unique $\varepsilon$-GNE $\bm{\Phi_{u}}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ with rate
$\cfrac{\|\bm{\Phi_{u,k}}-\bm{\Phi_{u}}^{*}\|_{\ell_{2}}}{\|\bm{\Phi_{u,0}}-\bm{\Phi_{u}}^{*}\|_{\ell_{2}}}\leq\big{(}(1{-}\eta)-\eta L_{\widehat{BR}_{\Phi}}\big{)}^{k}$ (42)
from any feasible initial $\bm{\Phi_{u,0}}$.
In general, determining a Lipschitz constant for such mappings is challenging. However, for linear-quadratic games $\mathcal{G}_{\infty}^{\text{LQ}}$ where $\mathcal{W}$ is a polyhedron, best-responses are piecewise-affine operators and their Lipschitz properties are straightforward. In the following, we use this fact to establish some conditions for convergence of the SLS-BRD for a specific class of games.
Consider the (approximately)best-response maps $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ and assume $N$ sufficiently large to ensure that $\|\Phi_{x,N}\|_{F}^{2}<\gamma$ is strictly satisfied. For notational convenience, let us introduce the operators $\bm{F_{\Phi}}^{p}=(I-\bm{S_{+}}\bm{A})^{-1}\bm{S_{+}}\bm{B}^{p}$ ($\forall p\in\mathcal{P}$) and signal $\bm{F_{\Phi}}^{0}=(I-\bm{S_{+}}\bm{A})^{-1}\bm{\delta}I_{N_{x}}$, and the objective-related mappings $\bm{C}^{p}:\bm{\Phi_{x}}\mapsto(C^{p}\Phi_{x,n})_{n\in\mathbb{N}}$ and $\bm{D}^{p\tilde{p}}:\bm{\Phi_{u}}^{\tilde{p}}\mapsto(D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}})_{n\in\mathbb{N}}$, with $\bm{D}^{p0}=0$. Moreover, for all $p,\tilde{p}\in\mathcal{P}$, define the operators
$\bm{H}^{p\tilde{p}}=(\bm{C}^{p}\bm{F_{\Phi}}^{p}+\bm{D}^{pp})^{{\mkern-1.5mu\mathsf{T}}}(\bm{C}^{p}\bm{F_{\Phi}}^{\tilde{p}}+\bm{D}^{p\tilde{p}}),$ (43)
and $\bm{H}^{p,-p}=(\bm{H}^{p,\tilde{p}})_{\tilde{p}\in\mathcal{P}}$. Because $\{\bm{\Phi}_{u}^{p}\}_{p\in\mathcal{P}}$ are FIR, all the elements defined above have equivalent matrix representations. Using this notation, Problem (33) can be reformulated as
$\underset{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})}{\text{minimize}}\leavevmode\nobreak$ $\mathrm{Tr}\big{[}{\bm{\Phi_{u}}^{p}}^{{\mkern-1.5mu\mathsf{T}}}\bm{H}^{pp}\bm{\Phi_{u}}^{p}$
$\hskip 40.00006pt+2(\textstyle\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\bm{H}^{p\tilde{p}}\bm{\Phi_{u}}^{\tilde{p}}+\bm{H}^{p0})^{{\mkern-1.5mu\mathsf{T}}}\bm{\Phi_{u}}^{p}\big{]}$ (44)
The maps $BR^{p}(\bm{\Phi_{u}}^{-p})$, $p\in\mathcal{P}$, thus correspond to the solution of quadratic programs with convex constraints $U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$, for $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$. In the case of $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$, and without structural constraints, that is, $\bm{S}_{x}=I$ and $\bm{S}_{u}^{p}=I$,
$U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$ =$:\displaystyle=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N)$ : (45)
$\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},\leavevmode\nobreak\ \Phi_{x,1}=I_{N_{x}},$
$\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{x}]_{i},$
$\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{u}^{p}]_{j},$
$\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{\mathcal{G}}]_{l}\}.$
Using standard techniques from optimisation, the constraints Eq. (45) can be incorporated into Problem (44) as linear inequalities. The solutions for these problems, that is, the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, are thus piecewise affine in $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$ [44]. Consequently, also $\widehat{BR}_{\Phi}$ must be piecewise affine. A local Lipschitz constant can then be derived for each region of $\bm{\mathcal{C}}_{u}^{-p}$ that leads to a subset of the operational constraints being active. For non-generalised games (i.e., when $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$), the structure of this affine mapping can be exploited to derive a global Lipschitz constant for $\widehat{BR}_{\Phi}$.
Theorem 8.
Consider $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$. Then, The map $\widehat{BR}_{\Phi}$ is $L_{\widehat{BR}_{\Phi}}$-Lipschitz with
$L_{\widehat{BR}_{\Phi}}=\sqrt{\textstyle\sum_{p\in\mathcal{P}}(L_{\widehat{BR}_{\Phi}}^{p})^{2}},$ (46)
given the player-specific constants
$L_{\widehat{BR}_{\Phi}}^{p}=\big{(}1+\kappa(\bm{H}^{pp})\big{)}\|(\bm{H}^{pp})^{\dagger}\bm{H}^{p,-p}\|_{2\to 2},$ (47)
with condition number $\kappa(\bm{H}^{pp})=\|(\bm{H}^{pp})^{\dagger}\|_{2\to 2}\|\bm{H}^{pp}\|_{2\to 2}$.
The proof of Theorem 8 is extensive: The reader is referred to Appendix -C for the full details. Importantly, this Lipschitz constant is not tight and thus $L_{\widehat{BR}_{\Phi}}<1$ using Eq. (46) is only a sufficient (but not necessary) condition for Theorem 7 to hold in practice. Regardless, it highlights some intuitive, but non-trivial, facts about the convergence of the SLS-BRD:
  • The condition $L_{\widehat{BR}_{\Phi}}<1$ is implies that the block-operator $\bm{H}=[H^{p\tilde{p}}]_{p,\tilde{p}\in\mathcal{P}}$ is diagonally dominant. This highlights the relationship between the SLS-BRD and the class of Jacobi iterative methods, where diagonal dominance of the linear system being solved is a strict requirement.
  • The convergence rate of the SLS-BRD is inversely proportional to $\|D^{pp}\|_{2}^{2}$ ($\forall p$) through $\bm{H}^{pp}$. As such, faster convergence is expected for games in which players apply strong penalties to their own actions.
  • The convergence rate of the SLS-BRD depends on the number of players: Let $L_{\widehat{BR}_{\Phi}}^{p}=\sqrt{\alpha+\beta^{p}}$ for all $p\in\mathcal{P}$ given some $\alpha,\beta^{p}>0$. Then, $L_{\widehat{BR}_{\Phi}}^{2}=(\alpha N_{p})^{2}\sum_{p}\beta^{p}$. In large-scale games, players might need to become more conservative in order to ensure that $\widehat{BR}$ is a contraction.
  • The convergence rate of the SLS-BRD is dominated by the slowest player: Whenever there exists a $p\in\mathcal{P}$ such that $L_{\widehat{BR}_{\Phi}}^{p}\gg L_{\widehat{BR}_{\Phi}}^{\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$, then $L_{\widehat{BR}_{\Phi}}\approx L_{\widehat{BR}_{\Phi}}^{p}$.
  • The norms $\|\bm{H}^{pp}\|_{2\to 2}=\|\bm{C}^{p}\bm{F_{\Phi}}^{p}+\bm{D}^{pp}\|_{2\to 2}^{2}$ increase with the FIR horizon $N$, specially for unstable games. In some cases, players might need to proportionally decrease the state-penalties $C^{p}$ to ensure (or improve) convergence.
At the cost of interpretability, similar Lipschitz constants as in Eqs. (46)–(47) can be obtained also when $(\mathcal{X},\mathcal{U}_{\mathcal{G}},\mathcal{U}^{p})$ and $P$ are general, and when structural constraints are present. Moreover, we remark that Theorem 8 is obtained under the assumption that $\|\Phi_{x,N}\|_{F}^{2}<\gamma$ holds strictly. The best-response maps $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ when this constraint is active, or when $\mathcal{W}$ is an ellipsoid, become quadratically-constrained quadratic programs and their Lipschitz properties are less intuitive.
IV Examples
In the following, we demonstrate the SLS-BRD algorithm in two exemplary problems: Decentralised control of an unstable network and management of a competitive market. In both problems, we set the initial profile $\bm{K}_{0}=(\bm{\Phi_{u,0}}^{1},\ldots,\bm{\Phi_{u,0}}^{N_{P}})\bm{\Phi_{x,0}}^{-1}$ by projecting the zero-response $\bm{\hat{\Phi}_{u,0}}=\bm{0}$ into the feasible set. [^2] That is, $\bm{\Phi_{u,0}}=\operatorname*{arg\,min}_{\bm{\Phi_{u}}\in\bm{\mathcal{U}}_{\Phi}}\|\bm{\Phi_{u}}-\bm{\hat{\Phi}_{u,0}}\|_{\ell_{2}}^{2}$ given the feasible set $\bm{\mathcal{U}}_{\Phi}=\{\bm{\Phi_{u}}\in\ell_{2}[0,N):\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p}),\leavevmode\nobreak\ p\in\mathcal{P}\}$. A SLS-BRD routine is then simulated by having players update their policies through best-responses to the parametrisation of their rivals’ policies, assumed to be available. Due to numerical limitations, we interrupt the updates whenever the condition $\|\bm{\Phi}_{u,k}^{p}-\bm{\Phi}_{u,k-1}^{p}\|_{\ell_{2}}/\|\bm{\Phi}_{u,k}^{p}\|_{\ell_{2}}\leq 10^{-8}$ ($\forall p\in\mathcal{P}$) is satisfied.
IV-A Stabilisation of a bidirectional chain network
Consider a game $\mathcal{G}_{\infty}^{\text{LQ}}=\{\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}\}$ with players $\mathcal{P}=\{1,2,3\}$ operating a chain network of $N_{x}=14$ single-state nodes whose dynamics are described by
$A=\begin{bmatrix}1&0.2&&\\ -0.2&\ddots&\ddots&\\ &\ddots&\ddots&0.2\\ &&-0.2&1\end{bmatrix};\leavevmode\nobreak\ \left\{B^{p}{=}\begin{bmatrix}\bm{0}_{6(p{-}1)\times 2}\\[6.45831pt] I_{2}\\[2.15277pt] \bm{0}_{6(3{-}p)\times 2}\end{bmatrix}\right\}_{p\in\mathcal{P}},$ (48)
where $A\in\mathbb{R}^{N_{x}\times N_{x}}$ and $B^{p}\in\mathbb{R}^{N_{x}\times N_{u}^{p}}$ with $N_{u}^{p}=2$ ($\forall p$). This game has unstable dynamics, since $\rho(A)=1.073>1$, however it is stabilisable for each $(A,B^{p})$. Moreover, the game is subjected to a noise process described by $w_{t}\sim\mathrm{Uniform}(\mathcal{W})$, $t\in\mathbb{N}$, defined over $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$. In this problem, players are interested in stabilising the game $\mathcal{G}_{\infty}^{\text{LQ}}$, while minimising their individual objective functionals,
$J^{p}(u^{p},u^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
[…]
$\mathcal{X}$ $=\mathbb{R}^{N_{x}};$
[…]
$\mathcal{U}_{\mathcal{G}}$ $=\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}},$
enforcing $-10\leq[u_{t}^{p}]_{1}+[u_{t}^{p}]_{2}\leq 10$ for every $t\in\mathbb{N}$ and $p\in\mathcal{P}$. We assume that players design their state-feedback policies, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$, considering a FIR horizon of $N=50$ and the constraints $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N}$,
$\mathcal{C}_{x,n}$ =$=\{\Phi_{x,n}\in\mathbb{R}^{N_{x}\times N_{x}}:\Phi_{x,n}=\texttt{Sp}(A^{n-1})\odot\Phi_{x,n}\};$
$\mathcal{C}_{u,n}^{p}$ =$=\{\Phi_{u,n}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}:\Phi_{u,n}^{p}=\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{n-1})\odot\Phi_{u,n}^{p}\}.$
Under this setup, $\mathcal{G}_{\infty}^{\text{LQ}}$ belongs to the class of dynamic potential games (DPG, [45]) and the (unique) GFNE can be obtained in advance by solving a centralised optimisation problem.
In this experiment, we simulate a GFNE seeking procedure for each different value $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$. The game $\mathcal{G}_{\infty}^{\text{LQ}}$ is executed alongside the updating of players’ policies according to some learning dynamics. The players seek $\varepsilon$-GFNE policies by adhering to the SLS-BRD routine (Algorithm 4) using their (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, with $\gamma=0.95$. The policies are updated simultaneously every $\Delta T=1$ stage with a learning rate of $\eta=1/2$.
The convergence of the SLS-BRD routine to the fixed-point $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 3. The results demonstrate that the players’ policies are sufficiently close to the $\epsilon$-GFNE profile after $260$, $420$, and $510$ iterations for each respective weighting configuration. In each case, the (soft) FIR constraints are satisfied strictly after the initial profile (that is, $\|\Phi_{x,k,N}^{p}\|_{F}^{2}<0.95$, $k>2$, $p\in\mathcal{P}$). The Lipschitz properties of the global update rule $T_{\Phi}$ seem to follow the intuition provided in Section III-B: The convergence rate improves when players apply stronger penalties to their actions ($\beta=(10,40,10)$) when compared to that obtained by weaker control penalties ($\beta=(0.4,1.6,0.4)$). However, we stress that the Lipschitz constants from Theorem 8 do not consider structural constraints ($\bm{S}_{x}$ and $\bm{S}_{u}^{p}$, $\forall p\in\mathcal{P}$) and cannot be applied to this experiment. Regardless, the convergence rates are shown to be geometric, indicating that the best-response maps $\widehat{BR}_{\Phi}$ in these scenarios are contractive. If not interrupted, and disregarding numerical limitations, the SLS-BRD should continue to approach the fixed-point $\bm{K}^{*}$ at this rate.
Figure 3: $N_{P}$-Chain game: Convergence of the SLS-BRD routine.
Figure 5: $N_{P}$-Chain game, $t\in(550,600]$: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for each execution of $\mathcal{G}_{\infty}^{\text{LQ}}$ with $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$. The vertical axis represents each node in the chain-networked system.
In Figure 4 (top), we show the relative distance between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ from each player. The updates are of similar magnitude for all players $p\in\mathcal{P}$, in each scenario, except for player $p=3$ which shows a slightly faster convergence. In general, these local changes become numerically negligible at a faster rate than the global convergence in Figure 3. Specifically, when the relative distance between updates approaches the aforementioned threshold of $10^{-8}$, the corresponding policy profile has become closer to the $\epsilon$-GFNE by a factor of $10^{-6}$. Finally, we consider the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between the responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ obtained from the system-level parametrisation associated with $\bm{\Phi_{u,k}}$ (Theorem 4). As shown in Figure 4 (bottom), these distances decrease at a similar rate and are relatively small since the initial stages of the game. The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are expected to stabilise the system during this learning dynamics.
Figure 4: $N_{P}$-Chain game: Relative distance (in $\log$-scale) between updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k+1}}^{p})$, top panels, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom panels.
The evolution of the game given the actions by each player is displayed in Figure 5 for the period $t\in(550,600]$, when the policy updates have been interrupted. The results demonstrate that the policies obtained by the SLS-BRD routine are capable of jointly stabilising the networked system, robustly against the random noise. These policies are also shown to satisfy the operational constraints: Each player’s actions become (roughly) symmetrical and satisfy $|[u_{t}]_{1}+[u_{t}]_{2}|\leq 5$, $t\in\mathbb{N}$, in all scenarios. The enforcement of this strategy highlights the conservativeness resulting from the robust operational constraints. Finally, note that the state- and control-trajectories from operating the system through these policies are similar for all $\beta$ configurations; however, the policies with $\beta=(2,8,2)$ and $\beta=(0.4,1.6,0.4)$ achieve a slightly better noise rejection at the expense of more aggressive control actions. In this experiment, the choice $\beta=(10,40,10)$ seems to be preferable as it converges faster while still achieving satisfactory performance.
IV-B Price competition in oligopolistic markets
Consider a game $\mathcal{G}_{\infty}^{\text{LQ}}{=}\{\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}\}$ consisting of a set of companies $\mathcal{P}=\{1,2,3,4\}$ participating in a single-product market. Assume that these companies have equivalent production capacities and are able to satisfy the demand for their products. The product offered by company $p\in\mathcal{P}$ has a daily local demand $\bm{d^{p}}=(d^{p}(t))_{t\in\mathbb{R}_{\geq 0}}$ which evolves according to the continuous-time dynamics
$\tau\dot{d}^{p}(t)=\underbrace{d_{\text{base}}^{p}(t)-\textstyle\sum_{\tilde{p}\in\mathcal{P}}\widetilde{B}^{p\tilde{p}}u^{\tilde{p}}(t)}_{\text{linear price-demand curve}}-d^{p}(t),$
where $\bm{u}^{\tilde{p}}=(u^{\tilde{p}}(t))_{t\in\mathbb{R}_{\geq 0}}$ are price changes around the value at which $\tilde{p}\in\mathcal{P}$ sell its products and $\bm{d_{\text{base}}}^{p}=(d_{\text{base}}^{p}(t))_{t\in\mathbb{R}_{\geq 0}}$ is some fluctuating baseline demand. Specifically, the baseline demands are of the form $d_{\text{base}}^{p}(t)=\bar{d}_{base}+v^{p}(t)$, for all $p\in\mathcal{P}$, given fixed $\bar{d}_{base}=10$ and noise process $\bm{v^{p}}=(v^{p}(t))_{t\in\mathbb{R}_{\geq 0}}$. The market parameters $\tau\in\mathbb{R}_{\geq 0}$ and $\widetilde{B}^{p\tilde{p}}\in\mathbb{R}_{\geq 0}$ ($\forall p,\tilde{p}\in\mathcal{P}$) describe how local demands respond to price changes: We set $\tau=1.2$ and sample $B^{p\tilde{p}}\sim\mathrm{Uniform}(0.5,1.5)$ for all $p,\tilde{p}\in\mathcal{P}$. In this problem, companies aim at devising pricing policies to stabilise their demands around $\bar{d}_{base}$, which provides a stable profit margin, while satisfying a price-cap regulation enforcing
$-\bar{u}_{\text{avg}}\leq\textstyle\frac{1}{N_{P}}\textstyle\sum_{p\in\mathcal{P}}u^{p}(t)\leq\bar{u}_{\text{avg}},\quad\bar{u}_{\text{avg}}=0.5.$
Define $\bm{x}=(\bm{x}^{1},\ldots,\bm{x}^{N_{P}})$, with $\bm{x}^{p}=(d^{p}(t)-\bar{d}_{base})_{t\in\mathbb{R}_{\geq 0}}$, and $\widetilde{B}^{p}=[\widetilde{B}^{p1}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \widetilde{B}^{pN_{P}}]$, for all $p\in\mathcal{P}$. Considering a zero-order hold of inputs with period $\Delta t=1/4$ [days], the game $\mathcal{G}_{\infty}^{\text{LQ}}$ can be described by the discrete-time dynamics [^3] With a slight abuse of notation, we use $t$ for both the continuous-time ($x(t)$, $t\in\mathbb{R}_{\geq 0}$) and discrete-time ($x_{t}$, $t\in\mathbb{N}$) signals.
$x_{t+1}=Ax_{t}+\sum_{p\in\mathcal{P}}B^{p}u^{p}_{t}+w_{t}$
with $A=\exp(-\tau\Delta t)I_{N_{P}}$ and $\{B^{p}=-\frac{1}{\tau}(A-I_{N_{x}})\widetilde{B}^{p}\}_{p\in\mathcal{P}}$, and the noise process $\bm{w}=\big{(}-\frac{1}{\tau}(A-I_{N_{x}})v_{t}\big{)}_{t\in\mathbb{N}}$ [46]. The companies assume that the baseline demand fluctuations satisfy $w_{t}\in\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$ for all $t\in\mathbb{N}$. Under this representation, each player’s objective is formulated as solving for a policy which minimises the functional
$J^{p}(u^{p},u^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\alpha^{p}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
given weights $\alpha^{p},\beta^{p}\in\mathbb{R}_{\geq 0}$. We sample $\alpha^{p}\sim\mathrm{Uniform}(5,15)$ and $\beta^{p}\sim\mathrm{Uniform}(0.3,0.6)$ for each $p\in\mathcal{P}$. The operational constraints, $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$, are defined by the constraint sets
$\mathcal{X}$ $=\mathbb{R}^{N_{x}};$
$\mathcal{U}^{p}$ $=\mathbb{R}^{N_{u}^{p}};$
$\mathcal{U}_{\mathcal{G}}$ =$=\big{\{}u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:\textstyle\frac{1}{N_{P}\cdot\bar{u}_{\text{avg}}}\big{[}\mathbf{1}_{N_{u}}\leavevmode\nobreak\ {-}\mathbf{1}_{N_{u}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}u_{t}\leq[1\leavevmode\nobreak\ 1]^{{\mkern-1.5mu\mathsf{T}}}\big{\}},$
We consider that players design their state-feedback policies, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$, with a FIR horizon of $N=16$ and no structural constraints, that is, $\bm{S}_{x}=I$ and $\bm{S}_{u}^{p}=I$ ($\forall p\in\mathcal{P}$). In practice, this implies that companies have perfect information of any demand $\bm{x}^{p}$ and price $\bm{u}^{p}$ changes in the market.
As in Section IV-A, we simulate an instance of the game $\mathcal{G}_{\infty}^{\text{LQ}}$ alongside the SLS-BRD routine (Algorithm 4) with players using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, given $\gamma=0.95$. Policies are updated simultaneously every $\Delta T=1$ stage with learning rate $\eta=1/4$. Under the above setup, the game $\mathcal{G}_{\infty}^{\text{LQ}}$ is not a potential game and thus a GFNE cannot be easily computed in advance. Finally, we remark that solutions to this problem are neither unique nor necessarily admissible: Different initial profiles result in convergence to different fixed-points, which might be unfair to a subset of players.
The convergence of the SLS-BRD routine to a fixed-point $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 6. Since the exact fixed-point to which the routine will converge is not known for this problem, we let $\bm{\Phi_{u}}^{*}\approx\bm{\Phi_{u,k_{f}}}$ for $k_{f}=828$ (when the policy updates are interrupted) and analyse the convergence with respect to this point. In this case, the iterates approach the fixed-point mostly at a geometric rate with the policies requiring roughly $800$ updates (or $200$ days in-game) before changes become numerically negligible. The rate of convergence is also observed to slow around $k=15$, when the generalised constraints ($\mathcal{U}_{\mathcal{G}}$) become active. We stress that the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ cannot be (globally) contractive in this case, as the set of fixed points $\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ is not a singleton.
Figure 6: Market game: Convergence of the SLS-BRD routine.
In Figure 7 (top), the relative distances between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ are displayed. As in the previous example, the updates are of similar magnitude for all players $p\in\mathcal{P}$ and they become numerically negligible at a faster rate than the global convergence in Figure 6. The convergence of the distance between updates is shown to slow considerably, especially for $p\in\{1,4\}$, around $k=50$. Thereafter, these relative distances continue to decrease at a steady rate. In Figure 7 (bottom), the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ are shown. As before, these distances decrease at a similar rate and are relatively small since the initial stages of the game. The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate the stabilising profile $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are expected to be stabilising for all iterations.
Figure 7: Market game: Relative distance between updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k+1}}^{p})$, top panels, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom panels, for all $p\in\mathcal{P}$.
The evolution of the game given each player’s actions is displayed Figure 5 for the in-game period $t\in(260,400]$ days, after the policy updates have been interrupted. We compare the performance of the policy profile $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}$ with the evolution obtained through the open-loop operation $u(t)=0$. During this period, we simulate a worst-case fluctuation on the baseline demand for the companies’ product by defining
$w(t)=\begin{cases}(\phantom{-}1,\phantom{-}1,\phantom{-}1,\phantom{-}1)&\text{for}\leavevmode\nobreak\ t\in[265,279];\\ (\phantom{-}1,-1,\phantom{-}1,-1)&\text{for}\leavevmode\nobreak\ t\in[293,307];\\ (-1,\phantom{-}1,-1,\phantom{-}1)&\text{for}\leavevmode\nobreak\ t\in[321,335];\\ (\phantom{-}1,\phantom{-}1,-1,-1)&\text{for}\leavevmode\nobreak\ t\in[349,363];\\ (-1,-1,-1,-1)&\text{for}\leavevmode\nobreak\ t\in[377,391],\end{cases}$
and $w(t)=0$ otherwise. The results show that the policy profile obtained by the SLS-BRD routine allows the companies to efficiently respond to changes in their local demands. In general, the players coordinate price changes to alleviate the deviations from the baseline demand caused by the disturbances. Moreover, the player’s actions satisfy the price-cap constraint $\frac{1}{N_{P}}|\sum_{p\in\mathcal{P}}u^{p}(t)|\leq 0.5$ in all disturbance scenarios. The best performance is observed for companies $p\in\{3,4\}$, whereas $p=1$ behaves noticeably worse than all players (especially for $w(t)=(1,-1,1,-1)$ and $w(t)=(-1,1,-1,1)$, when the open-loop operation seems to attain better results). This highlights the fact that fixed-points obtained through the SLS-BRD routine are not necessarily admissible GFNE, and thus might be unfavourable for a subset of players. Finally, we note that these policies are still not able to completely reject the effect of the extreme disturbances: Achieving zero-offset would require incorporating integral action into the player’s policies.
Figure 8: Market game, $t\in(260,400]$ days: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for an execution of the game $\mathcal{G}_{\infty}^{\text{LQ}}$. The dashed lines refer to the state-trajectories resulting from an open-loop operation of the market with $u(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$
V Concluding remarks
This work presents the SLS-BRD, an algorithm for generalised feedback Nash equilibrium seeking in $N_{P}$-players non-cooperative games. The method is based on the class of best-response dynamics algorithms for decentralised learning and consists of players updating and announcing a parametrisation of their policies until converging to an equilibrium. Because not updating control actions explicitly, this learning dynamics can be performed alongside the execution of the game. Using results from operator theory, we propose that convergence to a Nash equilibrium is geometric if certain conditions are met: Namely, that the best-response maps are contractive operators. Our approach leverages the System Level Synthesis framework to formulate each player’s best-response map as the solution to robust finite-horizon optimal control problems. This framework benefits the SLS-BRD algorithm also by allowing richer information patterns to be enforced directly at the synthesis level. After the main theoretical and practical aspects are discussed, the algorithm is demonstrated on an exemplary problems on the decentralised control of multi-agent systems.

### v3

@@
 Abstract
-This work proposes a policy learning algorithm for seeking generalised feedback Nash equilibria in $N_{P}$-players non-cooperative dynamic games.
-We consider linear-quadratic games with stochastic dynamics and design a best-response dynamics in which players update and communicate a parametrisation of their state-feedback policies.
-Our approach leverages the System Level Synthesis framework to formulate each player’s update rule as the solution to a tractable robust optimisation problem.
+This work proposes a policy learning algorithm for seeking generalised feedback Nash equilibria (GFNE) in $N_{P}$-player noncooperative dynamic games.
+We consider linear-quadratic games with stochastic dynamics and design a best-response dynamics in which players update and broadcast a parametrisation of their state-feedback policies.
+Our approach leverages the System Level Synthesis (SLS) framework to formulate each player’s update rule as the solution to a robust optimisation problem.
 Under certain conditions, rates of convergence to a feedback Nash equilibrium can be established.
 The algorithm is showcased in exemplary problems ranging from the decentralised control of unstable systems to competition in oligopolistic markets.
-Index Terms:
-Non-cooperative games, Feedback Nash equilibrium, Best-response dynamics, System level synthesis
-I Introduction
-Modern cyber-physical systems are often comprised of interacting subsystems operated locally by selfish decision-making agents.
-Ideally, agents operate these systems according to feedback policies that optimise local objectives, while satisfying global requirements and being robust to their rivals’ interference.
-However, the large-scale and decentralised nature of most applications, alongside the usual lack of coordination between agents, hinders most traditional approaches to policy design.
-Dynamic game theory provides an alternative framework based on the concept of competitive equilibria (e.g., the Nash equilibrium [1, 2]), which describe locally optimal, yet strategically stable, operating conditions for each non-cooperative agent.
-A methodology for policy design based on Nash equilibrium seeking thus presents a promising venue for enabling a decentralised control of cyber-physical systems.
-In general, solving a non-cooperative game can concern distinct goals: i) For the unbiased observer, to examine the behaviour of rational agents from a local to a global scale; ii) for the game designer, to actuate the self-interested players towards system-wide objectives; and iii) for the player, to determine locally optimal and robust policies in competitive environments.
-Regardless, computing a Nash equilibrium (NE) solution is a notoriously difficult task [3].
-Algorithmic game theory thus emerges as the field concerned with designing procedures to bridge this computational gap [4].
-An important class of algorithms are those which place to the players the task of converging towards a non-cooperative equilibrium: These include best-response dynamics (BRD, [5, 6, 7, 8, 9, 10]) and no-regret learning (NRL, [11, 12, 13, 14, 15]).
-Recently, online feedback optimisation has also been employed for equilibrium seeking problems [16].
-In general, these are fixed-point iteration methods centred on players updating their strategies, simultaneously and independently, using only the information available to them.
+This is the Accepted Manuscript of an article to appear in the IEEE Transactions on Automatic Control, available at: https://doi.org/10.1109/TAC.2025.3568560
+{IEEEkeywords}
+Noncooperative games, best-response dynamics, feedback Nash equilibrium, system level synthesis
+1 Introduction
+[…]
+Modern cyber-physical systems are often comprised of interacting subsystems operated locally by noncooperative decision-making agents.
+Ideally, each agent operate its subsystem according to a feedback policy that optimise local objectives, while satisfying global constraints and being robust to their rivals’ interference.
+However, the large-scale, decentralised, and multi-objective nature of such applications hinders most traditional approaches to policy design.
+Dynamic game theory provides an alternative framework through the concept of noncooperative equilibria (e.g., the generalized Nash equilibrium [1, 2]) describing efficient, yet strategically stable, strategies for each agent.
+Under a dynamic game setting, decision-making agents must design feedback policies (i.e., strategies) to operate their subsystems while aware that their choice affects (and is affected by) the other agents’ choice.
+An equilibrium would then correspond to a set of policies that satisfy the global constraints and is agreeable to all agents given their objectives.
+A policy design based on generalized Nash equilibrium seeking thus presents a promising venue for the decentralised control of cyber-physical systems.
+In general, solving a noncooperative game has distinct goals: i) For the agents, to design efficient and robust policies in competitive environments; ii) for the game designer, to examine (and potentially control) the local and global behaviour of rational agents.
+Regardless, obtaining a Nash equilibrium (NE) is a notoriously difficult task [3].
+Algorithmic game theory thus emerges as the field concerned with designing methods to bridge this computational gap [4].
+An important class of algorithms, denoted equilibrium-seeking methods, places the task of searching for an equilibrium on the players: These include best-response dynamics (BRD, [5, 6, 7]), no-regret learning (NRL, [8, 9, 10]), and operator-splitting [11, 12] methods.
+Broadly, these are fixed-point methods designed for (repeated) static games centred on the idea of players improving their strategies using only the information available to them.
+Aside from enabling the computation of Nash equilibria, these routines have the advantage of mimicking how noncooperative players would learn their policies in reality.
 In particular, best-response dynamics stands out as a simple, yet fundamental, model of policy learning for uncoordinated but communicating players.
-These algorithms have become important tools in economics and engineering, with applications including the design and control of networked systems [17, 18, 19], robotics [20, 21], and resource management [22, 23, 24].
-For dynamic games, any solution concept depends qualitatively on a description of the data available to each player; a notion termed information pattern [1].
-Under open-loop information patterns, when players only access the initial state of the game, the analysis and solution of NE problems are well-understood.
-Conversely, games under closed-loop (or feedback) information patterns, when players monitor the state of the game, are still under active research.
-A feedback Nash equilibrium (FNE) solution is often more attractive than its open-loop counterpart: It renders the players robust to disturbances and decision errors.
-However, designing routines for computing FNE is demanding for all but simple cases.
-A notably challenging class of games consists of generalised feedback Nash equilibrium (GFNE, [2]) problems: When players’ policies are required to satisfy coupled constraints (e.g., restrictions on the state of the game).
-To the best of our knowledge, methods for FNE seeking exist only for linear-quadratic games under restrictive assumptions [25, 26, 27, 28], unconstrained nonlinear and control-affine games [29, 30, 31, 32], and application-specific problems [33, 34].
-Recently, [35] expanded these results and presented a systematic (approximate) solution for GFNE problems.
-Despite remarkable, these efforts are restricted to finite-duration games and thus exclude decision-making processes in which agents are continuously operating the underlying system.
-Finally, a systematic solution is also lacking for GFNE games described by stochastic dynamics.
-In this work, we investigate algorithms for GFNE seeking in (stationary) difference games with linear-stochastic dynamics.
-Leveraging the System Level Synthesis (SLS, [36]) framework, we propose a best-response approach based on a parametrisation of all stabilising policies for the game.
-The algorithm consists of players iteratively, and simultaneously, updating the parameters of their individual policies, then announcing these changes through some communication network (Figure 1).
-Under our approach, each player’s update rule is formulated as the solution to a tractable robust optimisation problem.
-Moreover, these problems allow for constraints to be imposed directly on the policies’ structure, thus having the ability to encode information patterns at the synthesis level.
-Finally, this proposed learning dynamics do not depend on the state and actions applied to the system, and thus can be performed simultaneously with the game’s execution.
+This method has become an important tool in economics and engineering, with applications in networked systems [13, 14], robotics [15, 16, 17], and resource management [18, 19].
+We are interested in Nash equilibrium seeking algorithms for dynamic games in which the underlying system has linear, stochastic, and potentially unstable dynamics.
+The focus is on closed-loop perfect state information structures, as they yield equilibrium policies that are less sensitive to modelling and decision errors than their open-loop counterparts [1].
+Specifically, we consider the relevant task of players learning a generalised feedback Nash equilibrium (GFNE) of state-feedback policies which: i) stabilise the system against disturbances, ii) satisfy constraints on the control and state signals, and iii) incorporate some specific structure (e.g., encoding communication delays).
+In the current practice, feedback Nash equilibria is obtained by either applying dynamic programming principles [1, 20, 21, 22, 23, 24], deriving equivalent complementary problems [25, 26], or using iterative heuristics [27, 28].
+Recently, [29] extended the scope to provide a systematic method to compute (approximate) GFNE.
+While remarkable, these solutions cannot enforce either closed-loop stability, operational constraints, or design of the policy structure, in practice.
+Moreover, most of them still require some central coordinator to solve the game; the players themselves do not perform equilibrium-seeking.
+The available literature on decentralised policy learning is concentrated on the field of reinforcement learning, where the focus is on finite-duration games described by Markov decision processes [30, 31, 32, 33, 34].
+To the best of our knowledge, there are no equilibrium-seeking solutions to the aforementioned class of GFNE problems.
+In this work, we propose a GFNE seeking algorithm to bridge this research gap.
+Leveraging the System Level Synthesis (SLS, [35]) framework, we first recast each player’s policy synthesis problem as the search for an optimal system level response to disturbances.
+The system level response can be used to recover a corresponding optimal feedback policy, thus serving as its parametrisation.
+Under this setup, each player’s best-response strategy is the solution to a tractable optimisation problem which enforces closed-loop stability, operational constraints (on the control and state signals), and structural constraints (on the policy’s parameters), already at the synthesis level.
+We then design a best-response routine in which players update the parametrisation of their policies, in parallel, and then announce them through some communication network (Figure 1).
+The algorithm does not depend on the state and actions applied to the underlying system and thus can be executed simultaneously with its operation.
 In summary, our contributions are:
   1. (i)
-    A system-level best-response dynamics (SLS-BRD) algorithm for GFNE seeking in linear-stochastic dynamic games.
-Specifically, we design a class of best-response maps which compute an optimal system level parametrisation of a player’s policy in response to its rivals’ choices.
+    A realisation of the associated best-response mappings in GFNE problems as robust convex optimisation problems which are amenable to numerical solutions.
   2. (ii)
-    A realisation of the best-response mappings as robust finite-dimensional programs amenable to numerical solutions.
-Considering these best-responses, the SLS-BRD algorithm is formally an $\varepsilon$-GFNE seeking algorithm for an equilibrium gap $\varepsilon>0$ which can be explicitly established.
+    A system-level best-response dynamics (SLS-BRD) algorithm for GFNE seeking in dynamic games.
+In this equilibrium-seeking routine, players converge to a GFNE by iteratively updating the parametrisation of their policies as the best response to their rivals’ current policies.
   3. (iii)
-    For a specific but important class of dynamic games, an analysis of the conditions and rates of convergence of the SLS-BRD to feedback Nash equilibria.
-This policy learning algorithm is showcased in simulated experiments on the decentralised control of an unstable network and price management in a competitive oligopolistic market.
+    In the absence of shared constraints, a formal analysis of the convergence properties of the SLS-BRD algorithm.
+This policy learning algorithm is demonstrated in simulated experiments on the decentralised control of an unstable network and price management in a competitive oligopolistic market.
 Figure 1: SLS-BRD: Control architecture and the learning dynamics.
-The paper is organised as follows: Section II overviews the classes of (generalised) static and dynamic games, and the best-response dynamics algorithm.
-In Section III, we provide a system level parametrisation for linear-stochastic games, then design a best-response dynamics for GFNE seeking.
-Finally, Section IV illustrates this learning dynamics in simulated examples, and Section V provides some concluding remarks.
-Towards a concise presentation, only some essential theorems are proved in the main text: The remaining are in the Appendix.
-I-A Notation
-We use Latin letters to denote vectors (lowercase) and mappings (uppercase), and use boldface to distinguish signals, operators, and their respective spaces.
-Sets are in calligraphic font; exceptions are the usual $\mathbb{R}$ and $\mathbb{N}$, and the sets of $N\times N$ symmetric ($\mathbb{S}^{N}$), positive semidefinite ($\mathbb{S}^{N}_{+}$), and positive definite matrices ($\mathbb{S}^{N}_{++}$).
-In particular, sequences are denoted $\bm{x}=(x_{t})_{t\in\mathcal{I}}$ for a countable set $\mathcal{I}\subseteq\mathbb{N}$, or $\bm{x}=(x_{t})_{t=0}^{T}$ if =$\mathcal{I}=\{0,\ldots,T\}$.
-For $p\in(0,\infty)$, we use the space of $N_{x}$-dimensional vector-sequences $\ell^{N_{x}}_{p}(\mathcal{I})=\{\bm{x}:\|\bm{x}\|_{\ell_{p}}=(\sum_{t\in\mathcal{I}}\|x_{t}\|^{p})^{1/p}<\infty\}$, with $\ell^{N_{x}}_{\infty}$ and ,$\ell^{N_{x}}_{\infty,e}$ the space of all bounded sequences and all sequences, respectively.
-$\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{Y}})$ is the set of all bounded linear operators $\bm{A}:\bm{\mathcal{X}}\to\bm{\mathcal{Y}}$ and we sometimes denote transformed signals by $\bm{A}\bm{x}=(Ax_{t})_{t\in\mathcal{I}}$.
-Whenever $(\bm{\mathcal{X}},\bm{\mathcal{Y}})$ are finite-dimensional, we let $M_{A}$ refer to a matrix representation of $\bm{A}$.
-We use the standard definition of Hardy spaces $\mathcal{H}_{\infty}$ and $\mathcal{RH}_{\infty}$, and write $\frac{1}{z}\mathcal{RH}_{\infty}$ to refer to the set of real-rational strictly proper transfer functions.
-Finally, some standard signals and operators used in this paper are: The impulse signal $\bm{\delta}=(\delta_{t})_{t\in\mathcal{I}}$, the identity operator $I$ and the identity matrix $I_{N_{x}}$, and the shift operator $\bm{S_{+}}:(x_{0},x_{1},\ldots)\mapsto(0,x_{0},\ldots)$.
-We distinguish set-valued mappings from ordinary functions using the notation $F:\mathcal{X}\rightrightarrows\mathcal{Y}$.
-For any tuple $s=(s^{p})_{p\in\mathcal{P}}\in\mathcal{S}$ we frequently write $s=(s^{p},s^{-p})$ to highlight the element $s^{p}$; this should not be interpreted as a re-ordering.
-Similarly, if $\mathcal{S}=\prod_{p\in\mathcal{P}}\mathcal{S}^{p}$, we define the product $\mathcal{S}^{-p}=\prod_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\mathcal{S}^{\tilde{p}}$.
-II Non-cooperative games and best-response dynamics
+A set of players ($1,\ldots,N_{P}$) control a decentralised dynamical system by measuring its state ($\bm{x}$) and then applying actions ($\bm{u}^{1},\ldots,\bm{u}^{N_{P}}$) based on state-feedback policies ($\bm{K}^{1},\ldots,\bm{K}^{N_{P}}$) whose parametrisations ($\Phi^{1},\ldots,\Phi^{N_{P}}$) are chosen as best responses to each others’ strategies, simultaneously, and broadcasted via a communication network.
+The paper is organised as follows: Section 2 overviews the classes of (generalised) static and dynamic games, and the best-response dynamics algorithm.
+In Section 3, we provide a system level parametrisation for linear-quadratic games, then design a best-response dynamics for GFNE seeking.
+Finally, Section 4 illustrates this approach in exemplary problems and Section 5 provides concluding remarks.
+Towards a concise presentation, only essential results are given in the main text: We refer to the Supplementary Material for remaining details.
+1.1 Notation
+We use Latin letters to denote vectors and functions, and boldface to distinguish signals, operators, and their respective spaces.
+Sets are in calligraphic font; exceptions are the usual $\mathbb{R}$ and $\mathbb{N}$, and the sets of $(N{\times}N)$ symmetric ($\mathbb{S}^{N}$), positive semidefinite ($\mathbb{S}^{N}_{+}$), and positive definite matrices ($\mathbb{S}^{N}_{++}$).
+In particular, sequences are written as $\bm{x}=(x_{t})_{t\in\mathcal{I}}$ for a countable set $\mathcal{I}\subseteq\mathbb{N}$, or $\bm{x}=(x_{t})_{t=0}^{T}$ if =$\mathcal{I}=\{0,\ldots,T\}$.
+For $p\in(0,\infty)$, we define the space of $N_{x}$-dimensional vector-sequences $\ell^{N_{x}}_{p}(\mathcal{I})=\{\bm{x}:\|\bm{x}\|_{\ell_{p}}=(\sum_{t\in\mathcal{I}}\|x_{t}\|^{p})^{1/p}<\infty\}$, with $\ell^{N_{x}}_{\infty}$ the space of all bounded sequences.
+The set $\bm{\mathcal{Y}}^{\bm{\mathcal{X}}}$ denotes all relations $\bm{A}:\bm{\mathcal{X}}\to\bm{\mathcal{Y}}$ with $\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{Y}})\subseteq\bm{\mathcal{Y}}^{\bm{\mathcal{X}}}$ being the set of bounded linear operators.
+We sometimes write transformed signals as $\bm{A}\bm{x}=(Ax_{t})_{t\in\mathcal{I}}$.
+We use the standard definition of Hardy spaces $\mathcal{H}_{\infty}$ and $\mathcal{RH}_{\infty}$, and write $\frac{1}{z}\mathcal{RH}_{\infty}$ for the set of real-rational strictly-proper transfer functions.
+Finally, signals and operators used in this paper include: The impulse signal $\bm{\delta}=(\delta_{t})_{t\in\mathcal{I}}$, the identity operator $I$ and matrix $I_{N_{x}}$, the matrices $\bm{1}_{n\times m}$ and $\bm{0}_{n\times m}$ of all 1’s and 0’s, respectively, the shift operator $\bm{S_{+}}:(x_{0},x_{1},\ldots)\mapsto(0,x_{0},\ldots)$, and the Kronecker and Hadamard products $\otimes$ and $\odot$, respectively.
+[…]
+2 Noncooperative games and best-response dynamics
 A (static) $N_{P}$-player game, denoted by a tuple
@@
 The strategy spaces $S^{p}$ ($\forall p\in\mathcal{P}$) determine the actions available to the players, with the mappings $S^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ restricting this choice based on the actions from their opponents.
-As such, both the players’ objectives and feasible strategies depend explicitly on their rivals’ strategies.
-Finally, the players are assumed to be rational, non-cooperative, and acting simultaneously.
+As such, both the players’ objectives and feasible strategies depend explicitly on their competitors’ actions.
+Finally, the players are assumed to be rational, noncooperative, and acting simultaneously.
 A solution to the game $\mathcal{G}$ is understood as a strategy profile $s=(s^{1},\ldots,s^{N_{P}})\in\mathcal{S}$, $\mathcal{S}=\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}$, having some specified property that makes it agreeable to all players if they act rationally.
-In non-cooperative settings, a widely accepted solution concept is that of a generalized Nash equilibrium: The game is solved when no player can improve its objective by unilaterally deviating from the agreed strategy profile.
+In noncooperative settings, a widely accepted solution concept is that of a generalized Nash equilibrium: The game is solved when no player can improve its objective by unilaterally deviating from the agreed strategy profile.
 Formally,
 Definition 1.
-A strategy profile $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})\in\mathcal{S}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}$ if
-$L^{p}(s^{p^{*}},s^{-p^{*}})\leq\min_{s^{p}\in S^{p}(s^{-p^{*}})}L^{p}(s^{p},s^{-p^{*}}).$ (2)
+A strategy profile $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})\in\mathcal{S}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}$ if
+$L^{p}(s^{p^{\star}},s^{-p^{\star}})\leq\min_{s^{p}\in S^{p}(s^{-p^{\star}})}L^{p}(s^{p},s^{-p^{\star}}).$ (2)
 holds for every player $p\in\mathcal{P}$.
@@
 […]
-is not a singleton.
-As such, players might favour different solutions based on their individual objectives.
-We thus seek to characterise which solutions are “acceptable”, in the sense that no other equilibrium can improve the objective of all players simultaneously.
-Consider the partial ordering of $\Omega_{\mathcal{G}}$ defined by
-$s^{*}\prec\tilde{s}^{*}\leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ L^{p}(s^{p^{*}})\leq L^{p}(\tilde{s}^{p^{*}}),\quad\forall p\in\mathcal{P},$ (3)
-with at least one inequality being strict, for any pair $s^{*},\tilde{s}^{*}\in\Omega_{\mathcal{G}}$.
-A solution $s^{*}\in\Omega_{\mathcal{G}}$ is then characterised as an admissible GNE if there are no other equilibrium $\tilde{s}^{*}\in\Omega_{\mathcal{G}}$ such that $\tilde{s}^{*}\prec s^{*}$.
-Note that an admissible GNE can still be favourable to only a subset of players (that is, $\tilde{s}^{*}\nprec s^{*}\nprec\tilde{s}^{*}$), and thus it is possible that $\mathcal{G}$ has no solution that is “fair” to all players.
+is not a singleton and can include strategy profiles that favour a specific subset of players (that is, non-admissible GNEs [1]).
 The game might also not admit any GNE (i.e., $\Omega_{\mathcal{G}}=\emptyset$), then being characterised as unsolvable.
@@
 Although restrictive, these assumptions still cover a broad class of problems of practical relevance.
+Input: Game $\mathcal{G}\coloneqq(\mathcal{P},\{\mathcal{S}^{p}\}_{p\in\mathcal{P}},\{L^{p}\}_{p\in\mathcal{P}})$
+Output: GNE $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})$
+1
+2 Initialize $s_{0}\coloneqq(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k\coloneqq 0$;
+3 for $k=0,1,2,\ldots$ do
+4       if $s_{k}\in\Omega_{\mathcal{G}}$ then return $s_{k}$;
+5       for $p\in\mathcal{P}$ do
+6             Update $s^{p}_{k+1}\in T^{p}(s^{p}_{k},R^{p}(s_{k})\mid L^{p})$;
+7
+8
+Algorithm 1 Prototypical learning dynamics
 We investigate algorithms for solving $\mathcal{G}$.
@@
 In this direction, consider that $\mathcal{G}$ admits episodic repetitions, and let $s_{k}\coloneqq(s_{k}^{1},\ldots,s_{k}^{N_{P}})$ be the strategy profile taken by players $\mathcal{P}$ at the $k$-th episode.
-A prototypical learning procedure for equilibrium seeking is outlined in Algorithm 1, where
-  • $T^{p}:\mathcal{S}^{p}\times\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ describes how the $p$-th player updates its strategy, based on its opponents’ predicted next actions and given its individual objective;
-  • $R^{p}:\mathcal{S}\rightrightarrows\mathcal{S}^{-p}$ describes how the $p$-th player’s predicts its opponents’ strategies for the next episode, based on the strategy profile currently being played.
-Input: Game $\mathcal{G}\coloneqq(\mathcal{P},\{\mathcal{S}^{p}\}_{p\in\mathcal{P}},\{L^{p}\}_{p\in\mathcal{P}})$
-Output: GNE $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})$
-1
-2 Initialize $s_{0}=(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k=0$;
-3 for $k=1,2,\ldots$ do
-4       if $s_{k}\in T(s_{k})$ then return $s_{k}$;
-5       for $p\in\mathcal{P}$ do
-6             Update $s^{p}_{k+1}\in T^{p}(s^{p}_{k},R^{p}(s_{k})\mid L^{p})$;
-7
-8
-Algorithm 1 Prototypical learning dynamics
-The procedure in Algorithm 1 belongs to the class of fixed-point methods: Its termination implies that $s^{*}$ is a fixed-point of both $T^{p}$ and $R^{p}$, that is, $s^{*}\in T^{p}(s^{p^{*}},R^{p}(s^{*}))\subseteq T^{p}(s^{p^{*}},s^{-p^{*}})$.
-In its general form, the conditions (and convergence rates) for these learning dynamics to approach an equilibrium are difficult to establish.
-In this work, we build upon a fundamental instance from this class of algorithms: The best-response dynamics (BRD).
+A prototypical learning routine for equilibrium seeking is outlined in Algorithm 1, with
+  • $T^{p}:\mathcal{S}^{p}\times\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ describing how the $p$-th player updates its strategy, based on a prediction of its opponents’ next actions, given its individual objective;
+  • $R^{p}:\mathcal{S}\rightrightarrows\mathcal{S}^{-p}$ describing how the $p$-th player predicts its opponents’ strategies for the next episode, based on the strategy profile currently being played.
+Algorithm 1 belongs to the class of fixed-point methods: Its termination implies that $s^{\star}$ is a fixed-point of both $T^{p}$ and $R^{p}$, that is, $s^{\star}\in T^{p}(s^{p^{\star}},R^{p}(s^{\star}))\subseteq T^{p}(s^{p^{\star}},s^{-p^{\star}})$.
+In its general form, it is difficult to establish the conditions (and convergence rates) for these learning dynamics to approach an equilibrium.
+In this work, we build upon a specific yet fundamental instance of this algorithm: The best-response dynamics (BRD).
 This routine is overviewed in the following.
 Best-response dynamics
-Let the mapping $BR^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$,
+Let the map $BR^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$,
 $BR^{p}(s^{-p})\coloneqq\operatorname*{arg\,min}_{s^{p}\in S^{p}(s^{-p})}L^{p}(s^{p},s^{-p})$ (3)
@@
 Collectively, $BR(s)\coloneqq BR^{1}(s^{-1})\times\cdots\times BR^{N_{P}}(s^{-N_{P}})\subseteq\mathcal{S}$ is the joint best-response to any given profile $s\in\mathcal{S}$.
-Then,
-Theorem 1.
-A strategy profile $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})\in\mathcal{S}$ is a GNE for $\mathcal{G}$ if and only if $s^{*}\in BR(s^{*})$ or, equivalently,
-$s^{p^{*}}\in BR^{p}(s^{-p^{*}}),\quad\forall p\in\mathcal{P}.$ (5)
-The task of computing a Nash equilibrium can thus be translated into searching for a fixed-point of the set-valued mapping $BR:\mathcal{S}\rightrightarrows\mathcal{S}$.
-The set of GNE solutions for $\mathcal{G}$ is the set of all such fixed-points, $\Omega_{\mathcal{G}}\coloneqq\{s^{*}\in\mathcal{S}:s^{*}\in BR(s^{*})\}$.
-A natural procedure for GNE seeking thus consists of players adapting their strategies towards best-responses to their rivals’ strategies, which they assume will remain constant.
+From Definition 1, a strategy profile $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})\in\mathcal{S}$ is a GNE for $\mathcal{G}$ whenever $s^{\star}\in BR(s^{\star})$ or, equivalently,
+$s^{p^{\star}}\in BR^{p}(s^{-p^{\star}}),\quad\forall p\in\mathcal{P}.$ (4)
+The task of computing a Nash equilibrium can thus be translated into the search for a fixed-point of the set-valued mapping $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ [7].
+The set of GNE solutions for $\mathcal{G}$ is the set of all such fixed-points, $\Omega_{\mathcal{G}}\coloneqq\{s^{\star}\in\mathcal{S}:s^{\star}\in BR(s^{\star})\}$.
+A natural procedure for GNE seeking consists of players adapting their strategies towards best-responses to their rivals’ strategies, which they assume will remain constant.
 Formally,
 $T^{p}(s^{p}_{k},R^{p}(s_{k}))\coloneqq(1{-}\eta)s^{p}_{k}+\eta BR^{p}(R^{p}(s_{k})),$ (5)
-given $R^{p}(s_{k})=s_{k}^{-p}$ and a learning rate factor of $\eta\in(0,1)$. This learning dynamics, summarised in Algorithm 2, is known as best-response dynamics or fictitious play.
+given $R^{p}(s_{k})=s_{k}^{-p}$ and a learning rate factor of $\eta\in(0,1)$. This learning dynamics, summarised in Algorithm 2, is known as (discrete-time) best-response dynamics.
 Input: Game $\mathcal{G}\coloneqq(\mathcal{P},\{\mathcal{S}^{p}\}_{p\in\mathcal{P}},\{L^{p}\}_{p\in\mathcal{P}})$
-Output: GNE $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})$
+Output: GNE $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})$
 1
-2 Initialize $s_{0}=(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k=0$;
+2 Initialize $s_{0}\coloneqq(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k\coloneqq 0$;
 3 for $k=0,1,2,\ldots$ do
@@
 Algorithm 2 Best-Response Dynamics (BRD)
-After each episode, the strategy profile gets updated as
+After each episode, the strategy profile is updated to
 $s_{k+1}=T(s_{k})=(1-\eta)s_{k}+\eta BR(s_{k}),$ (6)
 given the global update rule $T=(1-\eta)I+\eta BR$. Notably, the mappings $T$ and $BR$ share the same set of fixed-points: The GNEs $\Omega_{\mathcal{G}}$. We can then establish the following result.
-Theorem 2.
-Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be a non-expansive mapping.
-Then, the best-response dynamics $s_{k+1}=T(s_{k})$ converge monotonically to a GNE solution $s^{*}\in\Omega_{\mathcal{G}}$, that is,
-$\lim_{k\to\infty}\inf_{s^{*}\in\Omega_{\mathcal{G}}}\|T(s_{k})-s^{*}\|=0$ (8)
+Lemma 1.
 […]
-This convergence result stems from fixed-point theory, where the BRD algorithm is interpreted as belonging to the class of averaged (or Mann-Krasnosel‘skii) iteration methods [38, 39].
-For non-generalised quadratic problems, we show in Section III-B that the best-response mapping $BR$ is non-expansive under specific conditions on $L^{1},\ldots,L^{N_{P}}$ Importantly, if the best-response mapping is a contraction, then the BRD converge geometrically to a GNE solution $s^{*}\in\Omega_{\mathcal{G}}$, which is unique.
-Theorem 3.
+This convergence result stems from fixed-point theory, where the BRD algorithm is interpreted as belonging to the class of averaged (or Krasnosel‘skii-Mann) iteration methods (see [37] for a formal proof).
+Moreover, if the best-response mapping $BR$ is a contraction then so is $T$ and the BRD must converge geometrically to a GNE $s^{\star}\in\Omega_{\mathcal{G}}$, which is unique [37]:
+Lemma 2.
 Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be $L_{BR}$-Lipschitz, $L_{BR}<1$.
-Then, from any feasible $s_{0}\in\mathcal{S}$, the best-response dynamics $s_{k+1}=T(s_{k})$ converge to the unique GNE $s^{*}\in\Omega_{\mathcal{G}}$ with rate
-$\cfrac{\|s_{k}-s^{*}\|}{\|s_{0}-s^{*}\|}\leq\big{(}(1{-}\eta)+\eta L_{BR}\big{)}^{k}$ (9)
+Then, from any feasible $s_{0}\in\mathcal{S}$, the best-response dynamics $s_{k+1}=T(s_{k})$ converge to the unique GNE $s^{\star}\in\Omega_{\mathcal{G}}$ with rate
+$\cfrac{\|s_{k}-s^{\star}\|}{\|s_{0}-s^{\star}\|}\leq\big{(}(1{-}\eta)+\eta L_{BR}\big{)}^{k}$ (7)
 […]
@@
 (3) numerically).
-However, such inexact averaged operators are still known to converge under specific assumptions on the accuracy of this approximation [40].
-The learning rate $\eta$ also plays a central role in the numerical stability of the BRD algorithm: A careful choice is required to ensure that strategy updates do not escape the feasible set, that is, to ensure that $T(s_{k})\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$ for all $s_{k}\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$.
-Otherwise, the learning dynamics become undefined and the aforementioned convergence results do not apply.
+However, such inexact averaged operators are still known to converge under reasonable assumptions on the accuracy of this approximation [38].
+The learning rate $\eta$ plays a central role in the numerical stability of the BRD algorithm: A careful choice is required to ensure that strategy updates do not escape the feasible set, that is, to ensure that $T(s_{k})\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$ for all $s_{k}\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$.
+The choice of $\eta$ can also ensure convergence in specific games in which $L_{BR}>1$ (see the Supplementary Material).
 For non-generalised games, $T$ trivially satisfy the constraints for any $\eta\in(0,1)$ and thus a careful design of the learning rate might not be necessary.
@@
 In this case, interrupting the best-response dynamics at some episode $k_{f}>0$ will produce a strategy profile $s_{k_{f}}\in\mathcal{S}$ for which
-$L^{p}(s_{k_{f}}^{p},s_{k_{f}}^{-p})\leq\min_{s^{p}\in S^{p}(s_{k_{f}}^{-p})}L^{p}(s^{p},s_{k_{f}}^{-p})+\varepsilon$ (10)
+$L^{p}(s_{k_{f}}^{p},s_{k_{f}}^{-p})\leq\textstyle\min_{s^{p}\in S^{p}(s_{k_{f}}^{-p})}L^{p}(s^{p},s_{k_{f}}^{-p})+\varepsilon$ (8)
 […]
-II-A Infinite-horizon dynamic games
-An infinite-horizon dynamic $N_{P}$-player game, denoted by tuple
+2.1 Infinite-horizon dynamic games
+A dynamic $N_{P}$-player game, denoted by a tuple
 $\mathcal{G}_{\infty}\coloneqq(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}),$ (9)
@@
 For each realisation $\bm{w}\in\bm{\mathcal{W}}$ and initial $x_{0}$, the state is explicitly expressed as $\bm{x}=\bm{F_{w}}\bm{u}$ via the causal affine operator
-$\bm{F_{w}}:\bm{u}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\displaystyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{u}^{p}+\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\Big{)},$ (13)
+$\bm{F_{w}}:\bm{u}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\textstyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{u}^{p}+\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\Big{)},$ (11)
 with $\bm{A}:\bm{x}\mapsto(Ax_{t})_{t\in\mathbb{N}}$ and $\bm{B}^{p}:\bm{u}^{p}\mapsto(B^{p}u^{p}_{t})_{t\in\mathbb{N}}$.
 Because known, the dependency on $x_{0}$ is omitted to simplify notation.
-Moreover, we assume $\mathrm{E}w_{t}=0$ and $\mathrm{E}(w_{t}w_{t^{\prime}}^{{\mkern-1.5mu\mathsf{T}}})=\delta_{t-t^{\prime}}\Sigma_{w}$, given a covariance matrix $\Sigma_{w}\in\mathbb{S}_{++}^{N_{x}}$, for every $t,t^{\prime}\in\mathbb{N}$.
+Moreover, we assume $\mathrm{E}w_{t}=0$ and $\mathrm{E}(w_{t+\tau}w_{t}^{{\mkern-1.5mu\mathsf{T}}})=\delta_{\tau}\Sigma_{w}$, given a covariance matrix $\Sigma_{w}\in\mathbb{S}_{++}^{N_{x}}$, for every $t,\tau\in\mathbb{N}$.
 Finally, the sets $\bm{\mathcal{X}}$, $\bm{\mathcal{U}}^{p}$ ($\forall p\in\mathcal{P}$), and $\bm{\mathcal{W}}$ define all permissible state, action, and noise sequences; they take the form
-$\bm{\mathcal{X}}$$\coloneqq\{\bm{x}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):x_{t}\in\mathcal{X},\leavevmode\nobreak\ t\in\mathbb{N}\};$
-$\bm{\mathcal{U}}^{p}$$\coloneqq\{\bm{u}^{p}\in\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):u^{p}_{t}\in\mathcal{U}^{p},\leavevmode\nobreak\ t\in\mathbb{N}\};$
-$\bm{\mathcal{W}}$$\coloneqq\{\bm{w}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):w_{t}\in\mathcal{W},\leavevmode\nobreak\ t\in\mathbb{N}\},$
+$\bm{\mathcal{X}}$$\coloneqq\{\bm{x}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):x_{t}\in\mathcal{X},~{}t\in\mathbb{N}\};$
+$\bm{\mathcal{U}}^{p}$$\coloneqq\{\bm{u}^{p}\in\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):u^{p}_{t}\in\mathcal{U}^{p},~{}t\in\mathbb{N}\};$
+$\bm{\mathcal{W}}$$\coloneqq\{\bm{w}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):w_{t}\in\mathcal{W},~{}t\in\mathbb{N}\},$
 given sets $\mathcal{X}\subseteq\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}\subseteq\mathbb{R}^{N_{u}^{p}}$ ($\forall p\in\mathcal{P}$), and $\mathcal{W}\subseteq\mathbb{R}^{N_{x}}$.
-In dynamic stochastic games, each player decides on a plan of action $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$ to minimize its objective functional
+In infinite-horizon games, each player chooses a plan of action $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$ to minimize its objective functional
 $J^{p}(\bm{u}^{p},\bm{u}^{-p})\coloneqq\mathrm{E}\left[\sum_{t=0}^{\infty}L^{p}(x_{t},u^{p}_{t},u^{-p}_{t})\right],$ (12)
 defined by cost function $L^{p}:\mathcal{X}\times\mathcal{U}^{1}\times\cdots\times\mathcal{U}^{N_{P}}\to\mathbb{R}$.
-The mappings $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ restrict the permissible strategies for each player based on their rivals’ strategies.
-Under this formulation, the dynamic game $\mathcal{G}_{\infty}$ is stationary and can be interpreted as a static game defined on the appropriate functional spaces.
-A plan of action $\bm{u}=(\bm{u}^{1},\ldots,\bm{u}^{N_{P}})\in\bm{\mathcal{U}}$ can then be characterised as a generalised Nash Equilibrium solution to $\mathcal{G}_{\infty}$ when no player can improve its objective by unilaterally deviating from this agreed profile.
+The mappings $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ restrict the permissible actions for each player based on its rivals’ strategies.
+Under this setup, the dynamic game $\mathcal{G}_{\infty}$ is stationary and can be interpreted as a static game on the appropriate functional spaces.
+A plan of action $\bm{u}=(\bm{u}^{1},\ldots,\bm{u}^{N_{P}})\in\bm{\mathcal{U}}$ can then be characterised as a GNE solution to $\mathcal{G}_{\infty}$ when no player can improve its objective by unilaterally deviating from this profile.
 Formally,
 Definition 2.
-A strategy profile $\bm{u}^{*}=(\bm{u}^{1^{*}},\ldots,\bm{u}^{N_{P}^{*}})\in\bm{\mathcal{U}}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}_{\infty}$ if
-$J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{*}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{*}})$ (15)
+A strategy profile $\bm{u}^{\star}=(\bm{u}^{1^{\star}},\ldots,\bm{u}^{N_{P}^{\star}})\in\bm{\mathcal{U}}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}_{\infty}$ if
+$J^{p}(\bm{u}^{p^{\star}},\bm{u}^{-p^{\star}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{\star}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{\star}})$ (13)
 holds for every player $p\in\mathcal{P}$.
@@
 […]
-is not necessarily a singleton and can include non-admissible equilibria.
-The game is considered unsolvable if $\Omega_{\mathcal{G}_{\infty}}=\emptyset$.
-The following assumptions are taken for this class of games:
+is not necessarily a singleton and the game is considered unsolvable if $\Omega_{\mathcal{G}_{\infty}}=\emptyset$. The following assumptions are taken:
 Assumption 2.
@@
     the mapping $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ takes the form
-    $U^{p}(\bm{u}^{-p})\coloneqq\{\bm{u}^{p}\in\bm{\mathcal{U}}^{p}:(\bm{u}^{p},\bm{u}^{-p})\in\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\},$
-    where $\bm{F_{w}}^{\dagger}$ is a left-inverse of the operator in Eq. (13) and
-    $\bm{\mathcal{U}}_{\mathcal{G}}=\{\bm{u}\in\textstyle\prod_{p\in\mathcal{P}}\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):(u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},\leavevmode\nobreak\ t\in\mathbb{N}\}$
-    is a global constraint set shared by all players.
-The sets $\mathcal{U}^{p}$, $\mathcal{U}_{\mathcal{G}}$, and $\mathcal{X}$ are all non-empty, compact, and convex.
-Finally, we have that $\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})\neq\emptyset$.
-These conditions are analogous to those of Assumption 1: They ensure the existence of GNE solutions to $\mathcal{G}_{\infty}$, that is, $\Omega_{\mathcal{G}_{\infty}}\neq\emptyset$.
-In this case, the players not only share a global constraint $\bm{\mathcal{U}}_{\mathcal{G}}$, but are also required to ensure that state trajectories lie in a feasible set, $x\in\bm{\mathcal{X}}$, against all possible noise realisations.
-These constraints often describe operational desiderata and/or limitations concerning the game $\mathcal{G}_{\infty}$.
-Moreover, note that Assumption 2 implies that the operator $\bm{F_{w}}$ is injective, that is, that distinct control signals cannot produce the same state trajectory.
-A sufficient condition for this property consists on the matrix $B=[B^{1}\leavevmode\nobreak\ B^{2}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ B^{N_{P}}]$ being full-rank.
-Any equilibrium $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ incorporates an open-loop information pattern: Actions $u_{t}^{*}=(u_{t}^{1^{*}},\ldots,u_{t}^{N_{P}^{*}})$ depend explicitly only on initial $x_{0}\in\mathcal{X}$ and stage $t\in\mathbb{N}$.
+    $U^{p}(\bm{u}^{-p})\coloneqq\{\bm{u}^{p}\in\bm{\mathcal{U}}^{p}:(\bm{u}^{p},\bm{u}^{-p})\in\bm{\mathcal{U}}_{\mathcal{G}}\},$
+    given the global constraint set
+    $\bm{\mathcal{U}}_{\mathcal{G}}=\{\bm{u}\in\ell_{\infty}^{N_{u}}(\mathbb{N}):(F_{w}u)_{t}\in\mathcal{X},~{}u_{t}\in\mathcal{U}_{\mathcal{G}},~{}t\in\mathbb{N}\}.$
+    The sets $\mathcal{U}^{p}$, $\mathcal{U}_{\mathcal{G}}$, and $\mathcal{X}$ are all nonempty, compact, and convex.
+Finally, we have that $\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})\neq\emptyset$.
+These conditions are analogous to those of Assumption 1: They aim to ensure the existence of GNE solutions to $\mathcal{G}_{\infty}$, that is, $\Omega_{\mathcal{G}_{\infty}}\neq\emptyset$.
+Here, the shared constraints $\bm{\mathcal{U}}_{\mathcal{G}}$ also require the players to ensure that state trajectories lie in a feasible set ($x\in\bm{\mathcal{X}}$) against all realisations of the noise.
+These constraints describe operational desiderata and/or limitations in the game.
+The equilibria $\bm{u}^{\star}\in\Omega_{\mathcal{G}_{\infty}}$ have an open-loop information pattern: Actions $u_{t}^{\star}=(u_{t}^{1^{\star}},\ldots,u_{t}^{N_{P}^{\star}})$ depend explicitly only on initial state $x_{0}\in\mathcal{X}$ and stage-index $t\in\mathbb{N}$.
 A plan of action with such representation is undesirable, as players become sensible to noise disturbances and decision errors.
-Conversely, a state-feedback decision policy, $\bm{u}^{*}=K(\bm{x})$ for some $K:\bm{\mathcal{X}}\to\bm{\mathcal{U}}$, often carry desirable robustness properties.
-In this sense, a feedback (respectively, open-loop) representation of $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ is said to be strongly (weakly) time consistent [1].
+Conversely, state-feedback policies $\bm{u}^{\star}=K(\bm{x})$, for some $K:\bm{\mathcal{X}}\to\bm{\mathcal{U}}$, can detect such errors and provide corrective actions.
+A feedback (respectively, open-loop) representation of $\bm{u}^{\star}\in\Omega_{\mathcal{G}_{\infty}}$ is thus said to be strongly (weakly) time consistent [1].
 In this work, we investigate state-feedback solutions to the game $\mathcal{G}_{\infty}$.
@@
 $\bm{u}^{p}\coloneqq\bm{K}^{p}\bm{x},\quad\bm{K}^{p}:\bm{x}\mapsto\bm{\Phi^{p}}*\bm{x},$ (14)
-given a causal operator $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\subseteq\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}}^{p})$ defined by its convolution kernel $\bm{\Phi}^{p}=(\Phi^{p}_{n})_{n\in\mathbb{N}}\in\ell_{1}(\mathbb{N})$.
+given a linear causal operator $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\subseteq\mathcal{L}(\ell_{\infty}^{N_{x}},\ell_{\infty}^{N_{u}^{p}})$ defined by its convolution kernel $\bm{\Phi}^{p}=(\Phi^{p}_{n})_{n\in\mathbb{N}}\in\ell_{1}(\mathbb{N})$.
 The sets $\{\bm{\mathcal{C}}^{p}\}_{p\in\mathcal{P}}$ describe the operators that satisfy some $\mathcal{G}_{\infty}$-related restrictions (e.g., information patterns incurred by communication, actuation, and sensing delays).
-In this setting, players do not devise their actions explicitly but rather by designing a state-feedback policy profile $\bm{K}\coloneqq(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})\in\bm{\mathcal{C}}$, with $\bm{\mathcal{C}}=\bm{\mathcal{C}}^{1}\times\cdots\times\bm{\mathcal{C}}^{N_{P}}$.
+In this setup, players do not plan their actions explicitly but rather by designing a state-feedback policy profile $\bm{K}\coloneqq(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})\in\bm{\mathcal{C}}$, with $\bm{\mathcal{C}}=\bm{\mathcal{C}}^{1}\times\cdots\times\bm{\mathcal{C}}^{N_{P}}$.
 The solution concept that naturally arises is that of a generalized feedback Nash equilibrium.
 Definition 3.
-A policy profile $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$ is a generalized feedback Nash equilibrium (GFNE) for $\mathcal{G}_{\infty}$ if
-$J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{*}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{*}}),$ (17)
-where $\bm{u}^{*}\in\mathrm{Ker}(I-\bm{K}^{*}\bm{F_{w}})$, holds for every $p\in\mathcal{P}$.
+A policy profile $\bm{K}^{\star}=(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$ is a generalized feedback Nash equilibrium (GFNE) for $\mathcal{G}_{\infty}$ if
+$J^{p}(\bm{u}^{p^{\star}},\bm{u}^{-p^{\star}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{\star}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{\star}}),$ (15)
+where $\bm{u}^{\star}\in\mathrm{Ker}(I-\bm{K}^{\star}\bm{F_{w}})$, holds for every $p\in\mathcal{P}$.
 The set of GFNE that solve $\mathcal{G}_{\infty}$ is defined as
 […]
-We consider $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ to be admissible only if it renders the game stable, that is, if the closed-loop evolution
-$\bm{x}^{*}=\big{(}I-\bm{S_{+}}(\bm{A}-\textstyle\sum_{p\in\mathcal{P}}\bm{B}^{p}\bm{K}^{p^{*}})\big{)}^{-1}\big{(}\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\big{)}$
-is bounded ($\bm{x}^{*}\in\ell_{\infty}^{N_{x}}$) for all bounded noise ($\bm{w}\in\ell_{\infty}^{N_{x}}$).
+We consider $\bm{K}^{\star}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ to be admissible only if it renders the game stable, that is, if the closed-loop evolution
+$\bm{x}^{\star}=\big{(}I-\bm{S_{+}}(\bm{A}-\textstyle\sum_{p\in\mathcal{P}}\bm{B}^{p}\bm{K}^{p^{\star}})\big{)}^{-1}\big{(}\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\big{)}$
+is bounded ($\bm{x}^{\star}\in\ell_{\infty}^{N_{x}}$) for all bounded noise ($\bm{w}\in\ell_{\infty}^{N_{x}}$).
 A policy satisfying this requirement is said to be stabilising.
-From Assumption 2, we have $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}\neq\emptyset$ when $\bm{\mathcal{C}}=\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}})$.
-For a more specific $\bm{\mathcal{C}}\subset\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}})$, establishing the existence (and, especially, uniqueness) of a solution is demanding [41].
-In practice, the set $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ can be constructed from open-loop equilibria $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ by i) parametrising the set of all possible trajectories $\{\bm{x}_{\bm{w}}^{*}=\bm{F_{w}}\bm{u}^{*}\}_{\bm{w}\in\bm{\mathcal{W}}}$, then ii) identifying policies $(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$ that satisfy $\{\bm{u}^{p^{*}}=\bm{K}^{p^{*}}\bm{x}_{\bm{w}}^{*}\}_{\bm{w}\in\bm{\mathcal{W}}}$, $p\in\mathcal{P}$.
-Highlighting this equivalence, we refer to such $\bm{u}^{*}$ as the open-loop realisation of the closed-loop policy $\bm{K}^{*}$, and vice-versa.
+From Assumption 2, we have that $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}\neq\emptyset$ when $\bm{\mathcal{C}}=\bm{\mathcal{U}}^{\bm{\mathcal{X}}}$.
+In the case of $\bm{\mathcal{C}}\subseteq\mathcal{L}(\ell_{\infty}^{N_{x}},\ell_{\infty}^{N_{u}})$, establishing the existence (and, especially, uniqueness) of a solution is demanding [39].
+In practice, the set $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ could be constructed from open-loop equilibria $\bm{u}^{\star}\in\Omega_{\mathcal{G}_{\infty}}$ by i) parametrising the set of all possible trajectories $\{\bm{x}_{\bm{w}}^{\star}=\bm{F_{w}}\bm{u}^{\star}\}_{\bm{w}\in\bm{\mathcal{W}}}$, then ii) identifying policies $(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$ that satisfy $\{\bm{u}^{p^{\star}}=\bm{K}^{p^{\star}}\bm{x}_{\bm{w}}^{\star}\}_{\bm{w}\in\bm{\mathcal{W}}}$, $p\in\mathcal{P}$.
+Highlighting this equivalence, we refer to such $\bm{u}^{\star}$ as an open-loop realisation of the closed-loop policy $\bm{K}^{\star}$, and vice-versa.
 Best-response dynamics for GFNE seeking
@@
 is the best-response of $p\in\mathcal{P}$ to other players’ plan of action.
-Under its feedback representation, $\bm{u}^{p}=\bm{K}^{p}\bm{x}\in BR^{p}(\bm{u}^{-p})$ is a solution to the infinite-horizon control problem
+Under its feedback representation, $\bm{u}^{p}=\bm{K}^{p}\bm{x}\in BR^{p}(\bm{u}^{-p})$ is a solution to the infinite-horizon stochastic control problem
 $\underset{\bm{u}^{p}\coloneqq\bm{K}^{p}\bm{x}}{\text{minimize}}\quad$ $\mathrm{E}\left[\sum_{t=0}^{\infty}L^{p}(x_{t},u^{p}_{t},u^{-p}_{t})\right]$ (17a)
 $\underset{\forall t\in\mathbb{N}}{\text{subject to}}\quad$ $x_{t{+}1}=Ax_{t}+\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}u^{\tilde{p}}_{t}+w_{t},$ (17b)
-$x_{t}\in\mathcal{X},\leavevmode\nobreak\ \leavevmode\nobreak\ u^{p}_{t}\in\mathcal{U}^{p},\leavevmode\nobreak\ \leavevmode\nobreak\ (u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},$ (19c)
+$x_{t}\in\mathcal{X},~{}~{}u^{p}_{t}\in\mathcal{U}^{p},~{}~{}(u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},$ (17c)
 $\bm{K}^{p}\in\bm{\mathcal{C}}^{p},$ (17d)
@@
 While posed in terms of action signals ($\bm{u}^{p}$, $p\in\mathcal{P}$), Problem (17) should be interpreted as the direct search for a best-response policy $\bm{K}^{p}$ against the (fixed) plan of action from other players, $\bm{u}^{-p}\coloneqq(\bm{u}^{\tilde{p}})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$.
-We slightly abuse notation and let $BR^{p}(\bm{K}^{-p})$ be its solutions, when the problem is parametrised by $\bm{u}^{-p}\coloneqq\bm{K}^{-p}\bm{x}=(\bm{K}^{\tilde{p}}\bm{x})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$.
-The mapping $BR:\bm{\mathcal{C}}\rightrightarrows\bm{\mathcal{C}}$, defined by $BR(\bm{K})=BR^{1}(\bm{K}^{-1})\times\cdots\times BR^{N_{P}}(\bm{K}^{-N_{P}})$, is the joint best-response to a strategy profile $\bm{K}$.
-The GFNEs of $\mathcal{G}_{\infty}$ are thus the fixed-points of this mapping: That is, $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}=\{\bm{K}^{*}\in\bm{\mathcal{C}}:\bm{K}^{*}\in BR(\bm{K}^{*})\}$.
+We slightly abuse notation and let $BR^{p}(\bm{K}^{-p})$ be its solutions when parametrised by $\bm{u}^{-p}\coloneqq\bm{K}^{-p}\bm{x}=(\bm{K}^{\tilde{p}}\bm{x})_{\tilde{p}\in\mathcal{P}\backslash\{p\}}$.
+The mapping $BR:\bm{\mathcal{C}}\rightrightarrows\bm{\mathcal{C}}$, defined by $BR(\bm{K})=BR^{1}(\bm{K}^{-1})\times\cdots\times BR^{N_{P}}(\bm{K}^{-N_{P}})$, is the joint best-response to a strategy profile $\bm{K}\in\bm{\mathcal{C}}$.
+The GFNE of $\mathcal{G}_{\infty}$ thus correspond to the fixed-points of this mapping: That is, $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}=\{\bm{K}^{\star}\in\bm{\mathcal{C}}:\bm{K}^{\star}\in BR(\bm{K}^{\star})\}$.
 Due to constraints ($\mathcal{X},\mathcal{U}^{p},\mathcal{U}_{\mathcal{G}}$) and $\bm{\mathcal{C}}^{p}$, an analytical solution to Problem (17) does not exist.
 Moreover, because infinite-dimensional, its numerical approximation cannot be obtained.
-The BRD method for GFNE seeking is given in Algorithm 3.
-As $\mathcal{G}_{\infty}$ is dynamic and stationary, the procedure needs not episodic repetitions of the game.
+A BRD for GFNE seeking is outlined in Algorithm 3.
+As $\mathcal{G}_{\infty}$ is dynamic and stationary, the procedure does not require episodic repetitions of the game.
 Instead, the learning dynamics occurs simultaneously with the game’s execution: Players learn and announce their new policies at stages $t\in\{(k{+}1)\Delta T\}_{k\in\mathbb{N}}$.
@@
 Input: Game $\mathcal{G}_{\infty}\coloneqq(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$
-Output: GFNE $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$
+Output: GFNE $\bm{K}^{\star}=(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$
 1
-2 Initialize $\bm{K}_{0}=(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k=0$;
+2 Initialize $\bm{K}_{0}\coloneqq(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k\coloneqq 0$;
 3 for $t=0,1,2,\ldots$ do
@@
 Algorithm 3 BRD for GFNE seeking (BRD-GFNE)
-If we verbally execute Algorithm 3, we have the following:
+A verbal execution of Algorithm 3 yields the following:
   • The players $p\in\mathcal{P}$ act on $\mathcal{G}_{\infty}$ according to the policies
@@
 The BRD-GFNE induces an operator $T=(1-\eta)I+\eta BR$ which is equivalent to the update rule of its static counterpart.
-Thus, it possesses the same properties: The learning dynamics converge if $BR$ is non-expansive and the convergence rate is geometric if $BR$ is also a contraction (Theorems 23).
-These properties can also be stated in terms of stage indices $t\in\mathbb{N}$ by substituting $k=\lfloor t/\Delta T\rfloor$.
+Thus, it possesses the same properties: The learning dynamics converge if $BR$ is nonexpansive and the convergence rate is geometric if $BR$ is also a contraction (Lemmas 12).
+These properties can also be stated in terms of stage indices $t\in\mathbb{N}$ by replacing $k=\lfloor t/\Delta T\rfloor$.
 As in the static case, a careful choice of the learning rate $\eta\in(0,1)$ is required to ensure that this fixed-point iteration is well-defined.
-Moreover, the BRD-GFNE can be interrupted at any episode $k_{f}>1$ (producing an $\epsilon$-GFNE, $\epsilon>0$) and, if necessary, a central coordinator assigned to project the resulting strategy profile $\bm{K}_{k_{f}}$ onto the feasible set $\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})$.
-Finally, we note that such considerations can be disregarded when the game is not generalised, that is, when $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$.
-III Best-response dynamics via system level synthesis
+Finally, we stress that the BRD-GFNE can be interrupted at any $k_{f}>1$, thus producing an $\epsilon$-GFNE policy $\bm{K}_{k_{f}}$ with associated equilibrium gap $\epsilon>0$.
+3 Best-response dynamics via system level synthesis
 In this section, we present an approach for GFNE seeking in (stationary) stochastic dynamic games.
@@
 Then, a modified BRD-GFNE procedure is proposed and its convergence properties are investigated.
-We focus on $N_{P}$-players linear-quadratic stochastic games $\mathcal{G}_{\infty}^{\text{LQ}}=(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$ with dynamics
+We focus on $N_{P}$-player linear-quadratic stochastic games $\mathcal{G}_{\infty}^{\text{LQ}}=(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$ with dynamics
 $x_{t{+}1}=Ax_{t}+\sum_{p\in\mathcal{P}}B^{p}u_{t}^{p}+w_{t},\quad x_{0}\text{ given},$ (18)
@@
 defined by matrices $C^{p}\in\mathbb{R}^{N_{z}\times N_{x}}$ and $D^{p\tilde{p}}\in\mathbb{R}^{N_{z}\times N_{u}^{\tilde{p}}}$ with dimension $N_{z}\geq N_{x}+N_{u}$.
-The following assumptions ensure that GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ are admissible (that is, stabilising):
+The following assumptions ensure that stabilising GFNE solutions to $\mathcal{G}_{\infty}^{\text{LQ}}$ exist:
 Assumption 3.
@@
 Moreover, $D^{p\tilde{p}^{{\mkern-1.5mu\mathsf{T}}}}C^{p}=0=C^{p^{{\mkern-1.5mu\mathsf{T}}}}D^{p\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$.
-Finally, the constraint sets $\mathcal{X}$ and $\mathcal{U}^{p}$ ($\forall p$) are convex polyhedra and contain their respective origins: $0\in\mathcal{X}$ and $0\in\mathcal{U}^{p}$.
-The class $\mathcal{G}_{\infty}^{\text{LQ}}$ describe problems in which $N_{P}$ non-cooperative agents have to agree on stationary policies that jointly stabilise a global system, robustly to the noise process, while penalising state- and input-deviations differently.
+Finally, the sets $\mathcal{X}$, $\mathcal{U}^{p}$ ($\forall p$), and $\mathcal{U}_{\mathcal{G}}$, are convex polyhedra satisfying $0\in\mathbf{relint}~{}\mathcal{X}$, $0\in\mathbf{relint}~{}\mathcal{U}^{p}$, and $0\in\mathbf{relint}~{}\mathcal{U}_{\mathcal{G}}$.
+The class $\mathcal{G}_{\infty}^{\text{LQ}}$ describe problems in which $N_{P}$ noncooperative agents have to agree on stationary policies that jointly stabilise a global system, robustly to the noise process, while penalising state- and input-deviations differently.
 While representative of many practically relevant problems, this choice is not restrictive.
 Our derivations should follow similarly for any collection of cost functions $\{L^{p}\}_{p\in\mathcal{P}}$ satisfying Assumption 2.
-III-A System-level best-response mappings
-System level synthesis (SLS, [36]) is a novel methodology for controller design which lifts the search for optimal controllers into the search of optimal closed-loop responses to disturbances.
-Under this approach, the synthesis of stationary control policies can be equivalently posed as the solution to scalable numerical problems, even when operational constraints on the state and control signals need to be enforced.
-Additionally, the method allows for constraints to be enforced directly on the structure of the control policies.
+3.1 System-level best-response mappings
+System level synthesis (SLS, [35]) is a novel methodology for controller design centred on the equivalent representation of control policies in terms of the closed-loop responses that they achieve.
+Unlike similar approaches, such as the Youla [40] and input-output (IOP, [41]) parametrisations, SLS allows the synthesis of state-feedback policies to be posed as the solution to convex optimisation problems, even when subjected to constraints on the state- and input-signals, and on the structure of the policy itself.
 In this section, we present a system-level parametrisation for the best-response mappings in $\mathcal{G}_{\infty}^{\text{LQ}}$.
@@
 the signals $(\bm{\hat{x}},\bm{\hat{u}}^{1},\ldots,\bm{\hat{u}}^{N_{P}})$ can be expressed in terms of $\bm{\hat{w}}$,
-$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}^{1}\\ \vdots\\ \bm{\hat{u}}^{N_{P}}\end{bmatrix}$ $=\begin{bmatrix}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\\ \bm{\hat{K}}^{1}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\\ \vdots\\ \bm{\hat{K}}^{N_{P}}(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}\end{bmatrix}\bm{\hat{w}};$
-$=\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}\bm{\hat{w}},$ (23)
-where the introduced transfer matrices ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}$) are denoted as system level responses or closed-loop maps.
+$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}^{1}\\ \vdots\\ \bm{\hat{u}}^{N_{P}}\end{bmatrix}$ $=\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}\bm{\hat{w}},$ (21)
+where $\bm{\hat{\Phi}}_{x}=(zI-A-\textstyle\sum_{p\in\mathcal{P}}B^{p}\bm{\hat{K}}^{p})^{-1}$ and $\bm{\hat{\Phi}}_{u}^{p}=\bm{\hat{K}}^{p}\bm{\hat{\Phi}}_{x}$ ($p\in\mathcal{P}$).
+The introduced transfer matrices ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}$) are referred to as system level responses or closed-loop maps.
 Under this representation, the following result holds.
-Theorem 4 (System level parametrisation, [36]).
+Theorem 1 (System level parametrisation).
 Consider the dynamics Eq. (20) under state-feedback $\bm{\hat{u}}^{p}=\bm{\hat{K}}^{p}\bm{\hat{x}}$ ($\forall p\in\mathcal{P}$). The following statements are true:
@@
     The affine space
-    $\begin{bmatrix}zI-A&{-}B^{1}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ {-}B^{N_{P}}\end{bmatrix}\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}=I,$ (24)
+    $\begin{bmatrix}zI-A&{-}B^{1}~{}\cdots~{}{-}B^{N_{P}}\end{bmatrix}\begin{bmatrix}\bm{\hat{\Phi}_{x}}\\ \bm{\hat{\Phi}_{u}}^{1}\\ \vdots\\ \bm{\hat{\Phi}_{u}}^{N_{P}}\end{bmatrix}=I,$ (22)
     with $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}}\in\frac{1}{z}\mathcal{RH}_{\infty}$, parametrizes all system responses from $\bm{\hat{w}}$ to $(\bm{\hat{x}},\bm{\hat{u}}^{1},\ldots,\bm{\hat{u}}^{N_{P}})$ achievable by internally stabilising policies $(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$.
@@
     $\bm{\hat{u}}^{p}$ $=\bm{\tilde{\Phi}_{u}}^{p}\bm{\hat{\xi}},$ (23b)
-    with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$ and $\bm{\tilde{\Phi}_{u}}^{p}=z\bm{\hat{\Phi}_{u}}^{p}$.
+    with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$ and $\bm{\tilde{\Phi}_{u}}^{p}=z\bm{\hat{\Phi}_{u}}^{p}$ (see Figure 2).
 Proof.
-To simplify the notation, we momentarily define the matrices $B\coloneqq[B^{1}\ B^{2}\ \cdots\ B^{N_{P}}]$, $\bm{\hat{\Phi}_{u}}\coloneqq\big{[}\bm{\hat{\Phi}_{u}}^{1^{{\mkern-1.5mu\mathsf{T}}}}\leavevmode\nobreak\ \bm{\hat{\Phi}_{u}}^{2^{{\mkern-1.5mu\mathsf{T}}}}\cdots\bm{\hat{\Phi}_{u}}^{N_{P}^{{\mkern-1.5mu\mathsf{T}}}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}$, and $\bm{\hat{K}}\coloneqq\big{[}\bm{\hat{K}}^{1^{\mkern-1.5mu\mathsf{T}}}\leavevmode\nobreak\ \bm{\hat{K}}^{2^{\mkern-1.5mu\mathsf{T}}}\cdots\bm{\hat{K}}^{N_{P}^{\mkern-1.5mu\mathsf{T}}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}$.
-For the first statement, consider internally stabilising policies $(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$ such that $\bm{\hat{u}}^{p}=\bm{\hat{K}}^{p}\bm{\hat{x}}$ ($\forall p\in\mathcal{P}$).
-Then, from Eq.
-(23), we have that
-$\begin{bmatrix}zI-A&-B\end{bmatrix}\begin{bmatrix}(zI-A-B\bm{\hat{K}})^{-1}\\ \bm{\hat{K}}(zI-A-B\bm{\hat{K}})^{-1}\end{bmatrix}$
-$\hskip 15.00002pt=(zI-A)(zI-A-B\bm{\hat{K}})^{-1}{-}B\bm{\hat{K}}(zI-A-B\bm{\hat{K}})^{-1}$
-$\hskip 15.00002pt=(zI-A-B\bm{\hat{K}})(zI-A-B\bm{\hat{K}})^{-1}$
-$\hskip 15.00002pt=I.$
-For the second statement, we first show that $\bm{\hat{K}}$ achieves the desired response, and then that these policies are internally stabilising.
-Note that Eq.
-(24) implies that $\bm{\hat{\Phi}_{x}}$ has the leading spectral component $\Phi_{x,1}=I_{N_{x}}$, which is invertible, and thus $\bm{\hat{\Phi}_{x}}^{-1}$ exists.
-Then, $\bm{\hat{K}}=\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1}$ is well-defined, and we have
-$\bm{\hat{x}}$ $=(zI-A-B\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1})^{-1}\bm{\hat{w}}$
-$=\bm{\hat{\Phi}_{x}}\left((zI-A)\bm{\hat{\Phi}_{x}}-B\bm{\hat{\Phi}_{u}}\right)^{-1}\bm{\hat{w}}$
-$=\bm{\hat{\Phi}_{x}}\bm{\hat{w}},$
-due to Eq.
-(24), and $\bm{\hat{u}}=\bm{\hat{K}}\bm{\hat{x}}=\bm{\hat{\Phi}_{u}}\bm{\hat{\Phi}_{x}}^{-1}\bm{\hat{\Phi}_{x}}\bm{\hat{w}}=\bm{\hat{\Phi}_{u}}\bm{\hat{w}}$.
-Thus, $\bm{\hat{K}}$ achieves the response ($\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}$), or, equivalently, $(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$ achieve $(\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}})$.
-To show that this policy is internally stabilising, consider its equivalent representation $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}$ with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$, $\bm{\tilde{\Phi}_{u}}=z\bm{\hat{\Phi}_{u}}$, and $\bm{\tilde{\Phi}_{y}}=I$, as depicted in Figure 2.
+Defining $B\coloneqq[B^{1}\ B^{2}\ \cdots\ B^{N_{P}}]$, and transfer matrices $\bm{\hat{\Phi}_{u}}\coloneqq\texttt{col}(\bm{\hat{\Phi}_{u}}^{1},\ldots,\bm{\hat{\Phi}_{u}}^{N_{P}})$ and $\bm{\hat{K}}\coloneqq\texttt{col}(\bm{\hat{K}}^{1},\ldots,\bm{\hat{K}}^{N_{P}})$, the proof is as in [35, Theorem 4.1].
+We refer to the Supplementary Material for the full details.
+In the second statement, we consider an alternative representation $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}$ with $\bm{\tilde{\Phi}_{x}}=z(I-\bm{\hat{\Phi}_{x}})$, $\bm{\tilde{\Phi}_{u}}=z\bm{\hat{\Phi}_{u}}$, and $\bm{\tilde{\Phi}_{y}}=I$: This leads to the transfer matrices from $(\bm{\hat{\delta}_{x}},\bm{\hat{\delta}_{u}},\bm{\hat{\delta}_{\xi}})$ to $(\bm{\hat{x}},\bm{\hat{u}},\bm{\hat{\xi}})$,
+$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}\\ \bm{\hat{\xi}}\end{bmatrix}=\begin{bmatrix}\bm{\hat{\Phi}_{x}}&\bm{\hat{\Phi}_{x}}B&\bm{\hat{\Phi}_{x}}(zI-A)\\ \bm{\hat{\Phi}_{u}}&I+\bm{\hat{\Phi}_{u}}B&\bm{\hat{\Phi}_{u}}(zI-A)\\ \frac{1}{z}I&\frac{1}{z}B&\frac{1}{z}(zI-A)\\ \end{bmatrix}\begin{bmatrix}\bm{\hat{\delta}_{x}}\\ \bm{\hat{\delta}_{u}}\\ \bm{\hat{\delta}_{\xi}}\end{bmatrix},$ (24)
+which are all stable due to $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}\in\frac{1}{z}\mathcal{RH}_{\infty}$.
+Thus, the policy $\bm{\hat{K}}=(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$ is internally stabilising.∎
 Figure 2: Feedback structure for the policy $\bm{\hat{K}}=\bm{\tilde{\Phi}_{u}}(zI-\bm{\tilde{\Phi}_{x}})^{-1}\bm{\tilde{\Phi}_{y}}=\bm{\hat{\Phi}_{u}^{p}}\bm{\hat{\Phi}_{x}^{-1}}$, equivalent to the internal representation in Eq.
 (23).
-Introducing external perturbations $\{\delta_{x},\delta_{u},\delta_{\xi}\}\subseteq\ell_{\infty}$ into the game, it suffices to verify that the transfer matrices from $(\bm{\hat{\delta}_{x}},\bm{\hat{\delta}_{u}},\bm{\hat{\delta}_{\xi}})$ to $(\bm{\hat{x}},\bm{\hat{u}},\bm{\hat{\xi}})$ are stable.
-In this case, we have
-$\begin{bmatrix}\bm{\hat{x}}\\ \bm{\hat{u}}\\ \bm{\hat{\xi}}\end{bmatrix}=\begin{bmatrix}\bm{\hat{\Phi}_{x}}&\bm{\hat{\Phi}_{x}}B&\bm{\hat{\Phi}_{x}}(zI-A)\\ \bm{\hat{\Phi}_{u}}&I+\bm{\hat{\Phi}_{u}}B&\bm{\hat{\Phi}_{u}}(zI-A)\\ \frac{1}{z}I&\frac{1}{z}B&\frac{1}{z}(zI-A)\\ \end{bmatrix}\begin{bmatrix}\bm{\hat{\delta}_{x}}\\ \bm{\hat{\delta}_{u}}\\ \bm{\hat{\delta}_{\xi}}\end{bmatrix}.$ (26)
-Since $\bm{\hat{\Phi}_{x}},\bm{\hat{\Phi}_{u}}\in\frac{1}{z}\mathcal{RH}_{\infty}$, the transfer matrices in Eq.
-(26) are all stable.
-We thus conclude that the policy $\bm{\hat{K}}$ is internally stabilising, and so are the policies $(\bm{\hat{K}}^{1},\bm{\hat{K}}^{2},\ldots,\bm{\hat{K}}^{N_{P}})$.
-∎
 We refer to the system level responses through their kernels, $\bm{\Phi_{x}}=(\Phi_{x,n})_{n\in\mathbb{N}}\in\ell_{2}(\mathbb{N})$ and $\bm{\Phi_{u}}^{p}=(\Phi^{p}_{u,n})_{n\in\mathbb{N}}\in\ell_{2}(\mathbb{N})$, for all $p\in\mathcal{P}$.
@@
 Corollary 1.1.
-A policy $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($\forall p\in\mathcal{P}$) is defined by the kernel $\bm{\Phi}^{p}=\bm{\Phi_{u}}^{p}*\bm{\Phi_{x}}^{-1}$, and implemented as
+A policy $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($p\in\mathcal{P}$) is defined by the kernel $\bm{\Phi}^{p}=\bm{\Phi_{u}}^{p}*\bm{\Phi_{x}}^{-1}$, and can be implemented as
 $\xi_{t}$ $=-\textstyle\sum_{\tau=1}^{t}\Phi_{x,\tau{+}1}\xi_{t-\tau}+x_{t};$ (25a)
@@
 The system level parametrisation enables a methodology for policy synthesis consisting of searching the space of stabilising policies (in)directly through $(\bm{\Phi_{x}},\bm{\Phi_{u}}^{p})$, $p\in\mathcal{P}$.
-In particular, this parametrisation can be leveraged to reformulate the best-response dynamics maps in $\mathcal{G}_{\infty}^{\text{LQ}}$ as tractable numerical programs.
-In this direction, consider that players design stabilising policies $\bm{K}=(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})$ by choosing their desired system level responses $\bm{\Phi_{u}}=(\bm{\Phi_{u}}^{1},\ldots,\bm{\Phi_{u}}^{N_{P}})$ simultaneously.
+In particular, this parametrisation can be leveraged to reformulate the best-response mappings in $\mathcal{G}_{\infty}^{\text{LQ}}$ as numerically tractable problems.
+In this direction, consider that players design stabilising policies $\bm{K}=(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})$ by choosing their desired system level responses $\bm{\Phi_{u}}=(\bm{\Phi_{u}}^{1},\ldots,\bm{\Phi_{u}}^{N_{P}})$, simultaneously.
 From the affine space Eq.
@@
 or $\bm{\Phi_{x}}=\bm{F_{\Phi}}\bm{\Phi_{u}}$ given the causal affine operator
-$\bm{F_{\Phi}}:\bm{\Phi_{u}}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\displaystyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{\Phi_{u}}^{p}+\bm{\delta}I_{N_{x}}\Big{)},$ (29)
-Using Parseval’s Theorem [42], we substitute Eq. (23) into Eq. (21) to redefine the objective functionals of $\mathcal{G}_{\infty}^{\text{LQ}}$ as
-$J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})$
-$\quad=\mathrm{E}\Big{[}\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}w_{n}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}w_{n}\|_{F}^{2}\Big{)}\Big{]}$
-$\quad=\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}.$
+$\bm{F_{\Phi}}:\bm{\Phi_{u}}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\textstyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{\Phi_{u}}^{p}+\bm{\delta}I_{N_{x}}\Big{)}.$ (27)
+Using the system level responses, the objective functionals of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be shown as equivalent to the functional
+$J^{p}(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})\\ =\sum_{n=1}^{\infty}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}.$
 The game $\mathcal{G}_{\infty}^{\text{LQ}}$ thus induces a system-level dynamic game,
@@
 Here, the set-valued mappings $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ are defined as
-$U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\coloneqq\{\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}:(\bm{\Phi_{u}}^{p},\bm{\Phi_{u}}^{-p})\in\bm{F_{\Phi}}^{\dagger}(\bm{\mathcal{C}}_{x}),\\ \bm{\Phi_{u}}^{p}*\bm{w}\in U^{p}(\bm{\Phi_{u}}^{-p}*\bm{w})\},$
+$U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})\coloneqq\{\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}:\bm{F_{\Phi}}\bm{\Phi_{u}}\in\bm{\mathcal{C}}_{x},\\ \quad\bm{\Phi_{u}}^{p}*\bm{w}\in U^{p}(\bm{\Phi_{u}}^{-p}*\bm{w})\},$
 which incorporate the constraints $U^{p}$ from the original $\mathcal{G}_{\infty}^{\text{LQ}}$.
@@
 $\Phi_{x,1}=I_{N_{x}}.$ (29g)
-Collectively, the mapping $BR_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$, defined by $BR_{\Phi}(\bm{\Phi_{u}})=BR^{1}_{\Phi}(\bm{\Phi_{u}}^{-1})\times\cdots\times BR^{N_{P}}_{\Phi}(\bm{\Phi_{u}}^{-N_{P}})$, is the joint best-response to a system-level strategy profile $\bm{\Phi_{u}}$.
-As before, the GNEs of $\mathcal{G}_{\infty}^{\Phi}$ are equivalent to the fixed-points of this map, $\Omega_{\mathcal{G}_{\infty}^{\Phi}}=\{\bm{\Phi_{u}}^{*}\in\bm{\mathcal{C}}_{u}:\bm{\Phi_{u}}^{*}\in BR_{\Phi}(\bm{\Phi_{u}}^{*})\}$.
+We refer to the Supplementary Material for a detailed derivation of Problem (29) from Problem (17).
+Collectively, the mapping $BR_{\Phi}(\bm{\Phi_{u}})=BR^{1}_{\Phi}(\bm{\Phi_{u}}^{-1})\times\cdots\times BR^{N_{P}}_{\Phi}(\bm{\Phi_{u}}^{-N_{P}})$, is the joint best-response to a system-level strategy profile $\bm{\Phi_{u}}$.
+The GNEs of $\mathcal{G}_{\infty}^{\Phi}$ are equivalent to the fixed-points of this map, $\Omega_{\mathcal{G}_{\infty}^{\Phi}}=\{\bm{\Phi_{u}}^{\star}\in\bm{\mathcal{C}}_{u}:\bm{\Phi_{u}}^{\star}\in BR_{\Phi}(\bm{\Phi_{u}}^{\star})\}$.
 Considering how $\mathcal{G}_{\infty}^{\text{LQ}}$ induces $\mathcal{G}_{\infty}^{\Phi}$, a relationship can be established between the best-responses $BR$ and $BR_{\Phi}$ and, consequently, between their fixed-points, $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$ and $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$.
-In Section III-B, we formalise this relationship and propose a learning dynamics for GFNE seeking based on the system-level best-response mappings.
-Remark 1.
-Problem (31) is independent of the matrix $\Sigma_{w}^{1\over 2}$ due to its linear-quadratic structure and the fact that $\Sigma_{w}\in\mathbb{S}^{N_{x}}$.
-Hence, we remove it from $J^{p}$ ($\forall p$) to simplify notation.
+In Section 3.2, we formalise this relationship and propose a learning dynamics for GFNE seeking based on the system-level best-response mappings.
 The best-responses $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ are still intractable: i) They are defined by infinite-dimensional problems with no general solution and ii) that require full knowledge of the noise process $(w_{n})_{n\in\mathbb{N}}$ to formulate the constraints $U_{\Phi}^{p}$.
-In the following, we tackle both issues and provide a class of finite-dimensional robust optimisation programs that approximate Problem (31).
+In the following, we tackle both issues and provide a class of finite-dimensional robust optimisation problems that approximate Problem (29).
 We conclude the section by presenting a class of system level constraints which enforce a richer feedback information pattern.
 Finite-horizon approximation
-The programs in $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ can be made finite-dimensional by restricting the closed-loop maps to the set of finite-impulse responses (FIR),
-$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}\in\ell_{2}[0,N]:\Phi_{x,n}\in\mathcal{C}_{x,n},\leavevmode\nobreak\ n{\in}[0,N),\leavevmode\nobreak\ \Phi_{x,N}=0\};$
-$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N):\Phi_{u,n}^{p}\in\mathcal{C}^{p}_{u,n},\leavevmode\nobreak\ n{\in}[0,N)\},$
-given horizon $N\in[2,\infty)$.
-We enforce $\bm{\Phi_{x}}\in\bm{\mathcal{C}}_{x}$ and $\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}$ in Problem (31) by adding the terminal constraint $\Phi_{x,N}=0$, and restricting the spectral factors to satisfy $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N-1}$.
-The constraint sets $\mathcal{C}_{x,n}\subseteq\mathbb{R}^{N_{x}\times N_{x}}$ and $\mathcal{C}_{u,n}^{p}\subseteq\mathbb{R}^{N_{u}\times N_{x}}$ ($\forall p$) have no restriction other than being compact convex sets.
-The resulting problems are of the form
-$\underset{\bm{\Phi_{u}}^{p}=(\Phi_{u,n}^{p})_{n=1}^{N-1}}{\text{minimize}}\quad$ $\sum_{n=1}^{N}\Big{(}\|C^{p}\Phi_{x,n}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\|_{F}^{2}\Big{)}$ (32a)
-$\underset{\forall n\in[1,N)}{\text{subject to}}\quad\leavevmode\nobreak\ \leavevmode\nobreak$ $\Phi_{x,{n+1}}=A\Phi_{x,n}+\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (32b)
-[…]
-$\Phi_{x,n}\in\mathcal{C}_{x,n},\quad\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (32f)
-$\Phi_{x,1}=I_{N_{x}},\quad\Phi_{x,N}=0,$ (32g)
-which define finite-dimensional convex programs that can be solved numerically.
-Since $(\bm{\mathcal{C}}_{x},\bm{\mathcal{C}}_{u}^{p})$ are finite-dimensional, the operators $(\bm{F_{\Phi}},\bm{K})$ and the convolutions in Eq.
-(32e) can be represented by matrix multiplications: Consider any kernel $\bm{\Phi}\in\{\bm{\Phi_{x}}\}\cup\{\bm{\Phi}^{p},\bm{\Phi_{u}}^{p}\}_{p\in\mathcal{P}}$ and some signal $\bm{z}\in\ell_{\infty}$.
-Then,
-$\bm{z}^{\text{out}}=\bm{\Phi}*\bm{z}\leavevmode\nobreak\ \leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ z^{\text{out}}=M_{\Phi}z,$
-given block-Toeplitz matrix $M_{\Phi}=[\Phi_{i-j}]_{i\in[1,2N+1],j\in[1,N+1]}$ and $(z,z^{\text{out}})$ being a vector representation of these signals, that is, $z=\texttt{col}(z_{0},\ldots,z_{N})$ and $z^{\text{out}}=\texttt{col}(z^{\text{out}}_{0},\ldots,z^{\text{out}}_{2N})$.
-Moreover, $z^{\text{out}}_{n}=(\Phi*z)_{n}=[M_{\Phi}]_{n}z$, where $[M_{\Phi}]_{n}$ is the $n$-th block-row of $M_{\Phi}$.
-Thus, Problem (32) can be reformulated into a program requiring only matrix algebra.
-Finally, we remark that policies ($\bm{K}^{p}$, $p\in\mathcal{P}$) are implemented either by using Corollary 4.1 or by directly computing kernels $\bm{\Phi}^{p}=([M_{\Phi_{u}^{p}}]_{n}\Phi_{x}^{-1})_{n\in\mathbb{N}}$.
-Although realising Problem (31) into a tractable program, the Problem (32) is only feasible when the pair $(A,B^{p})$ is full-state controllable.
-This is a difficult requirement in multi-agent settings, as often $N_{u}^{p}\ll N_{x}$ for all $p\in\mathcal{P}$, leading to Problem (32) becoming overdetermined.
+The programs in $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ can be made finite-dimensional by restricting the system level responses to the set of finite-impulse responses (FIR),
+$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}\in\ell_{2}[0,N]:\Phi_{x,n}\in\mathcal{C}_{x,n},~{}n{\in}[0,N),~{}\Phi_{x,N}=0\};$
+$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N):\Phi_{u,n}^{p}\in\mathcal{C}^{p}_{u,n},~{}n{\in}[0,N)\},$
+given horizon $N\in(1,\infty)$.
+We enforce $\bm{\Phi_{x}}\in\bm{\mathcal{C}}_{x}$ and $\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}$ in Problem (29) by adding a terminal constraint $\Phi_{x,N}=0$ then letting $\bm{\Phi_{x}}=(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $\bm{\Phi_{u}}=(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N-1}$.
+The sets $\mathcal{C}_{x,n}\subseteq\mathbb{R}^{N_{x}\times N_{x}}$ and $\mathcal{C}_{u,n}^{p}\subseteq\mathbb{R}^{N_{u}^{p}\times N_{x}}$ are required only to be compact convex sets.
+Under this condition, Problem (29) is finite-dimensional with $(N{-}1)N_{x}N_{u}^{p}$ decision variables (the entries of $\Phi_{u,1}^{p},\ldots,\Phi_{u,N{-}1}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}$) and thus can be solved numerically.
+We remark that the policies $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($p\in\mathcal{P}$) are still solutions to the infinite-horizon Problem (17) regardless of the closed-loop maps $\{\bm{\Phi_{x}},\bm{\Phi_{u}}^{p}\}$ being FIR.
+Although realising Problem (29) into a tractable program, the constraint $\Phi_{x,N}=0$ is only feasible when the pair $(A,B^{p})$ is full-state controllable.
+This is a difficult requirement in multi-agent settings, as often $N_{u}^{p}\ll N_{x}$ for all $p\in\mathcal{P}$, leading to overdetermined problems.
 Furthermore, enforcing FIR constraints is known to result in deadbeat policies: Control actions are excessively large in magnitude for small $N<\infty$.
 Alternatively, we restrict the system level responses to the sets
-$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}{\in}\ell_{2}[0,N]:\Phi_{x,n}{\in}\mathcal{C}_{x,n},\leavevmode\nobreak\ n{\in}[0,N),\leavevmode\nobreak\ \|\Phi_{x,N}\|_{F}^{2}\leq\gamma\};$
-$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}{\in}\ell_{2}[0,N):\Phi_{u,n}^{p}{\in}\mathcal{C}^{p}_{u,n},\leavevmode\nobreak\ n{\in}[0,N)\},$
+$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}{\in}\ell_{2}[0,N]:\Phi_{x,n}{\in}\mathcal{C}_{x,n},~{}n{\in}[0,N),~{}\|\Phi_{x,N}\|_{F}^{2}\leq\gamma\};$
+$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}{\in}\ell_{2}[0,N):\Phi_{u,n}^{p}{\in}\mathcal{C}^{p}_{u,n},~{}n{\in}[0,N)\},$
 with $\|\Phi_{x,N}\|_{F}^{2}=\sum_{i}\sigma_{i}(\Phi_{x,N})^{2}\leq\gamma$ for some factor $\gamma\in(0,1)$ and $\sigma_{i}(\cdot)$ denoting the $i$-th largest singular value of a matrix.
-These are reffered to as the set of approximately FIR for $N>0$.
-The $p$-th player best-response map thus takes the form
-$\underset{\bm{\Phi_{u}}^{p}=(\Phi_{u,n}^{p})_{n=1}^{N-1}}{\text{minimize}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{n=1}^{N-1}\Big{(}\|C^{p}\Phi_{x,n}\|_{F}^{2}{+}\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\|_{F}^{2}\Big{)}$
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \hskip 121.00024pt+\|C^{p}\Phi_{x,N}\|_{F}^{2}$ (33a)
-$\underset{\forall n\in[1,N)}{\text{subject to}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (33b)
+These are denoted as the set of (soft) FIR for $N>0$.
+The $p$-th player’s best-response map thus corresponds to the problem
+$\underset{\bm{\Phi_{u}}^{p}}{\text{minimize}}$ $~{}~{}\sum_{n=1}^{N-1}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}{+}\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}$
+$~{}~{}\hskip 124.00012pt+\|C^{p}\Phi_{x,N}\Sigma_{w}^{1\over 2}\|_{F}^{2}$ (30a)
+$\underset{\forall n\in[1,N)}{\text{subject to}}$ $~{}~{}\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (30b)
 […]
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,n}\in\mathcal{C}_{x,n},\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (33f)
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,1}=I_{N_{x}},\leavevmode\nobreak\ \leavevmode\nobreak\ \|\Phi_{x,N}\|_{F}^{2}\leq\gamma.$ (33g)
+$~{}~{}\Phi_{x,n}\in\mathcal{C}_{x,n},~{}~{}\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (30f)
+$~{}~{}\Phi_{x,1}=I_{N_{x}},~{}~{}\|\Phi_{x,N}\|_{F}^{2}\leq\gamma.$ (30g)
 The solutions to Problem (30) approximate those of the infinite-horizon Problem (29): With respect to $N$, the performance of the former converges to that achieved by the latter [35].
@@
 The map $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$, $\widehat{BR}_{\Phi}(\bm{\Phi_{u}})=\widehat{BR}_{\Phi}^{1}(\bm{\Phi_{u}}^{-1})\times\cdots\times\widehat{BR}_{\Phi}^{N_{P}}(\bm{\Phi_{u}}^{-N_{P}}),$ is the joint (approximately)best-response to the system-level profile $\bm{\Phi_{u}}$.
+The complexity of Problem (30) is agnostic to the number of players: The effect of other players’ strategies can always be condensed as affine terms (e.g., $z^{-p}_{n}=\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}B^{\tilde{p}}\Phi_{u,n}^{\tilde{p}}$).
+Conversely, it scales quickly with the FIR horizon $N$ and model dimensions $N_{x}$ and $N_{u}^{p}$.
+We remark, however, that this problem is still convex and can be solved efficiently by exploiting its structure using techniques from real-time optimisation of linear-quadratic control problems [42].
+Moreover, if the constraints Eq.
+(30e)–(30g) are column-separable (see [35]), then Problem (30) can be decomposed into smaller subproblems to be solved in parallel, substantially reducing its computational costs.
 Conversely to $BR_{\Phi}$, the fixed-points of $\widehat{BR}_{\Phi}$ do not coincide with the set of GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, but are rather contained in the set of $\varepsilon$-GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ for some equilibrium gap $\epsilon>0$.
@@
 with $\varepsilon=\max_{p\in\mathcal{P}}\gamma J^{p}(\bm{\Phi_{u}}^{p^{\varepsilon}},\bm{\Phi_{u}}^{-p^{\varepsilon}})$ for every player $p\in\mathcal{P}$.
-The equilibrium gap associated with $\bm{\Phi_{u}}^{\varepsilon}\in\widehat{BR}_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$ is thus proportional to the parameter $\gamma$, assuming that the terminal constraint Eq.
+The equilibrium gap associated with $\bm{\Phi_{u}}^{\varepsilon}\in\widehat{BR}_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$ thus depends on the choice of parameter $\gamma$, assuming that the terminal constraint Eq.
 (30g) also holds for solutions to the original Problem (29).
-Hereafter, $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ ($\forall p$) are assumed to incorporate the pair ($\bm{\mathcal{C}}_{x}$, $\bm{\mathcal{C}}_{u}^{p}$) defining the set of (soft-constrained) FIR approximations defined above for a $N>0$.
-We write $U_{\Phi,N}^{p}$ whenever this needs to be made explicit (e.g., to distinguish this choice from a more general $\bm{\mathcal{C}}_{x}$ and $\bm{\mathcal{C}}_{u}^{p}$).
+Hereafter, $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ ($\forall p$) are assumed to incorporate the sets ($\bm{\mathcal{C}}_{x}$, $\bm{\mathcal{C}}_{u}^{p}$) of (soft-constrained) FIR approximations as defined above for a $N>1$.
 Robust operational constraints
-From Assumption 3, the sets $\mathcal{X}$, $\mathcal{U}^{p}$, and $\mathcal{U}_{\mathcal{G}}$, can be expressed by linear inequalities,
-$\mathcal{X}$ =$=\{x_{t}\in\mathbb{R}^{N_{x}}:G_{x}x_{t}\leq g_{x}\};$
-$\mathcal{U}^{p}$ =$=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\leq g_{u}^{p}\};$
-$\mathcal{U}_{\mathcal{G}}$ =$=\{u_{t}\in\textstyle\prod_{\tilde{p}\in\mathcal{P}}\mathbb{R}^{N_{u}^{\tilde{p}}}:\textstyle\sum_{\tilde{p}\in\mathcal{P}}G_{\mathcal{G}}^{\tilde{p}}u^{\tilde{p}}_{t}\leq g_{\mathcal{G}}\},$
-given some matrices $G_{x}\in\mathbb{R}^{N_{\mathcal{X}}\times N_{x}}$, $G_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}^{p}}\times N_{u}^{p}}$ and $G_{\mathcal{G}}^{\tilde{p}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}\times N_{u}^{\tilde{p}}}$ ($\forall\tilde{p}\in\mathcal{P}$), and vectors $g_{x}\in\mathbb{R}^{N_{\mathcal{X}}}_{\geq 0}$, $g_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}}^{p}}_{\geq 0}$, and $g_{\mathcal{G}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}}_{\geq 0}$.
+From Assumption 3, the sets $\mathcal{X}$, $\mathcal{U}^{p}$ ($p\in\mathcal{P}$), and $\mathcal{U}_{\mathcal{G}}$, can be expressed by linear inequalities,
+$\mathcal{X}$ =$=\{x_{t}\in\mathbb{R}^{N_{x}}:G_{x}x_{t}\preceq\mathbf{1}_{N_{\mathcal{X}}}\};$
+$\mathcal{U}^{p}$ =$=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\preceq\mathbf{1}_{N_{\mathcal{U}}^{p}}\};$
+$\mathcal{U}_{\mathcal{G}}$ =$=\{u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:G_{\mathcal{G}}u_{t}\preceq\mathbf{1}_{N_{\mathcal{U}_{\mathcal{G}}}}\},$
+given some matrices $G_{x}\in\mathbb{R}^{N_{\mathcal{X}}\times N_{x}}$, $G_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}^{p}}\times N_{u}^{p}}$, and $G_{\mathcal{G}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}\times N_{u}}$.
 The map $U^{p}(u^{-p})$ is then equivalent to the actions $\bm{u}^{p}=\bm{\Phi_{u}}^{p}*\bm{w}$ whose associated response $\bm{\Phi_{u}}^{p}$ satisfy
-$[G_{x}]_{i}(\Phi_{x}*w)_{n}$ $\leq[g_{x}]_{i},\quad i=1,\ldots,N_{\mathcal{X}};$ (35)
-$[G_{u}^{p}]_{j}(\Phi_{u}^{p}*w)_{n}$ $\leq[g_{u}^{p}]_{j},\quad j=1,\ldots,N_{\mathcal{U}^{p}};$ (36)
-$\textstyle\sum_{\tilde{p}\in\mathcal{P}}[G_{\mathcal{G}}^{\tilde{p}}]_{l}(\Phi_{u}^{\tilde{p}}*w)_{n}$ $\leq[g_{\mathcal{G}}]_{l},\quad l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}},$ (37)
-with $([G_{x}]_{i},[g_{x}]_{i})$, $([G_{u}^{p}]_{j},[g_{u}^{p}]_{j})$, and $([G_{\mathcal{G}}^{\tilde{p}}]_{l},[g_{\mathcal{G}}]_{l})$ denoting, respectively, the $i$-th, $j$-th, and $l$-th rows of the corresponding matrices and vectors.
-In this work, players are assumed to synthesise policies that satisfy these constraints for any $\bm{w}\in\bm{\mathcal{W}}$.
-At the expense of conservativeness, we cast Problem (33) as a robust optimization problem by considering the worst-case realisation of the noise.
-Specifically, we reformulate Eq.
-(35),
-$\sum_{n^{\prime}=0}^{N}\sup_{w\in\mathcal{W}}[G_{x}]_{i}\Phi_{x,n^{\prime}}w\leq[g_{x}]_{i},\quad i=1,\ldots,N_{\mathcal{X}}$ (38)
-and similarly for Eqs. (36)–(37), then exploit our knowledge of $\bm{\mathcal{W}}$ to obtain an analytical solution for the supremum.
-A common instance of $\mathcal{G}_{\infty}^{\text{LQ}}$ considers the problem in which $(w_{n})_{n\in\mathbb{N}}$ is uniformly bounded.
-In these problems, the inequality Eq.
-(38) can be expressed in terms of a dual norm on the appropriate vector space.
-We highlight two important cases.
-  • $\mathcal{W}$ is an ellipsoid centred at zero. Formally, the set is
-    $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|Pw_{t}\|_{2}\leq 1\},$
-    given a matrix $P\in\mathbb{S}_{++}^{N_{x}}$.
-This corresponds to noise processes with uniformly bounded energy, as encoded by $\|\cdot\|_{2}$.
-For such cases, Eqs.
-(35)–(37) can be enforced by the second-order conic (SOC) constraints
-    $\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{x}]_{i},\quad(\forall i);$
-    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{u}^{p}]_{j},\quad(\forall j);$
-    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{\mathcal{G}}]_{l},\quad(\forall l).$
-  • $\mathcal{W}$ is a polyhedron, symmetric around zero. Formally,
-    $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|Pw_{t}\|_{\infty}\leq 1\},$
-    given a full-rank matrix $P\in\mathbb{R}^{N_{\mathcal{W}}\times N_{x}}$.
-This corresponds to noise processes with uniformly bounded intensity, as encoded by $\|\cdot\|_{\infty}$.
-For such cases, Eqs.
-(35)–(37) can be enforced by the first-order conic constraints
-    $\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{x}]_{i},\quad(\forall i);$
-    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{u}^{p}]_{j},\quad(\forall j);$
-    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{\mathcal{G}}]_{l},\quad(\forall l).$
+$[G_{x}]_{i}(\Phi_{x}*w)_{n}$ $\leq 1,\quad i=1,\ldots,N_{\mathcal{X}};$ (32)
+$[G_{u}^{p}]_{j}(\Phi_{u}^{p}*w)_{n}$ $\leq 1,\quad j=1,\ldots,N_{\mathcal{U}^{p}};$ (33)
+$[G_{\mathcal{G}}]_{l}(\Phi_{u}*w)_{n}$ $\leq 1,\quad l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}},$ (34)
+with $([G_{x}]_{i},[G_{u}^{p}]_{j},[G_{\mathcal{G}}]_{l})$ being the $i$-th, $j$-th, and $l$-th rows of the corresponding matrices.
+In this work, players are assumed to synthesise policies satisfying these constraints for any $\bm{w}\in\bm{\mathcal{W}}$.
+Equivalently, we cast Problem (30) as a robust optimization problem by considering the worst-case realisation of the noise.
+Specifically, we reformulate the constraint Eq.
+(32) as
+$\sup_{\bar{w}\in\mathcal{W}^{N}}\left\{\textstyle\sum_{n^{\prime}=0}^{N}[G_{x}]_{i}\Phi_{x,n^{\prime}}\bar{w}_{n^{\prime}}\right\}\leq 1,\quad i=1,\ldots,N_{\mathcal{X}},$
+and similarly for Eqs.
+(33)–(34).
+Since $(\bm{\Phi}_{x},\bm{\Phi}_{u}^{p})$ are FIR maps, it suffices to consider noise sequences of length $N$, $\bar{w}\in\mathcal{W}^{N}=\prod_{n^{\prime}=1}^{N}\mathcal{W}$, then exploit our knowledge of $\mathcal{W}$ to analytically solve this supremum.
+A common instance of $\mathcal{G}_{\infty}^{\text{LQ}}$ consists on the problems in which $(w_{n})_{n\in\mathbb{N}}$ is known to satisfy
+$w_{n}\in\mathcal{W}=\{P\zeta\in\mathbb{R}^{N_{x}}:\|\zeta\|_{q}\leq 1\},\quad\forall n\in\mathbb{N},$
+given a full-column-rank matrix $P\in\mathbb{R}^{N_{x}\times N_{\zeta}}$ and the $\ell_{q}$-norm $\|\cdot\|_{q}$.
+Using standard results from linear algebra [43], the robust counterpart of constraints Eqs.
+(32)–(34) are the constraints
+$N^{1/q}\big{\|}(I_{N}\otimes P^{{\mkern-1.5mu\mathsf{T}}})([G_{x}]_{i}\bar{\Phi}_{x})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{q}^{*}$ $\leq 1,\quad(\forall i);$ (35)
+$(N{-}1)^{1/q}\big{\|}(I_{N{-}1}\otimes P^{{\mkern-1.5mu\mathsf{T}}})([G_{u}^{p}]_{j}\bar{\Phi}_{u}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{q}^{*}$ $\leq 1,\quad(\forall j);$ (36)
+$(N{-}1)^{1/q}\big{\|}(I_{N{-}1}\otimes P^{{\mkern-1.5mu\mathsf{T}}})([G_{\mathcal{G}}]_{l}\bar{\Phi}_{u})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{q}^{*}$ $\leq 1,\quad(\forall l),$ (37)
+in which $\bar{\Phi}_{x}=[\Phi_{x,1}~{}\cdots~{}\Phi_{x,N}]$, $\bar{\Phi}_{u}^{p}=[\Phi_{u,1}^{p}~{}\cdots~{}\Phi_{u,N-1}^{p}]$, and $\bar{\Phi}_{u}=[\Phi_{u,1}~{}\cdots~{}\Phi_{u,N-1}]$.
+We refer to the Supplementary Material for a detailed derivation.
+The Problem (30) is thus rendered robust to the uncertainty in $\bm{w}\in\bm{\mathcal{W}}$ by incorporating the worst-case constraints Eqs.
+(35)–(37) in place of Eq.
+(30e).
 Remark 1.
-If $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}=\mathbb{R}^{N_{u}^{p}}$, and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, these constraints are trivially satisfied for any $w\in\bm{\mathcal{W}}$, and thus can be removed from Problem (31).
-Conversely, if $\mathcal{W}=\mathbb{R}^{N_{x}}$ when either $\mathcal{X}$, $\mathcal{U}^{p}$, or $\mathcal{U}_{\mathcal{G}}$ is bounded, then no stabilising policy can enforce those constraints for all possible $\bm{w}$.
+If $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}=\mathbb{R}^{N_{u}^{p}}$, or $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, then the corresponding constraints are trivially satisfied for any $w\in\bm{\mathcal{W}}$ and can be removed from Problem (29).
+Conversely, if $\mathcal{W}=\mathbb{R}^{N_{x}}$ when either $\mathcal{X}$, $\mathcal{U}^{p}$, or $\mathcal{U}_{\mathcal{G}}$ is bounded, then no stabilising policy can enforce those constraints for all $\bm{w}$.
+Remark 2.
+The constraints Eq.
+(35)–(37) are equivalent to those with $(-[G_{x}]_{i},-[G_{u}^{p}]_{j},-[G_{\mathcal{G}}]_{l})$.
+As such, matrices in the form $\widetilde{G}=\texttt{col}(G,-G)$ (i.e., enforcing $-\bm{1}_{N_{z}}\preceq Gz\preceq\bm{1}_{N_{z}}$) can be replaced by $G$ ($Gz\preceq\bm{1}_{N_{z}}$) and yield the same results.
 Structural constraints
 The sets $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ ($\forall p\in\mathcal{P}$) are designed to impose some structure directly on the policy $\bm{K}^{p}$ (Corollary 1.1), often in the form of sparsity constraints.
-An important class of such structural constraints encodes information patterns incurred by the presence of actuation and sensing delays: Let $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) satisfy the restrictions
-$\bm{\mathcal{C}}^{p}=\{\bm{K}^{p}\in\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}}^{p}):$ Actions [$\text{Actions }[B^{p}(K^{p}x)_{t}]_{i}\text{ respectively}$ (39)
-affect and feedback the component
-$[x_{t}]_{i}\text{ with the delays }d_{a},d_{s}>0\},$
-and consider the operators $\bm{S}_{x}:\bm{\Phi_{x}}\mapsto(S_{x,n}\odot\Phi_{x,n})_{n\in\mathbb{N}_{+}}$ and $\bm{S}_{u}^{p}:\bm{\Phi_{u}}^{p}\mapsto(S_{u,n}^{p}\odot\Phi_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ ($\forall p$), given the matrices
-$(S_{x,n})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}(A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{s}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
-$(S_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{s}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
-with $\texttt{Sp}(\cdot)$ denoting the sparsity pattern of a matrix [^1] The operator $\texttt{Sp}(A)$ produces a binary matrix such that $[\texttt{Sp}(A)]_{i,j}=1$ if $[A]_{i,j}\neq 0$ and $[\texttt{Sp}(A)]_{i,j}=0$ otherwise..
-It can be shown that $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) if its parametrisation satisfy $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n\in\mathbb{N}^{+}}$ and $(\Phi^{p}_{u,n}\in\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ with
+We consider the class of structural constraints which encode information patterns incurred by actuation and communication delays: Let $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) satisfy the restrictions
+$\bm{\mathcal{C}}^{p}=\{\bm{K}^{p}\in\mathcal{L}(\ell_{\infty}^{N_{x}},\ell_{\infty}^{N_{u}^{p}}):$ The state [$\text{The state }[x_{t}]_{i}\text{ is propagated to}$
+and affected by [$\text{and affected by }[B^{p}K^{p}x_{t}]_{i}\text{ with}$
+$\text{delays }d_{c},d_{a}>0\text{, respectively}\}.$
+and consider the operators $\bm{S}_{x}:\bm{\Phi_{x}}\mapsto(S_{x,n}\odot\Phi_{x,n})_{n\in\mathbb{N}_{+}}$ and $\bm{S}_{u}^{p}:\bm{\Phi_{u}}^{p}\mapsto(S_{u,n}^{p}\odot\Phi_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ ($\forall p\in\mathcal{P}$), given the signals
+$(S_{x,n})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}(A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{c}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}};$
+$(S_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{c}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
+with $\texttt{Sp}(\cdot)$ denoting the sparsity pattern of a matrix, that is, $[\texttt{Sp}(X)]_{i,j}=1$ if $[X]_{i,j}\neq 0$ and $[\texttt{Sp}(X)]_{i,j}=0$ otherwise, for any matrix $X$.
+It can be shown that $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$ if $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n\in\mathbb{N}^{+}}$ and $(\Phi^{p}_{u,n}\in\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ with convex sets
 $\mathcal{C}_{x,n}$ =$=\{\Phi_{x,n}\in\mathbb{R}^{N_{x}\times N_{x}}:\Phi_{x,n}=S_{x,n}\odot\Phi_{x,n}\};$ (38a)
@@
 (38) enforce that the closed-loop response to the noise obeys an information pattern induced by the dynamics of the game $\mathcal{G}_{\infty}^{\text{LQ}}$.
-Specifically, $[S_{x,n}]_{i,j}=0$ (respectively, $[S_{u,n}^{p}]_{i,j}=0$) indicates that any disturbance to the $j$-th component of the state, $[x_{t}]_{j}$, does not affect $[x_{t+n}]_{i}$ (does not feedback into action $[u_{t+n}^{p}]_{i}$).
-Finally, note that $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ are compact convex sets, as they correspond to the kernels of the linear operators ($I-\bm{S}_{x})$ and $(I-\bm{S}_{u}^{p}$).
+Specifically, $[S_{x,n}]_{i,j}=0$ (resp., $[S_{u,n}^{p}]_{i,j}=0$) implies that disturbances to the $j$-th component of the state, $[x_{t}]_{j}$, should not affect the state $[x_{t+n}]_{i}$ (the action $[u_{t+n}^{p}]_{i}$) when the policies $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ are employed.
+Enforcing sparsity also benefits the tractability of the best-response maps, as this reduces the effective number of decision variables.
 The ability to impose a desired policy structure using convex constraints is a central feature of the SLS framework.
 In the context of dynamic games, it allows for describing and, most importantly, solving problems where players have asymmetric information patterns; a major challenge for feedback Nash equilibrium problems [44, 39].
-For $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ from Eq.
-(40), conditions for the existence of GFNE (i.e., $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$), can be stated in terms of the delay parameters $d_{a}$ and $d_{s}$.
-Unless stated otherwise, the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ (and $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$) are assumed to include these constraints with $d_{a}=d_{s}=1$
-III-B System-level best-response dynamics
+Considering $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ from Eq.
+(38), the feasibility of the best-response maps $BR^{p}$ ($\forall p$) become dependent on the parameters $d_{a}$ and $d_{c}$: The larger their values the more restrictive the search space of matrices $\{\Phi_{x,n},\Phi_{u,n}^{p}\}_{n=1}^{N}$.
+As such, the actuation and communication delays associated with the game directly affect the existence of GFNE (that is, $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$).
+In practice, whether these structural constraints are consistent or not can be assessed by solving the feasibility problem associated with Problem (30) [43].
+A formal analysis of the requirements for $d_{a}$ and $d_{c}$ to ensure $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$ is beyond the scope of this paper.
+3.2 System-level best-response dynamics
 A learning dynamics based on the system-level best-responses $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ relies on the following central result.
 Theorem 3.
-A policy profile $\bm{K}^{*}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{*^{-1}}\in\bm{\mathcal{C}}$, is a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ if $\bm{\Phi_{u}}^{*}\in BR_{\Phi}(\bm{\Phi_{u}}^{*})$, or, equivalently,
-$\bm{\Phi_{u}}^{p^{*}}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{*}}),\quad\forall p\in\mathcal{P}.$ (41)
+A policy profile $\bm{K}^{\star}=(\bm{\Phi_{u}}^{1^{\star}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\star}})\bm{\Phi_{x}}^{\star^{-1}}\in\bm{\mathcal{C}}$, is a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ if $\bm{\Phi_{u}}^{\star}\in BR_{\Phi}(\bm{\Phi_{u}}^{\star})$, or, equivalently,
+$\bm{\Phi_{u}}^{p^{\star}}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{\star}}),\quad\forall p\in\mathcal{P}.$ (39)
 Proof.
-Consider an arbitrary fixed-point $\bm{\Phi_{u}}^{*}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{*})$.
-From Theorem 4, we have $\bm{\Phi_{x}}^{*}=\bm{F_{\Phi}}\bm{\Phi_{u}}^{*}$.
-Now, consider policies $\bm{K}^{p^{*}}=\bm{\Phi_{u}}^{p^{*}}(\bm{\Phi_{x}}^{*})^{-1}$, $p\in\mathcal{P}$.
-Clearly, $\bm{\Phi_{u}}^{p^{*}}=\bm{K}^{p^{*}}\bm{\Phi_{x}}^{*}$.
+Consider an arbitrary fixed-point $\bm{\Phi_{u}}^{\star}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{\star})$.
+From Theorem 1, we have $\bm{\Phi_{x}}^{\star}=\bm{F_{\Phi}}\bm{\Phi_{u}}^{\star}$.
+Now, consider policies $\bm{K}^{p^{\star}}=\bm{\Phi_{u}}^{p^{\star}}(\bm{\Phi_{x}}^{\star})^{-1}$, $p\in\mathcal{P}$.
+Clearly, $\bm{\Phi_{u}}^{p^{\star}}=\bm{K}^{p^{\star}}\bm{\Phi_{x}}^{\star}$.
 As a consequence, for any $\bm{w}\in\bm{\mathcal{W}}$,
-$\bm{\Phi_{u}}^{p^{*}}\bm{w}=\bm{K}^{p^{*}}\bm{\Phi_{x}}^{*}\bm{w}\leavevmode\nobreak\ \leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{u}^{p^{*}}=\bm{K}^{p^{*}}\bm{x}^{*}.$
-and, by definition, $\bm{\Phi_{u}}^{p^{*}}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{*}})$ imply $\bm{u}^{p^{*}}\in U^{p}(\bm{u}^{-p^{*}})$.
-Thus, $\bm{u}^{*}$ is the open-loop realisation of the policy $\bm{K}^{*}$.
-Finally, since $J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\cong J^{p}(\bm{\Phi_{u}}^{p^{*}},\bm{\Phi_{u}}^{-p^{*}})$ and $\bm{\Phi_{u}}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, we conclude that no player can obtain an admissible policy that unilaterally improves its cost, that is, $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$.
+$\bm{\Phi_{u}}^{p^{\star}}\bm{w}=\bm{K}^{p^{\star}}\bm{\Phi_{x}}^{\star}\bm{w}~{}~{}\iff~{}~{}\bm{u}^{p^{\star}}=\bm{K}^{p^{\star}}\bm{x}^{\star}.$
+and, by definition, $\bm{\Phi_{u}}^{p^{\star}}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{\star}})$ imply $\bm{u}^{p^{\star}}\in U^{p}(\bm{u}^{-p^{\star}})$.
+Thus, $\bm{u}^{\star}$ is an open-loop realisation of the policy $\bm{K}^{\star}$.
+Finally, since $J^{p}(\bm{u}^{p^{\star}},\bm{u}^{-p^{\star}})\cong J^{p}(\bm{\Phi_{u}}^{p^{\star}},\bm{\Phi_{u}}^{-p^{\star}})$ and $\bm{\Phi_{u}}^{\star}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, we conclude that no player can obtain an admissible policy that unilaterally improves its cost, that is, $\bm{K}^{\star}\in\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$.
 ∎
-The relationship between $BR$ and $BR_{\Phi}$ implies that a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be obtained analytically from a GNE of $\mathcal{G}_{\infty}^{\Phi}$.
-This fact allows us to adapt the BRD-GFNE procedure (Algorithm 3) and propose a procedure for GFNE seeking in constrained infinite-horizon dynamic games based on the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$.
-This system-level best-response dynamics (SLS-BRD) approach is given in Algorithm 4.
-We remark on some technical aspects:
-  • The pair $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{u,k}}^{p})$ defines the $p$-th player’s parametrisation after $k\in\mathbb{N}$ updates, that is, $\bm{\Phi_{x,k}}^{p}=(\Phi_{x,k,t}^{p})_{t\in\mathbb{N}}$ and $\bm{\Phi_{u,k}}^{p}=(\Phi_{u,k,t}^{p})_{t\in\mathbb{N}}$.
-The update index $k\in\mathbb{N}$ should not be mistaken for the stage index $t\in\mathbb{N}$.
-  • Updating the policy $\bm{K}_{k+1}^{p}$ consists of either employing Corollary 4.1 with maps $(\bm{\Phi_{x,k+1}}^{p},\bm{\Phi_{u,k+1}}^{p})$, or directly computing its kernel $\bm{\Phi_{k+1}}^{p}=([M_{\Phi_{u,k+1}^{p}}]_{n}\Phi_{x,k+1}^{-1})_{n\in\mathbb{N}}$.
-  • Responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ are most likely distinct at $k<\infty$, that is, $\bm{\Phi_{x,k}}^{p}\neq\bm{\Phi_{x,k}}^{\tilde{p}}$ for $p\neq\tilde{p}$.
-Consequently, the system level parametrisation (Eq.
-24) might not hold for the profile $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ and this could lead to stability issues.
-However, this policy still satisfies a robust variant of Theorem 4 when the distances $\|\bm{\Phi_{x,k}}^{p}-\bm{F_{\Phi}}\bm{\Phi_{u,k}}\|$ ($\forall p\in\mathcal{P}$) are sufficiently small [36].
 Input: Game $\mathcal{G}_{\infty}^{\text{LQ}}\coloneqq(\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}^{p}}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}})$
-Output: GFNE $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$
+Output: GFNE $\bm{K}^{\star}=(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$
 1
-2 Initialize $\bm{K}_{0}=(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k=0$;
+2 Initialize $\bm{K}_{0}\coloneqq(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k\coloneqq 0$;
 3 for $t=0,1,2,\ldots$ do
@@
 Algorithm 4 System-level BRD (SLS-BRD)
+The relationship between $BR$ and $BR_{\Phi}$ implies that a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be obtained analytically from a GNE of $\mathcal{G}_{\infty}^{\Phi}$.
+This allows us to adapt the BRD-GFNE procedure from Algorithm 3 into a procedure for GFNE seeking in constrained infinite-horizon dynamic games based on the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$.
+This system-level best-response dynamics (SLS-BRD) approach is given in Algorithm 4.
+We remark on some technical aspects:
+  • The pair $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{u,k}}^{p})$ is the parametrisation of $p$-th player’s policy after $k\in\mathbb{N}$ updates, that is, the signals $\bm{\Phi_{x,k}}^{p}=(\Phi_{x,k,t}^{p})_{t\in\mathbb{N}}$ and $\bm{\Phi_{u,k}}^{p}=(\Phi_{u,k,t}^{p})_{t\in\mathbb{N}}$.
+The index $k\in\mathbb{N}$ should not be mistaken for the stage index $t\in\mathbb{N}$.
+  • Updating the policy to $\bm{K}_{k+1}^{p}$ consists of employing the Corollary 1.1 with the updated maps $(\bm{\Phi_{x,k+1}}^{p},\bm{\Phi_{u,k+1}}^{p})$.
+  • Responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ are most likely distinct at $k<\infty$, that is, $\bm{\Phi_{x,k}}^{p}\neq\bm{\Phi_{x,k}}^{\tilde{p}}$ for $p\neq\tilde{p}$.
+Consequently, the system level parametrisation (Eq.
+22) might not hold for the profile $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ and this could lead to stability issues.
+However, this policy still satisfies a robust variant of Theorem 1 when the distances $\|\bm{\Phi_{x,k}}^{p}-\bm{F_{\Phi}}\bm{\Phi_{u,k}}\|$ ($\forall p\in\mathcal{P}$) are sufficiently small [35].
 The SLS-BRD induces an operator $T_{\Phi}=(1-\eta)I+\eta BR_{\Phi}$, which defines the global update $\bm{\Phi_{u,k+1}}$ from $\bm{\Phi_{u,k}}$.
 As before, $T_{\Phi}$ and $BR_{\Phi}$ share the same fixed-points: The GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$.
-From Theorem 6, convergence to a response $\bm{\Phi_{u}}^{*}\in T_{\Phi}(\bm{\Phi_{u}}^{*})$ then implies convergence to a policy $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$.
+From Theorem 3, convergence to a response $\bm{\Phi_{u}}^{\star}\in T_{\Phi}(\bm{\Phi_{u}}^{\star})$ then implies convergence to a policy $\bm{K}^{\star}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$.
 Hence, this learning dynamics is a formal procedure for GFNE seeking.
@@
 Convergence of the SLS-BRD
-The convergence of Algorithm 4 depends on $BR_{\Phi}$ (or $\widehat{BR}_{\Phi}$, for its tractable version) being at least non-expansive (Theorems 23). Formally,
-Theorem 7.
-Let the map $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$ be $L_{\widehat{BR}_{\Phi}}$-Lipschitz, with $L_{\widehat{BR}_{\Phi}}<1$.
-Then, the SLS-BRD $\bm{\Phi_{u,k+1}}=\widehat{T}_{\Phi}(\bm{\Phi_{u,k}})$ converge to the unique $\varepsilon$-GNE $\bm{\Phi_{u}}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ with rate
-$\cfrac{\|\bm{\Phi_{u,k}}-\bm{\Phi_{u}}^{*}\|_{\ell_{2}}}{\|\bm{\Phi_{u,0}}-\bm{\Phi_{u}}^{*}\|_{\ell_{2}}}\leq\big{(}(1{-}\eta)-\eta L_{\widehat{BR}_{\Phi}}\big{)}^{k}$ (42)
+The convergence of the Algorithm 4 depends on $BR_{\Phi}$ (or $\widehat{BR}_{\Phi}$, for its tractable version) being at least nonexpansive (Lemmas 12). Formally,
+Corollary 3.1.
+Let $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$ be $L_{\widehat{BR}_{\Phi}}$-Lipschitz, with $L_{\widehat{BR}_{\Phi}}<1$.
+Then, the SLS-BRD $\bm{\Phi_{u,k+1}}=\widehat{T}_{\Phi}(\bm{\Phi_{u,k}})$ converge to the unique $\varepsilon$-GNE $\bm{\Phi_{u}}^{\star}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ with rate
+$\cfrac{\|\bm{\Phi_{u,k}}-\bm{\Phi_{u}}^{\star}\|_{\ell_{2}}}{\|\bm{\Phi_{u,0}}-\bm{\Phi_{u}}^{\star}\|_{\ell_{2}}}\leq\big{(}(1{-}\eta)-\eta L_{\widehat{BR}_{\Phi}}\big{)}^{k}$ (40)
 from any feasible initial $\bm{\Phi_{u,0}}$.
@@
 Using this notation, Problem (30) can be reformulated as
-$\underset{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})}{\text{minimize}}\leavevmode\nobreak$ $\mathrm{Tr}\big{[}{\bm{\Phi_{u}}^{p}}^{{\mkern-1.5mu\mathsf{T}}}\bm{H}^{pp}\bm{\Phi_{u}}^{p}$
-$\hskip 40.00006pt+2(\textstyle\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\bm{H}^{p\tilde{p}}\bm{\Phi_{u}}^{\tilde{p}}+\bm{H}^{p0})^{{\mkern-1.5mu\mathsf{T}}}\bm{\Phi_{u}}^{p}\big{]}$ (44)
+$\underset{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})}{\text{minimize}}~{}$ $\mathrm{Tr}\big{[}{\bm{\Phi_{u}}^{p}}^{{\mkern-1.5mu\mathsf{T}}}\bm{H}^{pp}\bm{\Phi_{u}}^{p}$
+$\quad+2(\textstyle\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\bm{H}^{p\tilde{p}}\bm{\Phi_{u}}^{\tilde{p}}+\bm{H}^{p0})^{{\mkern-1.5mu\mathsf{T}}}\bm{\Phi_{u}}^{p}\big{]}$ (42)
 The maps $BR^{p}(\bm{\Phi_{u}}^{-p})$, $p\in\mathcal{P}$, thus correspond to the solution of quadratic programs with convex constraints $U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$, for $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$.
-In the case of $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$, and without structural constraints, that is, $\bm{S}_{x}=I$ and $\bm{S}_{u}^{p}=I$,
+In the case of $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$ and no structural constraints ($\bm{S}_{x}=\bm{S}_{u}^{p}=I$), we have
 $U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$ =$:\displaystyle=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N)$ : (43)
-$\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},\leavevmode\nobreak\ \Phi_{x,1}=I_{N_{x}},$
-$\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{x}]_{i},$
-$\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{u}^{p}]_{j},$
-$\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{\mathcal{G}}]_{l}\}.$
+$\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},~{}\Phi_{x,1}=I_{N_{x}},$
+$\big{\|}([G_{x}]_{i}\bar{\Phi}_{x})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq 1,~{}i=1,\ldots,N_{\mathcal{X}};$
+$\big{\|}([G_{u}^{p}]_{j}\bar{\Phi}_{u}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq 1,~{}j=1,\ldots,N_{\mathcal{U}^{p}};$
+$\big{\|}([G_{\mathcal{G}}]_{l}\bar{\Phi}_{u})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq 1,~{}l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}}\}.$
 Using standard techniques from optimisation, the constraints Eq.
 (43) can be incorporated into Problem (42) as linear inequalities.
-The solutions for these problems, that is, the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, are thus piecewise affine in $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$ [44].
-Consequently, also $\widehat{BR}_{\Phi}$ must be piecewise affine.
-A local Lipschitz constant can then be derived for each region of $\bm{\mathcal{C}}_{u}^{-p}$ that leads to a subset of the operational constraints being active.
-For non-generalised games (i.e., when $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$), the structure of this affine mapping can be exploited to derive a global Lipschitz constant for $\widehat{BR}_{\Phi}$.
+The best-responses mappings $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, are thus piecewise affine in $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$ [45].
+Consequently, also $\widehat{BR}_{\Phi}$ must be piecewise affine: A local Lipschitz constant can then be derived for each region of $\bm{\mathcal{C}}_{u}^{-p}$ that leads to a subset of the operational constraints being active.
+For non-generalised games, this fact can be exploited to derive a global Lipschitz constant for $\widehat{BR}_{\Phi}$.
 Theorem 4.
-Consider $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$. Then, The map $\widehat{BR}_{\Phi}$ is $L_{\widehat{BR}_{\Phi}}$-Lipschitz with
-$L_{\widehat{BR}_{\Phi}}=\sqrt{\textstyle\sum_{p\in\mathcal{P}}(L_{\widehat{BR}_{\Phi}}^{p})^{2}},$ (46)
-given the player-specific constants
-$L_{\widehat{BR}_{\Phi}}^{p}=\big{(}1+\kappa(\bm{H}^{pp})\big{)}\|(\bm{H}^{pp})^{\dagger}\bm{H}^{p,-p}\|_{2\to 2},$ (47)
-with condition number $\kappa(\bm{H}^{pp})=\|(\bm{H}^{pp})^{\dagger}\|_{2\to 2}\|\bm{H}^{pp}\|_{2\to 2}$.
-The proof of Theorem 8 is extensive: The reader is referred to Appendix -C for the full details.
-Importantly, this Lipschitz constant is not tight and thus $L_{\widehat{BR}_{\Phi}}<1$ using Eq.
-(46) is only a sufficient (but not necessary) condition for Theorem 7 to hold in practice.
+Consider $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, and $\mathcal{U}^{p}=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\preceq\mathbf{1}_{N_{\mathcal{U}}^{p}}\}$ with $G_{u}^{p}$ full-row-rank.
+Then, the best-response map $\widehat{BR}_{\Phi}$ is $L_{\widehat{BR}_{\Phi}}$-Lipschitz with
+$L_{\widehat{BR}_{\Phi}}^{2}=\textstyle\sum_{p\in\mathcal{P}}(L_{\widehat{BR}_{\Phi}}^{p})^{2},$ (44)
+given the player-specific $L_{\widehat{BR}_{\Phi}}^{p}=\sigma_{\max}(\bm{H}^{p,-p})/\sigma_{\min}(\bm{H}^{pp})$.
+The proof of Theorem 4 is extensive: The reader is referred to the Supplementary Material for the full details.
+The constants $L_{\widehat{BR}_{\Phi}}^{p}$ ($p\in\mathcal{P}$) quantify the ability of each player to optimize its objective in face of its rivals’ interference.
+Importantly, this Lipschitz constant is not tight and $L_{\widehat{BR}_{\Phi}}<1$ using Eq.
+(44) is only a sufficient (but not necessary) condition for Corollary 3.1 to hold in practice.
 Regardless, it highlights some intuitive, but non-trivial, facts about the convergence of the SLS-BRD:
-  • The condition $L_{\widehat{BR}_{\Phi}}<1$ is implies that the block-operator $\bm{H}=[H^{p\tilde{p}}]_{p,\tilde{p}\in\mathcal{P}}$ is diagonally dominant.
-This highlights the relationship between the SLS-BRD and the class of Jacobi iterative methods, where diagonal dominance of the linear system being solved is a strict requirement.
-  • The convergence rate of the SLS-BRD is inversely proportional to $\|D^{pp}\|_{2}^{2}$ ($\forall p$) through $\bm{H}^{pp}$.
+  • The condition $L_{\widehat{BR}_{\Phi}}<1$ is satisfied if the block-operator $\bm{H}=[H^{p\tilde{p}}]_{p,\tilde{p}\in\mathcal{P}}$ is diagonally dominant.
+This highlights the relationship between the SLS-BRD and classical Jacobi iterative methods, where diagonal dominance of the linear system being solved is a known requirement.
+  • The convergence rates of the SLS-BRD depend directly on $\|D^{pp}\|_{2}$ ($\forall p$) through $\sigma_{\min}(\bm{H}^{pp})=\sigma_{\min}^{2}(\bm{C}^{p}\bm{F_{\Phi}}^{p}+\bm{D}^{pp})$.
 As such, faster convergence is expected for games in which players apply strong penalties to their own actions.
-  • The convergence rate of the SLS-BRD depends on the number of players: Let $L_{\widehat{BR}_{\Phi}}^{p}=\sqrt{\alpha+\beta^{p}}$ for all $p\in\mathcal{P}$ given some $\alpha,\beta^{p}>0$.
-Then, $L_{\widehat{BR}_{\Phi}}^{2}=(\alpha N_{p})^{2}\sum_{p}\beta^{p}$.
-In large-scale games, players might need to become more conservative in order to ensure that $\widehat{BR}$ is a contraction.
-  • The convergence rate of the SLS-BRD is dominated by the slowest player: Whenever there exists a $p\in\mathcal{P}$ such that $L_{\widehat{BR}_{\Phi}}^{p}\gg L_{\widehat{BR}_{\Phi}}^{\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$, then $L_{\widehat{BR}_{\Phi}}\approx L_{\widehat{BR}_{\Phi}}^{p}$.
-  • The norms $\|\bm{H}^{pp}\|_{2\to 2}=\|\bm{C}^{p}\bm{F_{\Phi}}^{p}+\bm{D}^{pp}\|_{2\to 2}^{2}$ increase with the FIR horizon $N$, specially for unstable games.
-In some cases, players might need to proportionally decrease the state-penalties $C^{p}$ to ensure (or improve) convergence.
-At the cost of interpretability, similar Lipschitz constants as in Eqs.
-(46)–(47) can be obtained also when $(\mathcal{X},\mathcal{U}_{\mathcal{G}},\mathcal{U}^{p})$ and $P$ are general, and when structural constraints are present.
-Moreover, we remark that Theorem 8 is obtained under the assumption that $\|\Phi_{x,N}\|_{F}^{2}<\gamma$ holds strictly.
-The best-response maps $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ when this constraint is active, or when $\mathcal{W}$ is an ellipsoid, become quadratically-constrained quadratic programs and their Lipschitz properties are less intuitive.
-IV Examples
+  • The convergence rates of the SLS-BRD depend on the number of players since $L_{\widehat{BR}_{\Phi}}^{p}>0$ for all $p\in\mathcal{P}$.
+In large-scale games, players might need to apply strong penalties to their actions (through $D^{pp}$) to ensure convergence.
+  • The convergence rate of the SLS-BRD can be dominated by the slowest player: Whenever there exists a $p\in\mathcal{P}$ such that $L_{\widehat{BR}_{\Phi}}^{p}\gg L_{\widehat{BR}_{\Phi}}^{\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$, then $L_{\widehat{BR}_{\Phi}}\approx L_{\widehat{BR}_{\Phi}}^{p}$.
+At the cost of interpretability, similar Lipschitz constants as in Eq.
+(44) can also be obtained for general $(\mathcal{X},\mathcal{U}_{\mathcal{G}},\mathcal{U}^{p})$ and $P$, and when structural constraints are present.
+We stress that Theorem 4 is obtained under the assumption that $\|\Phi_{x,N}\|_{F}^{2}<\gamma$ holds strictly.
+The best-response maps $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ when this constraint is active, or when $\mathcal{W}$ is an ellipsoid, are solutions to quadratically-constrained quadratic programs (QCQP) and their Lipschitz properties are less intuitive.
+4 Examples
 In the following, we demonstrate the SLS-BRD algorithm in two exemplary problems: Decentralised control of an unstable network and management of a competitive market.
 In both problems, we set the initial profile $\bm{K}_{0}=(\bm{\Phi_{u,0}}^{1},\ldots,\bm{\Phi_{u,0}}^{N_{P}})\bm{\Phi_{x,0}}^{-1}$ by projecting the zero-response $\bm{\hat{\Phi}_{u,0}}=\bm{0}$ into the feasible set.
-[^2] That is, $\bm{\Phi_{u,0}}=\operatorname*{arg\,min}_{\bm{\Phi_{u}}\in\bm{\mathcal{U}}_{\Phi}}\|\bm{\Phi_{u}}-\bm{\hat{\Phi}_{u,0}}\|_{\ell_{2}}^{2}$ given the feasible set $\bm{\mathcal{U}}_{\Phi}=\{\bm{\Phi_{u}}\in\ell_{2}[0,N):\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p}),\leavevmode\nobreak\ p\in\mathcal{P}\}$.
 A SLS-BRD routine is then simulated by having players update their policies through best-responses to the parametrisation of their rivals’ policies, assumed to be available.
 Due to numerical limitations, we interrupt the updates whenever the condition $\|\bm{\Phi}_{u,k}^{p}-\bm{\Phi}_{u,k-1}^{p}\|_{\ell_{2}}/\|\bm{\Phi}_{u,k}^{p}\|_{\ell_{2}}\leq 10^{-8}$ ($\forall p\in\mathcal{P}$) is satisfied.
-IV-A Stabilisation of a bidirectional chain network
+4.1 Stabilisation of a bidirectional chain network
 Consider a game $\mathcal{G}_{\infty}^{\text{LQ}}=\{\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}\}$ with players $\mathcal{P}=\{1,2,3\}$ operating a chain network of $N_{x}=14$ single-state nodes whose dynamics are described by
-$A=\begin{bmatrix}1&0.2&&\\ -0.2&\ddots&\ddots&\\ &\ddots&\ddots&0.2\\ &&-0.2&1\end{bmatrix};\leavevmode\nobreak\ \left\{B^{p}{=}\begin{bmatrix}\bm{0}_{6(p{-}1)\times 2}\\[6.45831pt] I_{2}\\[2.15277pt] \bm{0}_{6(3{-}p)\times 2}\end{bmatrix}\right\}_{p\in\mathcal{P}},$ (48)
+$A=\begin{bmatrix}1&0.2&&\\ -0.2&\ddots&\ddots&\\ &\ddots&\ddots&0.2\\ &&-0.2&1\end{bmatrix};~{}\left\{B^{p}{=}\begin{bmatrix}\bm{0}_{6(p{-}1)\times 2}\\[6.45831pt] I_{2}\\[2.15277pt] \bm{0}_{6(3{-}p)\times 2}\end{bmatrix}\right\}_{p\in\mathcal{P}},$
 where $A\in\mathbb{R}^{N_{x}\times N_{x}}$ and $B^{p}\in\mathbb{R}^{N_{x}\times N_{u}^{p}}$ with $N_{u}^{p}=2$ ($\forall p$).
-This game has unstable dynamics, since $\rho(A)=1.073>1$, however it is stabilisable for each $(A,B^{p})$.
+This game has unstable dynamics, since $\|A\|_{2}=1.073>1$, however it is stabilisable for each $(A,B^{p})$.
 Moreover, the game is subjected to a noise process described by $w_{t}\sim\mathrm{Uniform}(\mathcal{W})$, $t\in\mathbb{N}$, defined over $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$.
 In this problem, players are interested in stabilising the game $\mathcal{G}_{\infty}^{\text{LQ}}$, while minimising their individual objective functionals,
-$J^{p}(u^{p},u^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
-[…]
+$J^{p}(\bm{u}^{p},\bm{u}^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
+equivalent to Eq.
+(19) after setting $C^{p}=[I_{N_{x}}~{}\bm{0}_{N_{x}{\times}N_{u}}]^{{\mkern-1.5mu\mathsf{T}}}$, $D^{pp}=[\bm{0}_{N_{u}{\times}N_{x}}~{}\sqrt{\beta^{p}}I_{Nu}]^{{\mkern-1.5mu\mathsf{T}}}$, and $D^{p\tilde{p}}=0$ for all $\tilde{p}\in\mathcal{P}\backslash\{p\}$.
+The players’ actions are subjected to operational constraints, $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$, defined by the constraint sets
 $\mathcal{X}$ $=\mathbb{R}^{N_{x}};$
-[…]
+$\mathcal{U}^{p}$ =$=\big{\{}u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:\begin{bmatrix}\tfrac{1}{12}I_{N_{u}^{p}}\end{bmatrix}u^{p}_{t}\preceq\bm{1}_{N_{u}^{p}}\big{\}};$
 $\mathcal{U}_{\mathcal{G}}$ $=\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}},$
-enforcing $-10\leq[u_{t}^{p}]_{1}+[u_{t}^{p}]_{2}\leq 10$ for every $t\in\mathbb{N}$ and $p\in\mathcal{P}$.
+enforcing $\|u_{t}^{p}\|_{\infty}\leq 12$ for each $t\in\mathbb{N}$ and $p\in\mathcal{P}$ (Remark 2).
 We assume that players design their state-feedback policies, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$, considering a FIR horizon of $N=50$ and the constraints $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N}$,
 $\mathcal{C}_{x,n}$ =$=\{\Phi_{x,n}\in\mathbb{R}^{N_{x}\times N_{x}}:\Phi_{x,n}=\texttt{Sp}(A^{n-1})\odot\Phi_{x,n}\};$
-$\mathcal{C}_{u,n}^{p}$ =$=\{\Phi_{u,n}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}:\Phi_{u,n}^{p}=\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{n-1})\odot\Phi_{u,n}^{p}\}.$
+$\mathcal{C}_{u,n}^{p}$ =$=\{\Phi_{u,n}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}:\Phi_{u,n}^{p}=\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{n-1})\odot\Phi_{u,n}^{p}\},$
+defined by setting the delay parameters $d_{a}=d_{c}=1$.
 Under this setup, $\mathcal{G}_{\infty}^{\text{LQ}}$ belongs to the class of dynamic potential games (DPG, [46]) and the (unique) GFNE can be obtained in advance by solving a centralised optimisation problem.
-In this experiment, we simulate a GFNE seeking procedure for each different value $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$.
+In this experiment, we simulate a GFNE seeking procedure for each value $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$.
 The game $\mathcal{G}_{\infty}^{\text{LQ}}$ is executed alongside the updating of players’ policies according to some learning dynamics.
-The players seek $\varepsilon$-GFNE policies by adhering to the SLS-BRD routine (Algorithm 4) using their (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, with $\gamma=0.95$.
-The policies are updated simultaneously every $\Delta T=1$ stage with a learning rate of $\eta=1/2$.
-The convergence of the SLS-BRD routine to the fixed-point $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 3.
-The results demonstrate that the players’ policies are sufficiently close to the $\epsilon$-GFNE profile after $260$, $420$, and $510$ iterations for each respective weighting configuration.
-In each case, the (soft) FIR constraints are satisfied strictly after the initial profile (that is, $\|\Phi_{x,k,N}^{p}\|_{F}^{2}<0.95$, $k>2$, $p\in\mathcal{P}$).
-The Lipschitz properties of the global update rule $T_{\Phi}$ seem to follow the intuition provided in Section III-B: The convergence rate improves when players apply stronger penalties to their actions ($\beta=(10,40,10)$) when compared to that obtained by weaker control penalties ($\beta=(0.4,1.6,0.4)$).
-However, we stress that the Lipschitz constants from Theorem 8 do not consider structural constraints ($\bm{S}_{x}$ and $\bm{S}_{u}^{p}$, $\forall p\in\mathcal{P}$) and cannot be applied to this experiment.
-Regardless, the convergence rates are shown to be geometric, indicating that the best-response maps $\widehat{BR}_{\Phi}$ in these scenarios are contractive.
-If not interrupted, and disregarding numerical limitations, the SLS-BRD should continue to approach the fixed-point $\bm{K}^{*}$ at this rate.
+The players are assumed to seek $\varepsilon$-GFNE policies by adhering to the SLS-BRD routine (Algorithm 4) using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, with $\gamma=0.95$.
+The policies are updated simultaneously every $\Delta T=1$ stage with learning rate $\eta=1/2$.
+The convergence of the SLS-BRD routine to the fixed-point $\bm{K}^{\star}=\bm{\Phi_{u}}^{\star}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{\star}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\star}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 3.
+The results demonstrate that the players’ policies are sufficiently close to the $\epsilon$-GFNE profile after $260$, $370$, and $425$ iterations for each respective weighting configuration.
+In each case, the (soft) FIR constraints are satisfied strictly after the initial profile (that is, $\|\Phi_{x,k,N}^{p}\|_{F}^{2}<0.95$, $k>2$, $\forall p\in\mathcal{P}$).
+Moreover, this fixed-point iteration seems to follow the behaviour discussed in Section 3.2: The convergence improves when players apply stronger penalties to their actions ($\beta=(10,40,10)$) when compared to that obtained by weaker control penalties ($\beta=(0.4,1.6,0.4)$).
+However, we stress that the Lipschitz constants from Theorem 4 do not consider structural constraints ($\bm{S}_{x}$, $\bm{S}_{u}^{p}$) and thus cannot be applied to this experiment.
+Regardless, the convergence is shown to be geometric, indicating that the best-response maps $\widehat{BR}_{\Phi}$ are contractive.
+If not interrupted, and disregarding numerical limitations, the SLS-BRD should continue to approach the fixed-point $\bm{K}^{\star}$ at this rate.
 Figure 3: $N_{P}$-chain game: Convergence of the SLS-BRD routine.
+Figure 4: $N_{P}$-chain game: Relative distance between the local updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p})$, top row, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom row.
 Figure 5: $N_{P}$-chain game, $t\in(550,600]$: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for each execution of $\mathcal{G}_{\infty}^{\text{LQ}}$ with $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$.
@@
 As shown in Figure 4 (bottom), these distances decrease at a similar rate and are relatively small since the initial stages of the game.
-The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are expected to stabilise the system during this learning dynamics.
-Figure 4: $N_{P}$-Chain game: Relative distance (in $\log$-scale) between updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k+1}}^{p})$, top panels, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom panels.
-The evolution of the game given the actions by each player is displayed in Figure 5 for the period $t\in(550,600]$, when the policy updates have been interrupted.
+The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are thus expected to stabilise the system during this learning dynamics.
+The evolution of the game given is displayed in Figure 5 for the period $t\in(550,600]$, after policy updates are interrupted.
 The results demonstrate that the policies obtained by the SLS-BRD routine are capable of jointly stabilising the networked system, robustly against the random noise.
-These policies are also shown to satisfy the operational constraints: Each player’s actions become (roughly) symmetrical and satisfy $|[u_{t}]_{1}+[u_{t}]_{2}|\leq 5$, $t\in\mathbb{N}$, in all scenarios.
+These policies are also shown to satisfy the operational constraints: Each player’s actions satisfy $\|u_{t}^{p}\|_{\infty}\leq 5$, $t\in\mathbb{N}$, in all scenarios.
 The enforcement of this strategy highlights the conservativeness resulting from the robust operational constraints.
@@
 In this experiment, the choice $\beta=(10,40,10)$ seems to be preferable as it converges faster while still achieving satisfactory performance.
-IV-B Price competition in oligopolistic markets
+4.2 Price competition in oligopolistic markets
 Consider a game $\mathcal{G}_{\infty}^{\text{LQ}}{=}\{\mathcal{P},\bm{\mathcal{X}},\{\bm{\mathcal{U}}^{p}\}_{p\in\mathcal{P}},\bm{\mathcal{W}},\{J^{p}\}_{p\in\mathcal{P}}\}$ consisting of a set of companies $\mathcal{P}=\{1,2,3,4\}$ participating in a single-product market.
@@
 $-\bar{u}_{\text{avg}}\leq\textstyle\frac{1}{N_{P}}\textstyle\sum_{p\in\mathcal{P}}u^{p}(t)\leq\bar{u}_{\text{avg}},\quad\bar{u}_{\text{avg}}=0.5.$
-Define $\bm{x}=(\bm{x}^{1},\ldots,\bm{x}^{N_{P}})$, with $\bm{x}^{p}=(d^{p}(t)-\bar{d}_{base})_{t\in\mathbb{R}_{\geq 0}}$, and $\widetilde{B}^{p}=[\widetilde{B}^{p1}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \widetilde{B}^{pN_{P}}]$, for all $p\in\mathcal{P}$.
-Considering a zero-order hold of inputs with period $\Delta t=1/4$ [days], the game $\mathcal{G}_{\infty}^{\text{LQ}}$ can be described by the discrete-time dynamics [^3] With a slight abuse of notation, we use $t$ for both the continuous-time ($x(t)$, $t\in\mathbb{R}_{\geq 0}$) and discrete-time ($x_{t}$, $t\in\mathbb{N}$) signals.
-$x_{t+1}=Ax_{t}+\sum_{p\in\mathcal{P}}B^{p}u^{p}_{t}+w_{t}$
-with $A=\exp(-\tau\Delta t)I_{N_{P}}$ and $\{B^{p}=-\frac{1}{\tau}(A-I_{N_{x}})\widetilde{B}^{p}\}_{p\in\mathcal{P}}$, and the noise process $\bm{w}=\big{(}-\frac{1}{\tau}(A-I_{N_{x}})v_{t}\big{)}_{t\in\mathbb{N}}$ [46].
+Define $\bm{x}=(\bm{x}^{1},\ldots,\bm{x}^{N_{P}})$, with $\bm{x}^{p}=(d^{p}(t)-\bar{d}_{base})_{t\in\mathbb{R}_{\geq 0}}$, and $\widetilde{B}^{p}=[\widetilde{B}^{p1}~{}\cdots~{}\widetilde{B}^{pN_{P}}]$, for all $p\in\mathcal{P}$.
+Considering a zero-order hold of inputs with period $\Delta t=1/4$ [days], the game $\mathcal{G}_{\infty}^{\text{LQ}}$ can be described by the discrete-time dynamics [^1] With a slight abuse of notation, we use $t$ to index both the continuous-time ($x(t)$, $t\in\mathbb{R}_{\geq 0}$) and discrete-time ($x_{t}$, $t\in\mathbb{N}$) signals in this example.
+$x_{t+1}=Ax_{t}+\textstyle\sum_{p\in\mathcal{P}}B^{p}u^{p}_{t}+w_{t}$
+with $A=\exp(-\tau\Delta t)I_{N_{x}}$ and $\{B^{p}=-\frac{1}{\tau}(A-I_{N_{x}})\widetilde{B}^{p}\}_{p\in\mathcal{P}}$, and the noise process $\bm{w}=\big{(}-\frac{1}{\tau}(A-I_{N_{x}})v_{t}\big{)}_{t\in\mathbb{N}}$ [47].
 The companies assume that the baseline demand fluctuations satisfy $w_{t}\in\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$ for all $t\in\mathbb{N}$.
 Under this representation, each player’s objective is formulated as solving for a policy which minimises the functional
-$J^{p}(u^{p},u^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\alpha^{p}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
+$J^{p}(\bm{u}^{p},\bm{u}^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\alpha^{p}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
 given weights $\alpha^{p},\beta^{p}\in\mathbb{R}_{\geq 0}$.
@@
 $\mathcal{U}^{p}$ $=\mathbb{R}^{N_{u}^{p}};$
-$\mathcal{U}_{\mathcal{G}}$ =$=\big{\{}u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:\textstyle\frac{1}{N_{P}\cdot\bar{u}_{\text{avg}}}\big{[}\mathbf{1}_{N_{u}}\leavevmode\nobreak\ {-}\mathbf{1}_{N_{u}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}u_{t}\leq[1\leavevmode\nobreak\ 1]^{{\mkern-1.5mu\mathsf{T}}}\big{\}},$
+$\mathcal{U}_{\mathcal{G}}$ =$=\big{\{}u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:\begin{bmatrix}\tfrac{1}{N_{P}\cdot\bar{u}_{\text{avg}}}\bm{1}_{N_{u}}^{{\mkern-1.5mu\mathsf{T}}}\end{bmatrix}u_{t}\leq 1\big{\}}.$
 We consider that players design their state-feedback policies, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$, with a FIR horizon of $N=16$ and no structural constraints, that is, $\bm{S}_{x}=I$ and $\bm{S}_{u}^{p}=I$ ($\forall p\in\mathcal{P}$).
 In practice, this implies that companies have perfect information of any demand $\bm{x}^{p}$ and price $\bm{u}^{p}$ changes in the market.
-As in Section IV-A, we simulate an instance of the game $\mathcal{G}_{\infty}^{\text{LQ}}$ alongside the SLS-BRD routine (Algorithm 4) with players using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, given $\gamma=0.95$.
+As in Section 4.1, we simulate an instance of the game $\mathcal{G}_{\infty}^{\text{LQ}}$ alongside the SLS-BRD routine (Algorithm 4) with players using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, given $\gamma=0.95$.
 Policies are updated simultaneously every $\Delta T=1$ stage with learning rate $\eta=1/4$.
 Under the above setup, the game $\mathcal{G}_{\infty}^{\text{LQ}}$ is not a potential game and thus a GFNE cannot be easily computed in advance.
-Finally, we remark that solutions to this problem are neither unique nor necessarily admissible: Different initial profiles result in convergence to different fixed-points, which might be unfair to a subset of players.
-The convergence of the SLS-BRD routine to a fixed-point $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 6.
-Since the exact fixed-point to which the routine will converge is not known for this problem, we let $\bm{\Phi_{u}}^{*}\approx\bm{\Phi_{u,k_{f}}}$ for $k_{f}=828$ (when the policy updates are interrupted) and analyse the convergence with respect to this point.
-In this case, the iterates approach the fixed-point mostly at a geometric rate with the policies requiring roughly $800$ updates (or $200$ days in-game) before changes become numerically negligible.
-The rate of convergence is also observed to slow around $k=15$, when the generalised constraints ($\mathcal{U}_{\mathcal{G}}$) become active.
+Finally, we remark that solutions to this problem are not unique: Different initial profiles may result in convergence to different equilibrium profiles.
+Figure 6: Market game: Convergence of the SLS-BRD routine.
+Figure 7: Market game: Relative distance between the local updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p})$, left row, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, right row.
+The convergence of the SLS-BRD routine to a fixed-point $\bm{K}^{\star}=\bm{\Phi_{u}}^{\star}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{\star}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\star}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 6.
+Since the exact fixed-point to which the routine will converge is not known for this problem, we let $\bm{\Phi_{u}}^{\star}\approx\bm{\Phi_{u,k_{f}}}$ for $k_{f}=1080$ (when the policy updates are interrupted) and analyse the convergence with respect to this point.
+In this case, the iterates approach the fixed-point mostly at a geometric rate with the policies requiring roughly $1000$ updates (or $250$ days in-game) before changes become numerically negligible.
 We stress that the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ cannot be (globally) contractive in this case, as the set of fixed points $\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ is not a singleton.
-Figure 6: Market game: Convergence of the SLS-BRD routine.
-In Figure 7 (top), the relative distances between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ are displayed.
+Finally, this experiment imply that multi-agent markets might require a half-year of learning dynamics (if policies are updated every $\Delta t=1/4$ [days]) before converging to an equilibrium.
+In Figure 7 (left), the relative distances between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ are displayed.
 As in the previous example, the updates are of similar magnitude for all players $p\in\mathcal{P}$ and they become numerically negligible at a faster rate than the global convergence in Figure 6.
@@
 Thereafter, these relative distances continue to decrease at a steady rate.
-In Figure 7 (bottom), the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ are shown.
+In Figure 7 (right), the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ are shown.
 As before, these distances decrease at a similar rate and are relatively small since the initial stages of the game.
-The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate the stabilising profile $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are expected to be stabilising for all iterations.
-Figure 7: Market game: Relative distance between updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k+1}}^{p})$, top panels, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom panels, for all $p\in\mathcal{P}$.
-The evolution of the game given each player’s actions is displayed Figure 5 for the in-game period $t\in(260,400]$ days, after the policy updates have been interrupted.
-We compare the performance of the policy profile $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}$ with the evolution obtained through the open-loop operation $u(t)=0$.
+The evolution of the game given each player’s actions is displayed Figure 8 for the in-game period $t\in(280,420]$ days, after the policy updates have been interrupted.
+We compare the performance of the policy profile $\bm{K}^{\star}=\bm{\Phi_{u}}^{\star}\bm{\Phi_{x}}^{-1}$ with the evolution obtained by the open-loop operation $u(t)=0$.
 During this period, we simulate a worst-case fluctuation on the baseline demand for the companies’ product by defining
-$w(t)=\begin{cases}(\phantom{-}1,\phantom{-}1,\phantom{-}1,\phantom{-}1)&\text{for}\leavevmode\nobreak\ t\in[265,279];\\ (\phantom{-}1,-1,\phantom{-}1,-1)&\text{for}\leavevmode\nobreak\ t\in[293,307];\\ (-1,\phantom{-}1,-1,\phantom{-}1)&\text{for}\leavevmode\nobreak\ t\in[321,335];\\ (\phantom{-}1,\phantom{-}1,-1,-1)&\text{for}\leavevmode\nobreak\ t\in[349,363];\\ (-1,-1,-1,-1)&\text{for}\leavevmode\nobreak\ t\in[377,391],\end{cases}$
+$w(t)=\begin{cases}(\phantom{-}1,\phantom{-}1,\phantom{-}1,\phantom{-}1)&\text{for}~{}t\in[285,299];\\ (\phantom{-}1,-1,\phantom{-}1,-1)&\text{for}~{}t\in[313,327];\\ (-1,\phantom{-}1,-1,\phantom{-}1)&\text{for}~{}t\in[341,355];\\ (\phantom{-}1,\phantom{-}1,-1,-1)&\text{for}~{}t\in[369,383];\\ (-1,-1,-1,-1)&\text{for}~{}t\in[397,411],\end{cases}$
 and $w(t)=0$ otherwise.
 The results show that the policy profile obtained by the SLS-BRD routine allows the companies to efficiently respond to changes in their local demands.
-In general, the players coordinate price changes to alleviate the deviations from the baseline demand caused by the disturbances.
-Moreover, the player’s actions satisfy the price-cap constraint $\frac{1}{N_{P}}|\sum_{p\in\mathcal{P}}u^{p}(t)|\leq 0.5$ in all disturbance scenarios.
-The best performance is observed for companies $p\in\{3,4\}$, whereas $p=1$ behaves noticeably worse than all players (especially for $w(t)=(1,-1,1,-1)$ and $w(t)=(-1,1,-1,1)$, when the open-loop operation seems to attain better results).
-This highlights the fact that fixed-points obtained through the SLS-BRD routine are not necessarily admissible GFNE, and thus might be unfavourable for a subset of players.
-Finally, we note that these policies are still not able to completely reject the effect of the extreme disturbances: Achieving zero-offset would require incorporating integral action into the player’s policies.
-Figure 8: Market game, $t\in(260,400]$ days: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for an execution of the game $\mathcal{G}_{\infty}^{\text{LQ}}$.
-The dashed lines refer to the state-trajectories resulting from an open-loop operation of the market with $u(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$
-V Concluding remarks
-This work presents the SLS-BRD, an algorithm for generalised feedback Nash equilibrium seeking in $N_{P}$-players non-cooperative games.
-The method is based on the class of best-response dynamics algorithms for decentralised learning and consists of players updating and announcing a parametrisation of their policies until converging to an equilibrium.
+In general, the players coordinate price changes to alleviate the deviations from the baseline demand caused by the disturbances, while still satisfying the price-cap constraint $\frac{1}{N_{P}}|\sum_{p\in\mathcal{P}}u^{p}(t)|\leq 0.5$.
+The best performance is observed for the companies $p\in\{3,4\}$, whereas $p=1$ behaves noticeably worse than all players (especially for $w(t)=(1,-1,1,-1)$ and $w(t)=(-1,1,-1,1)$, when the open-loop operation attains better results).
+This highlights the fact that fixed-points obtained through the SLS-BRD routine are not necessarily admissible GFNE, and might be unfavourable for a subset of players.
+Finally, we note that these policies are still not able to completely reject the effect of the noise: Achieving zero-offset requires incorporating integral action into the feedback policies.
+Figure 8: Market game, $t\in(280,420]$ days: State $\bm{x}$ (top panels) and applied control $\bm{Bu}$ (bottom panels) trajectories for an execution of the game $\mathcal{G}_{\infty}^{\text{LQ}}$.
+The dashed lines refer to the state-trajectories resulting from an open-loop operation of the market with $u(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$.
+5 Concluding remarks
+This work presents the SLS-BRD, an algorithm for generalised feedback Nash equilibrium seeking in $N_{P}$-player noncooperative games.
+The method is based on the best-response dynamics class of algorithms for Nash equilibrium seeking and consists of players updating and announcing a parametrisation of their policies until converging to a fixed-point.
+Our approach leverages the System Level Synthesis framework to formulate each player’s best-response map as the solution to robust finite-horizon optimal control problems.
 Because not updating control actions explicitly, this learning dynamics can be performed alongside the execution of the game.
-Using results from operator theory, we propose that convergence to a Nash equilibrium is geometric if certain conditions are met: Namely, that the best-response maps are contractive operators.
-Our approach leverages the System Level Synthesis framework to formulate each player’s best-response map as the solution to robust finite-horizon optimal control problems.
-This framework benefits the SLS-BRD algorithm also by allowing richer information patterns to be enforced directly at the synthesis level.
-After the main theoretical and practical aspects are discussed, the algorithm is demonstrated on an exemplary problems on the decentralised control of multi-agent systems.
+This framework also benefits the SLS-BRD by allowing richer information patterns to be enforced directly at the synthesis level.
+Using results from operator theory, we then established sufficient conditions for this procedure to converge to a generalised feedback Nash equilibrium.
+After the theoretical aspects are discussed, the algorithm is showcased on exemplary problems on the noncooperative control of multi-agent systems.